### v2

# HCT-QA: A Benchmark for Question Answering on Human-Centric Tables

Abstract
Tabular data embedded in PDF files, web pages, and other document types is prevalent in different sectors such as government, science, and business. These tables, which we call human-centric tables (HCTs), are dense in information but often exhibit both structurally and semantically complex layouts. This complexity poses significant challenges to querying them using traditional data processing systems. While some existing solutions focus on transforming these tables into relational formats, they fail to handle the diverse and complex layouts of HCTs, making them not amenable to easy querying with SQL-based approaches. Another promising option is to use Large Language Models (LLMs) and Vision Language Models (VLMs). To settle this debate and make progress on answering queries on this type of table, we propose HCT-QA, an extensive benchmark of HCTs, natural language queries, and their respective answers on thousands of tables. Our dataset includes 1,880 real-world HCTs with 9,835 QA pairs in addition to 4,679 synthetic HCTs with 67.7K QA pairs. Furthermore, we show through a series of extensive experiments the performance of a large number of LLMs (21) and VLMs (13) in properly answering the different questions in our benchmark. We also show how fine-tuned models can significantly outperform their pretrained versions on this task.
1 Introduction
Tabular data embedded in PDF files, web pages, and other types of documents are commonly present in numerous domains. Figure 1 shows a few examples of such data obtained from the US Census (uscensus, ), the Pakistan Bureau of Statistics (pakcensus, ), the Qatar National Planning Council (qnpc, ), and scientific paper archives (arxiv, ). These tabular data representations have rich semantics and are usually used to make strategic decisions without revisiting the raw, low-level data, often unavailable to the end user. They represent a type of data modality that encompasses a unique set of characteristics that make them very valuable to businesses, but extremely challenging to process systematically.
Challenging layouts.
The data is visualized in a non-standard, often complex tabular layout with artistic adornment to deliver a wealth of information summarized in a concise and dense way targeted primarily for direct human understanding and consumption. In Figure 1, we label several of the common table layout properties for ease of reference throughout the paper, including column nesting (L1), row nesting (L2) or row group labels (L3), embedded aggregations either row-wise (L4) or column-wise (L5). These properties may rely on visual traits such as differentiating text or row colors (L3 in table (d) & L4 in table (c)), bold text (L4 in tables (a) & (c)), or positioning within the cell to carry some semantic meaning (L2 in table (b)). Some other layout properties include nested table cells, local aggregations for specific subset of columns or rows, cell dividers, and row-group clustering (See examples in Appendix A). Although suitable for humans, such complex layouts introduce numerous challenges for systematic and automated methods for processing and querying the data. For instance, these layouts make it challenging to automatically answer questions like “What is the amount of dates imported in 2015?” (Table (a)), or “Which subject had the highest number of hardcover books sold in 2000” (Table (c)).
Untapped huge potential.
Tabular data embedded in documents are often treated as terminal data, i.e., no further processing is expected beyond eyeball analysis, which limits their full potential, especially at scale. For example, a single organization might have tens of thousands of such tables, and a big data lake could hold millions eltabakh2023crossmodaldatadiscovery . Moreover, for many applications, these tabular representations (and their container documents) are the only source of such aggregated data. Indeed, the base data (e.g., relational databases, CSVs, or Excel sheets) may not be available due to undocumented or lost sources, restricted access to the underlying databases, or missing ETL generating processes. Additionally, the content in these tables could be computed and assembled from multiple disjoint sources and augmented with external domain-expert knowledge. Hence, such data is unique and highly valuable but hardly exploitable at scale. In the rest of the paper, we refer to this tabular modality intended primarily for human consumption as “Human Centric Tables” ((HCTs for short)).
Figure 1: Examples of Human-Centric Tables (HCTs) from different real-world data sources exhibiting various levels of structural and semantic complexity: column nesting (L1); row nesting (L2); row group labels (L3); row aggregation (L4); column aggregation (L5).
Although tools exist to extract HCTs from PDF documents or web pages aspose ; innodata ; pdftable-tableextractor ; pdfplumber , end-users are on their own with no systematic support for any further processing or analytics. Research works typically study ways to transform non-relational layouts into relational tables using heuristics or data-driven transformation techniques AutoTables_57 ; retclean-paper ; tabnet-paper ; tab-cleaner-paper , then use NL-to-SQL approaches NLtoSQLsurvey2024 to query them. This process is complex, costly, error-prone, and may not work for elaborate HCTs.
A new perspective.
Recent advances in Large Language Models (LLMs) and Vision Language Models (VLMs) motivate a new “LLM-oriented” approach for querying HCTs, which could (1) efficiently handle HCTs among other human-centric data formats (images, charts, videos…), (2) avoid the complex and error-prone NL-to-SQL pipeline, and (3) directly support natural language queries, particularly crucial for multilingual tables such as the ones from the Qatar National Planning Council (Figure 1a). However, a systematic evaluation for identifying the strengths, weaknesses, and potential improvements of LLM-oriented approaches remains uncharted. Hence, we propose a novel benchmark for assessing LLMs as the engine for understanding, processing, and querying HCTs.
Challenges and contributions. Devising the proposed benchmark involves several challenges: (1) collecting or generating HCTs with an extensive diversity of structures from relational table layouts to very complex layouts as illustrated in Figure 1; (2) collecting or generating an extensive variety of natural language questions that require steps from simple selection to complex aggregation, joins, and ranking questions; (3) getting complete and accurate ground truth answers to each table and question pair; and (4) exploring strategies like fine-tuning and impact on performance. Here are our key contributions:
  • We propose a large HCT and question-answer (QA) benchmark containing 1,880 HCTs from four real-world data sources spanning over fifty semantic domains, both as images and extracted CSV formats, along with 9,835 QA pairs crafted by experts.
  • We develop a novel synthetic HCT generator to complement the real-world datasets with 4,679 HCTs, along with 67,747 template-based QA pairs covering seven semantic domains.
  • We propose layout and question properties capturing the most common elements of HCTs and natural language questions to identify performance gaps and opportunities for improvement of processing engines.
  • We use our benchmark and properties to evaluate $21$ LLMs and $13$ VLMs ranging from $3B$ to $100B+$ parameters. Our evaluation reveals that large closed-weight models perform the best overall, gpt-o1-mini being the best-performing model with a Recall score of 68.5%. Also, the small Pixtral-12B VLM outperforms the best Qwen2.5-7B LLM of comparable size by 7 points in Recall, showing promise of processing HCT images directly.
  • We fine-tuned two models and showed that fine-tuning is indeed effective and can improve performance by up to 50 points compared to the pretrained versions.
Our benchmark is publicly available on HuggingFace (HCT_QA_hf, ). All experimentation codes and the synthetic HCT generator are publicly available on our GitHub repository (HCT_QA_github, ) .
2 Related work
There is a flurry of research and commercial systems on document understanding and information extraction from complex documents (especially PDF) Ahmed2016_56 . However, none of this work tries to answer the questions targeted in the proposed benchmark, i.e., the potential and the hurdles that face the LLM-based approach as the query engine for HCTs.
Table extraction techniques transform tables from PDF documents or images into machine-readable format like CSV, HTML, and XML. There are various commercial tools  (aspose, ; azure_document_intelligence, ; innodata, ; docugami, ) and recent literature (tableformer_2022, ; robustTableDetection_Ma_2022, ; omniParser_2024, ; smolDocling_2025, ; Khan2021ICDAR_13, ; Oro2008ICDAR_14, ; Qiao2021ICDAR_22, ) that show promise on extracting tables like HCTs into these formats. Nevertheless, beyond the extraction step, end-users are on their own with no systematic support for further processing, querying, analytics, or discovery.
NL-to-SQL Weir2020ACM_25 ; Basik2018ACM_26 ; Wang2019_27 ; Brunner2021ValueNetAN_28 ; sen2020VLDB_29 ; liu2024surveynl2sqllargelanguage These techniques assume clean and simple relational tables as input and focus on translating natural language queries to SQL. Several techniques are proposed to transform non-relational tabular data into relational tables 10.1431751.3231766 ; patsql_vldb2021 ; 10.147407790.3407831 ; AutoTables_57 . However, these techniques deal with much simpler layouts compared to the HCTs presented in Figure 1, and the transformation accuracy reported in the most recent work AutoTables_57 is about 60%.
Existing Table QA. Most existing Table QA benchmarks use flat Wikipedia tables which, when compared to HCTs, are significantly simpler to understand tablebench_2025 ; wikitablequestions_2017 ; fetaqa_2021 ; tabfact_2020 ; ottqa_2021 ; hybridqa_2021 . Other benchmarks use tables from one or two semantic domains and use a very small number of tables.  aitqa_2021 ; TAT-QA ; finqa_2022 ; scitat_2024 ; multitat_2025 ; tablevqa_2024 ; spiqa_neurips2024 . Some benchmarks focus on QA on documents that include different modalities such as tables, charts, and text paragraphs but not on isolated HCTs TAT-QA ; scitat_2024 ; MMLongbenchdoc_neurips2024 . In comparison, HCT-QA uses 1,880 real HCTs from 4 data sources and 4,679 synthetic HCTs that cover 50+ domains, and have questions that require only the HCT as the context to be answered.
3 HCT-QA benchmark datasets
In this section we describe the construction of our benchmark datasets. Each element in the datasets consists of a triplet in the form of $(T_{HCT},Q_{NL},A)$ along with rich metadata for facilitating deeper analysis. We focus on the real-world HCT datasets collected from thousands of PDF documents from diverse domains. We present the extraction and transformation processes that we applied to them, and the most common table patterns present in such data. We then introduce our proposed characterization of query properties. Finally we present the synthetic table and QA generator. Further details are given in Appendix A,  Appendix B, and Appendix C
Overall, our benchmark consists of real-world and synthetic tables and QA summarized in Figure 2.
3.1 Real-world dataset collection
Data sources:
We collected tens of thousands of PDF documents from diverse sources: Qatar National Planning Council (QNPC) qnpc (Figure 1(a)), ArXiv Research Papers from ArXiv.org, biorxiv.org, and medarch.org (Figure 1(b)), US Census covering PDF reports from the US Census Bureau uscensus (Figure 1(c)) and Pakistan Census from the official website of the Pakistan government pakcensus (Figure 1(d)). These sources cover more than 50 domains ranging from education, disabilities, land usage, capital gain, machine learning, vision, data management, finance, astrophysics, nanoscale and nuclear physics, materials science, phenomenology, to cryptography and security, lipid disorders, hypertension, diabetology, preventive cardiology, and many others.
Extraction and transformation:
We propose a canonical representation of HCTs in CSV format wherein (i) for column nesting (L1) all parent cell values are above the left most child cell, (ii) for row nesting (L2) all parent cell values are in the same row as the top-most child cell, (iii) row group labels (L3) are added in the left most cell of that row with the remaining row cells as blank (L3 $\rightarrow$ L2), and (iv) single cell values in borderless tables that span multiple lines are represented only in the cell that corresponds to the top line (see Appendix E for illustrative examples). Other properties, such as aggregation rows (L4) and columns (L5) are explicitly represented as regular well-formatted rows and columns in the CSV. This representation in CSV format allows for easy understanding of the table hierarchy and enables further rule-based transformations on the HCT if needed.
After exploring various open-source and commercial tools such as Tabula (tabula-py, ), PDF Plumber pdfplumber , Azure AI Document Intelligence azure_document_intelligence , Aspose aspose , Docling smolDocling_2025 ; doclingModels_2025 , CascadeTabNet (cascadeTabNet_2020, ), we found that Azure AI Document Intelligence was satisfactory in its consistency at outputting the HCT image in PNG or JPG format and its canonical CSV representation, and being the cheapest and most convenient to use as an API service. See Appendix E for details on the extraction performance.
Characterization of table properties:
Across domains and sources, we found that certain structural layouts are very dominant (e.g., column nesting, and row nesting) compared to rare and outlier patterns (e.g., row group labels that are also local row aggregations). In the current benchmark, we focused only on the common patterns and excluded tables exhibiting outlier patterns. We determined properties to characterize HCTs related to the number of levels and symmetry of column and row headers (L1, L2, L3), the presence of aggregate rows or columns (L4, L5) explicitly labeled or not, and the indentation structure to display them, as well as split lines or color adornments to distinguish semantically similar areas in the table. Figure 2 gives the distribution of main Real-World HCT structural properties (a,d). The full list of properties and the web interface we developed to annotate them manually are described in Appendix A.
Refer to caption Refer to caption
(a) Statistics of HCTs and QA pairs (b) Main properties of real-world QA pairs
Refer to caption Refer to caption
(c) Main properties of synthetic HCTs (d) Main properties of real-world HCTs
Figure 2: Main properties of the HCTs and QA pairs in our benchmark datasets.
3.2 Query and ground-truth answer generation
In our context, a query on an HCT object, is a natural language question to extract some information, either directly (e.g., selecting some values from the table), or after applying some computations (e.g., expression, aggregation). We focus on single-table queries. Our benchmark has two types of queries: (1) Human-Generated are curated by domain experts; (2)  Model-Generated are generated by an LLM. All query-answer pairs underwent a human verification to ensure high quality.
Characterization of query properties:
We developed a set of properties, inspired by typical SQL operations, that can be used to describe each query. These properties capture information about the number and nature of the operations, such as aggregation and filtering, that are required to answer the query. For instance, the number of filtered rows (WHERE clause) and projected columns (Condition), the number of returned columns (SELECT clause), or more complex properties involving aggregations and ranks (see Appendix B for more details). These properties are useful in understanding what sorts of queries models may struggle with the most. In our benchmark, properties are only provided for human-generated queries.
Human-Generated Queries:
To generate a query and its corresponding answer, the domain expert uses a custom annotation tool (Appendix B) to see the image of the table, create the query and answer, and annotate the properties of that query. We also allow the user to select among a set of pre-defined query templates to create a question that adheres to the query properties of that template and find its answer in the HCT, enabling automatic filling of the query properties. The expert may see the type of queries they have created so far as a nudge to create a more balanced set of queries that require various operators, expressions, and aggregations. Each query was verified by another domain expert for quality control. We ensure that each HCT has at least one human-generated query. We created 3,248 such queries on Real World HCTs and found that generating each query takes around 2.5 minutes.
Model-Generated Queries:
To expand our query set, we leverage the OpenAI’s GPT model  openai2023gpt4-turbo . The model is provided with the HCT image, an example human-generated query-answer pair (one-shot prompting), and instructions containing the constraints the query must adhere to. The model is then prompted to return a set of 8 query-answer pairs for each table. This method allows us to scale the number of queries very well. We generated 6,587 such queries on Real World HCTs in approximately 30 minutes.
Human Verification:
Model-generated queries were given to a team of annotators to: (i) rate the quality of the question as ”good”, ”ambiguous”, ”nonsensical” and (ii) given the table and query, correct the corresponding answer if it is wrong (check Appendix B for details). We dropped ”nonsensical” queries, rewrote ”ambiguous” queries and kept the ”good” ones. Each query was independently verified by two annotators. We included the human-generated queries as well to gauge the quality of the annotators and removed annotators that marked more than 1 manually generated queries as ”nonsensical” or its answer as incorrect. This process ensured that we have a set of high-quality query-answer pairs. We found that only 4% of model-generated queries were marked as ”nonsensical”, 24% as ambiguous, and 72% as ”good”. Moreover, the model required answer corrections for nearly 50% of the queries. Finally, we obtained $1,880$ tables and $9835$ related question and answer pairs from QNPC, Pakistan Census, US Census, and ARXIV real-world sources (Figure 2).
3.3 Synthetic tables and QA generation
To expand the dataset, we developed a synthetic data generation tool to produce HCTs in CSV, HTML, and PDF formats, along with their queries and answers, aw well as their properties. Synthesizing datasets in benchmarking is significantly useful in various ways including filling in gaps such as specific table patterns that do not sufficiently exist in real-world data, stress testing certain combinations, scaling up the data size, and finally by controlling the data generation (for both HCTs and their associated queries) the ground truth answers become computationally known without the need for manual validation.
We generated $4,679$ unique synthetic HCTs ($T_{HCT}$) across $7$ semantics, with up to $15$ template-based natural language questions $Q_{NL}$ per table, leading to $67,747$ triplets $(T_{HCT},Q_{NL},A)$ (Figure 2) with the corresponding ground-truth answer $A$. Each synthetic HCT contains no duplicate values to avoid ambiguities when evaluating the matching between the ground-truth and LLM-predicted answers. In practice, HCTs are produced by pivoting pivotTable2005 standard relational tables $T_{REL}$ that we first generated from custom semantic data templates. Attributes of $T_{REL}$ are randomly selected and ordered for the pivoting process (pivottabler, ) to create a variety of the HCT properties described in Section 3.1, such as forming nested rows and columns of $T_{HCT}$, as well as adding, naming, and positioning aggregate rows or columns. We automatically get the ground-truth answer $A$ to a SQL query $Q_{SQL}$ on the relational table $T_{REL}$ using a $SQL$ engine. We then transform the SQL query $Q_{SQL}$ into $Q_{NL}$ using custom templates to get the desired triplet $(T_{HCT},Q_{NL},A)$.
Our synthetic generator can be customized to produce 16 out of the 18 layout characteristics detailed in subsection 3.1, and the table semantics can be easily extensible to any domain. Further details are given in Appendix C.
4 Experimental evaluation
Our goal is to build a benchmark dataset on which one can test how natural language queries can be answered on HCTs using any data processing method or engine. In our experiments, we evaluate our benchmark on Large Language Models (LLMs) and Vision Language Models (VLMs). We want to answer the following questions: (RQ1) How do these models perform overall on HCT-QA? (RQ2) Are models improving over time? (RQ3) Does the input format of the HCT matter in text-modality? (RQ4) Does fine-tuning models for this task improve performance? (RQ5) How do HCT properties affect model performance? We also explored four additional questions (described in Appendix D.2): (RQ6) Does model verbosity affect the output correctness? (RQ7) How do question properties affect model performance? (RQ8) How are model size and performance correlated? (RQ9) In what cases do VLMs perform better than LLMs?
4.1 Experimental setup
Models.
We tested our benchmark on a large variety of open- and closed-weight LLMs and VLMs. Details about these models are given in Table 8 (Appendix D.1).
Datasets.
Our benchmark comprises four real-world data sources (Real World HCTs) and one synthetically generated dataset (Synthetic HCTs) with their corresponding QA Pairs (Table 2). There are 1,880 real world HCTs with 9,835 QA pairs on them and 4,679 synthetic HCTs with 67,747 queries on them (total of 6,559 HCTs with 77,582 queries).
Prompting approach and hyperparameters.
We experimented with zero-shot and one-shot prompting and found that there was no clear indication that the latter improved model performance. Thus, for all subsequent experiments, we used zero-shot prompting. Additional details on the prompting approaches and tests can be found in Appendix D.3. The temperature was set to 0, and the maximum output tokens to 128. We post-process the model responses using a rule-based approach to extract the actual answer if one was provided. All inference and fine-tuning experiments were run on 4 NVIDIA A100 GPUs.
Evaluation metrics.
We used three standard metrics to evaluate the performance of the model: (i) Precision, (ii) Recall, and (iii) Complete Containment (CC Score), which equals 1 if the model outputs the complete answer (Recall = 100%) and 0 otherwise. It is important to note that Precision may be unfairly deflated due to the verbose nature of language models. In the prompt, we specify outputting the answer in an easy-to-parse set-like format. However, although most models follow the specified set format, Precision would not capture the correctness of the answer if the model provides it as part of a sentence. Additionally, we checked the robustness of using an LLM-as-a-judge for evaluation and found it to closely align with the CC Score metric. We thus opted not to use it (see Appendix D.4). Our analysis focuses primarily on Recall scores to compare different models.
Table 1: Overall performance of models. All text-only and vision-text models are evaluated on the appropriate datasets. Confidence intervals are given in Appendix D.7
Model Name QNPC PAK US CENSUS arXiv Synthetic Tables All Data
Precision Recall CC Score Precision Recall CC Score Precision Recall CC Score Precision Recall CC Score Precision Recall Precision Recall
TEXT-ONLY LLMs
Large Models – 100B+ (Closed)
gpt-o1-mini 0.67 0.665 0.617 0.741 0.742 0.708 0.775 0.756 0.716 0.74 0.733 0.699 - - 0.691 0.685
gpt-4o 0.627 0.606 0.515 0.691 0.677 0.6 0.601 0.591 0.535 0.646 0.63 0.565 0.521 0.439 0.635 0.616
gpt-35 0.412 0.415 0.315 0.443 0.447 0.356 0.375 0.395 0.333 0.45 0.468 0.386 0.391 0.309 0.419 0.425
Average (Large) 0.57 0.562 0.482 0.625 0.622 0.555 0.584 0.581 0.528 0.612 0.61 0.55 0.304 0.249 0.582 0.575
Medium Models – 27B-70B (Open)
Qwen2.5-72B-Instruct 0.574 0.582 0.479 0.622 0.629 0.541 0.505 0.519 0.465 0.594 0.595 0.52 0.502 0.491 0.579 0.586
Llama-3.1-70B-Instruct 0.489 0.474 0.374 0.531 0.515 0.413 0.476 0.473 0.411 0.469 0.454 0.386 0.451 0.427 0.49 0.475
Llama-3.3-70B-Instruct 0.467 0.494 0.401 0.531 0.555 0.475 0.456 0.489 0.431 0.453 0.476 0.415 0.402 0.436 0.471 0.497
gemma-2-27b-it 0.473 0.466 0.356 0.536 0.525 0.415 0.396 0.406 0.351 0.479 0.464 0.379 0.378 0.322 0.477 0.469
gemma-3-27b-it 0.441 0.424 0.338 0.5 0.479 0.404 0.419 0.417 0.359 0.46 0.44 0.377 0.34 0.297 0.448 0.431
Mixtral-8x7B-Instruct-v0.1 0.279 0.278 0.197 0.307 0.307 0.229 0.248 0.274 0.227 0.306 0.323 0.259 0.17 0.159 0.284 0.286
Average (Medium) 0.454 0.453 0.358 0.505 0.502 0.413 0.417 0.43 0.374 0.46 0.459 0.389 0.374 0.355 0.458 0.457
Small Models – 3B-12B (Open)
Meta-Llama-3-8B-Instruct 0.309 0.285 0.181 0.324 0.293 0.201 0.237 0.245 0.191 0.269 0.255 0.184 0.235 0.181 0.302 0.28
Llama-3.1-8B-Instruct 0.223 0.215 0.138 0.235 0.214 0.127 0.204 0.204 0.165 0.248 0.237 0.177 0.153 0.132 0.227 0.218
Qwen1.5-7B-Chat 0.181 0.168 0.14 0.188 0.172 0.145 0.156 0.157 0.14 0.194 0.184 0.169 0.138 0.104 0.183 0.17
Qwen2-7B-Instruct 0.24 0.297 0.208 0.274 0.309 0.214 0.227 0.288 0.243 0.235 0.302 0.251 0.278 0.266 0.242 0.299
Qwen2.5-7B-Instruct 0.35 0.38 0.264 0.396 0.429 0.311 0.27 0.336 0.284 0.337 0.373 0.291 0.257 0.242 0.35 0.382
gemma-2-9b-it 0.377 0.353 0.243 0.409 0.368 0.254 0.287 0.299 0.253 0.347 0.345 0.265 0.349 0.275 0.373 0.351
gemma-3-12b-it 0.421 0.372 0.261 0.471 0.41 0.297 0.355 0.352 0.295 0.386 0.352 0.28 0.425 0.29 0.418 0.372
Mistral-7B-Instruct-v0.3 0.208 0.201 0.151 0.216 0.221 0.174 0.188 0.205 0.176 0.215 0.222 0.185 0.144 0.115 0.209 0.206
Mathstral-7B-v0.1 0.274 0.263 0.192 0.301 0.284 0.213 0.266 0.268 0.233 0.29 0.278 0.237 0.211 0.19 0.278 0.267
FanarC 0.27 0.243 0.178 0.238 0.21 0.151 0.271 0.26 0.235 0.273 0.258 0.215 0.179 0.145 0.267 0.243
aya-23-8B 0.113 0.098 0.069 0.111 0.095 0.063 0.11 0.101 0.077 0.134 0.135 0.111 0.068 0.058 0.115 0.103
Phi-3.5-mini-instruct 0.114 0.102 0.078 0.137 0.118 0.092 0.081 0.076 0.062 0.161 0.145 0.115 0.162 0.125 0.121 0.108
Average (Small) 0.257 0.248 0.175 0.275 0.26 0.187 0.221 0.233 0.196 0.257 0.257 0.207 0.195 0.157 0.257 0.25
VISION-TEXT VLMs/MLMs
Large Models – 100B+ (Closed)
gpt-4o-vision 0.486 0.615 0.513 0.583 0.704 0.613 0.31 0.396 0.328 0.571 0.674 0.597 - - 0.487 0.623
Small Models – 3B-12B (Open)
Pixtral-12B-2409 - 0.435 0.312 - 0.532 0.405 - 0.4 0.344 - 0.547 0.465 - - - 0.459
InternVL2-4B - 0.29 0.193 - 0.372 0.262 - 0.254 0.202 - 0.385 0.298 - - - 0.309
InternVL2_5-8B-MPO - 0.239 0.152 - 0.302 0.203 - 0.228 0.181 - 0.376 0.298 - - - 0.264
Phi-3-vision-128k-instruct - 0.408 0.3 - 0.445 0.338 - 0.273 0.23 - 0.442 0.357 - - - 0.411
Phi-3.5-vision-instruct - 0.244 0.159 - 0.365 0.262 - 0.178 0.14 - 0.375 0.29 - - - 0.272
Idefics3-8B-Llama3 - 0.382 0.269 - 0.458 0.337 - 0.351 0.284 - 0.438 0.354 - - - 0.396
Qwen2-VL-7B-Instruct - 0.06 0.045 - 0.066 0.057 - 0.071 0.052 - 0.045 0.038 - - - 0.059
llava-1.5-7b-hf - 0.094 0.068 - 0.109 0.095 - 0.085 0.062 - 0.067 0.056 - - - 0.091
llava-v1.6-mistral-7b-hf - 0.139 0.092 - 0.174 0.119 - 0.105 0.088 - 0.15 0.113 - - - 0.143
llava-v1.6-vicuna-7b-hf - 0.158 0.106 - 0.204 0.136 - 0.132 0.098 - 0.169 0.125 - - - 0.163
Molmo-7B-D-0924 - 0.376 0.276 - 0.472 0.368 - 0.282 0.225 - 0.465 0.383 - - - 0.394
paligemma2-10b-ft-docci-448 - 0.102 0.066 - 0.111 0.081 - 0.093 0.062 - 0.065 0.049 - - - 0.098
Average (Small) - 0.244 0.17 - 0.301 0.222 - 0.204 0.164 - 0.294 0.235 - - - 0.255
4.2 Analysis and findings
(RQ1) How do these models perform overall on HCT-QA?
As we can see in Table 1, the latest versions of large closed-weight models like gpt-o1-mini (text modality) perform reasonably well, achieving 69% Precision and 68% Recall. This shows that even for state-of-the-art models, there is ample room for improvement on this task. Furthermore, Qwen2.5-72B-Instruct, a considerably smaller model, achieves a Recall of 58%, which is only 3 points less than the Recall of gpt-4o. This showcases the potential for medium-sized models in this task. On average, large models achieve a Recall of 57% compared to 45% for medium-sized models, showing that the performance gap between models of medium and large sizes is still small. However, with an average Recall of 25%, small models lag behind their larger versions. Small vision models with an average Recall of 25.5% are on par with small text-only models. But Pixtral-12B VLM outperforms the best Qwen2.5-7B LLM of comparable size by 7 points in Recall, emphasizing the potential of vision models in capturing the structural complexity of HCTs that may be missed in a textual representation (check Appendix D.5 for details). Overall, there remains clear room for improvement for both LLMs and VLMs.
(RQ2) Are models improving over time?
Newer models are improving their reasoning capabilities and thus their performance on HCT-QA. The Qwen family of models shows continuous improvement as Qwen1.5-7B, Qwen2-7B, and Qwen2.5-7B respectively achieve recall of 17%, 29%, and 38%. The same trend can be seen with the gpt and gemma families of models (Table 1). It should be noted that the Llama-3 family is an exception to this trend; Llama-3-8B (28% Recall) performs 6 points better than its later iteration Llama-3.1-8B (21.8% Recall). On the VLM side, Phi models also display this trend of decreasing performance for newer iterations (Phi-3-vision has 41% Recall while Phi3.5-vision reaches only 27% Recall).
Table 2: Model Precision & Recall across different HCTs input formats. Confidence intervals given in Appendix D.7.
Model Nonsensical-CSV Regular-CSV HTML
Precision Recall Precision Recall Precision Recall
gpt-4o 0.378 0.264 0.476 0.407 0.559 0.491
Llama-3.3-70B-Instruct 0.216 0.226 0.38 0.407 0.456 0.497
gemma-3-27b-it 0.211 0.17 0.344 0.305 0.457 0.429
gemma-3-12b-it 0.218 0.136 0.417 0.31 0.483 0.399
Qwen2.5-7B-Instruct 0.257 0.183 0.339 0.285 0.431 0.38
Llama-3.1-8B-Instruct 0.062 0.051 0.155 0.139 0.274 0.241
Meta-Llama-3-8B-Instruct 0.207 0.149 0.243 0.193 0.322 0.28
Mathstral-7B-v0.1 0.16 0.13 0.2 0.188 0.271 0.276
Average 0.214 0.164 0.319 0.279 0.366 0.339
Table 3: Fine-tuned model performances on real HCTs. Confidence intervals given in Appendix D.7
Model Fine-tuning Set Precision Recall CC Score
No FT 0.304 $|$ 0.262 0.279 $|$ 0.234 0.183 $|$ 0.171
Llama-3-8B $|$ FanarC-9B Real 0.600 $|$ 0.638 0.617 $|$ 0.646 0.525 $|$ 0.552
Real + Synth. 0.841 $|$ 0.864 0.849 $|$ 0.867 0.706 $|$ 0.734
(RQ3) Does the input format of the HCT matter in text-modality?
For the text models, we tested the effect of representing the table in HTML format instead of CSV. We used 1,000 synthetic HCTs (15,000 QA pairs) using 8 text-modality LLMs of different sizes. The results are shown in Table 2. We can see that the HTML format shows an improvement of 6% compared to canonical Regular-CSV, which is intuitive since HTML is a more expressive format that can better retain the complex structure of the HCTs.
We also explore the importance of having semantically coherent and meaningful wording within the HCTs. We created Nonsensical-CSV versions of the Regular-CSV tables where we replace all the words like ”Meat”, ”2024”, or ”Gender” in the HCTs and their corresponding QA pairs, with nonsensical terms like ”AAjw”, ”CCab”, and ”DDnq”. Although the complexity of the questions and table structure remains unchanged, we see that the semantics drastically impact the performance, with a decrease of Recall by 12 points on Nonsensical-CSV compared to Regular-CSV. Even large models like gpt-4o cannot compensate for the lack of semantics as its Recall score drops by 14 points. This is in stark contrast to SQL-based query engines where semantics have no effect whatsoever. This emphasizes the importance of having clear and semantically coherent names and values in HCTs when using LLMs for QA tasks.
(RQ4) Does fine-tuning models for this task improve performance?
We fully fine-tuned the FanarC-9B and Meta-Llama-3-8B-Instruct models on two SFT sets: (i) Real HCTs which has 7,867 (80% of total) HCT-QA triplets on Real World HCTs only, and (ii) Real + Synthetic HCTs which has 7,867 triplets on Real World HCTs and 62,000 triplets on Synthetic HCTs (total 69,867). The test set contains 992 triplets on Real World HCTs only. We ensure no table in the test set is seen during training (Details in Appendix D.6). The results are shown in Table 3. Fine-tuning drastically increases the performance of both models. Llama-3-8B Recall increases by 34 points when fine-tuned on Real HCTs only, and by 57 points when fine-tuned on Real + Synthetic HCTs compared to its off-the-shelf (No FT) Recall. FanarC-9B Recall increases by 41 points and 63 points, respectively, for each fine-tuning set. These results showcase the potential of models to learn HCT-QA tasks given enough training data. Furthermore, it reflects the importance of our synthetic data generator since including them more than tripled all scores on Real World HCTs.
Table 4: Effect of table properties on Recall. Only properties with significant effect ($p<0.05$) are shown
Table Properties Coef. (p-value) Table Properties Coef. (p-value)
Symmetric Row Nesting -0.103 (0.016) Global Column Aggreg. -0.035 (0.028)
Asymmetric Row Nesting -0.142 (0.001) Implicit Col. Aggreg. 0.081 (0.001)
Balanced Col. Nesting -0.270 (0.022) Implicit Row Aggreg. -0.065 (0.007)
Unbalanced Col. Nesting -0.315 (0.009) Row Group Label -0.069 (0.000)
(RQ5) How do HCT properties affect model performance? Table 4 shows the regression coefficient of various HCT properties against Recall (Properties are detailed in Appendix A). We can see that the presence of most of these properties, except Implicit Column Aggregation Terms, decreases the Recall on average. We also see a clear trend in models being able to better understand symmetrical and balanced nesting patterns compared to asymmetrical and unbalanced ones: Symmetric Row Nesting decreases Recall by 10% whereas Asymmetric Row Nesting decreases it by 14%. The same trend is seen when comparing Balanced Column Nesting with Unbalanced Column Nesting with the former having a decrease of 27% on Recall against 31.5% for the latter. Column and row nesting increases the HCT difficulty more than any type of aggregation, having the largest negative impact on Recall. Regression analysis on the full set of properties is given in Appendix D.2.
5 Limitations
The real-world HCTs were collected through various sources, but we could not guarantee a priori a uniform coverage across all layout properties (Section 3.1). However, we controlled the diversity and coverage of the natural language questions using predefined templates to guide their generation. These properties allowed us to characterize specific limitations of the evaluated models (Section 4). The synthetic HCT and QA generators (Section 3.3) allowed scaling up the size of the benchmark datasets at the cost of a lower diversity of the questions and layout structures compared to real-world data. Still, our fine-tuning experiments demonstrate the benefits of the synthetic datasets to improve LLM performances on real HCT-QA test data (Section 4).
6 Conclusion and future work
In this paper, we introduced a benchmark for querying human-centric tables found in PDFs, web pages, and similar documents. Our dataset includes a rich set of real-world and synthetically generated HCTs in addition to thousands of QA pairs, providing a rich resource to evaluate HCTs question-answering methods. Our results show that larger, closed-weight models (e.g., gpt-o1-mini) perform best overall, but struggle with complex queries—especially those involving HCTs with nested structures and queries that require aggregations. Performance improves with increased model size, instruction tuning, and more expressive HTML input formats. Notably, VLMs show promise by effectively processing HCT images. Still, the best LLM (gpt-o1-mini) tops at 69.1% precision and 68.5% recall, while the best VLM (gpt-4o-vision) tops at 48.7% and 62.3%, respectively, letting ample room for improvement. Our preliminary fine-tuning experiments show that fine-tuning Llama-3-8B and FanarC-9B using real-world and synthetic data more than triples their performance compared to no fine-tuning, demonstrating the potential of this approach and the benefit of our synthetic dataset.
Several research avenues could be explored. First, using our HCT-QA benchmark datasets to pre-train or fine-tune foundation models is an important endeavor to enhance their capacities to process HCTs. Our fine-tuning experiments show very promising results in this direction. Second, our work opens the way to address more complex QA tasks on multiple related HCTs spread within and across PDFs and other documents in a data lake. Finally, our synthetic generator could be used for the quantitative evaluation of NL-to-SQL (liu2024surveynl2sqllargelanguage, ), table transformation (Abedjan2016DataXFormerAR_46, ; AutoTables_57, ), and table extraction (Smock2021PubTables1MTC_15, ; tabnet-paper, ) techniques targeting HCTs.
APPENDIX
Appendix A HCT properties
A.1 HCT Property definitions and examples
Table 5 presents some summary statistics about the HCTs in our dataset. Definitions of HCT specific table properties are given in Table 6 and expanded upon with examples in this section.
Table 5: Summary statistics of the proposed benchmark datasets
QNPC Pak. Census US Census ArXiv SYNTHETIC
Average No. Rows 16 20 40 12 11
Average No. Columns 9 7 10 8 14
Average No. Cells 145 146 437 102 151
Average Table Word Count 201 187 575 138 153
Total Number of Tables 958 128 219 575 4,679
Total Number of QA Pairs 7,128 979 387 1,341 67,747
SIGNIFICANT TABLE PROPERTIES FREQUENCY (%)
Standard Relational Table 2.3% 7.0% 6.0% 18.6% 0%
Symmetric Row Nesting 7.4% 0.8% 0.0% 16.2% 100%
Asymmetric Row Nesting 3.2% 0.0% 0.4% 12.0% 0%
Balanced Multi Level Column 60.0% 66.4% 43.0% 59.1% 100%
Unbalanced Multi Level Column 0.8% 3.1% 22.6% 0.8% 0%
Implicit Column Aggregation Terms 6.8% 3.1% 2.6% 0.8% 0%
Implicit Row Aggregation Terms 10.1% 8.6% 43.0% 1.5% 0%
Global Column Aggregation 54.6% 13.3% 4.7% 3.1% 47%
Row Group Label 6.4% 36.7% 35.7% 8.0% 0%
*The percentages represent the occurrence of each table property across datasets.
*This subset of table properties has statistically significant effect on model performance.
Table 6: Definitions of HCT (table) properties.
Property Name Definition
Column Nesting When headers/column names are spread across more than one level (row).
Balanced Column Nesting When the depth of each branch within the nesting is the same throughout the table.
Unbalanced Column Nesting When the depth of each branch within the nesting is not the same.
Symmetric Column Nesting When the child column names are repeated across the table uniformly in columns where nesting is present.
Asymmetric Column Nesting When the child column names are different across the table in columns where nesting is present.
Row Nesting When a cell spans multiple rows but is displayed as one merged cell.
Balanced Row Nesting When the depth of each row branch is the same.
Unbalanced Row Nesting When the depth of each row branch is not the same.
Symmetric Row Nesting When the child cell values are repeated across the table uniformly in rows where nesting is present.
Asymmetric Row Nesting When the child cell values are different across the table in rows where nesting is present.
Column Aggregation Columns that aggregate all or a subset of other columns
Global Column Aggregation A column that aggregates all other numerical columns in the table (excluding local-group aggregation columns).
Local Column-Group Aggregation A column that aggregates a subset of numerical columns in the table.
Explicit Column Aggregation Terms An aggregation column where the column name clearly states it’s an aggregation (e.g., total, sum, percentage).
Implicit Column Aggregation Terms An aggregation column where the column name is ambiguous about whether it’s an aggregation.
Row Aggregation Rows that aggregate all or a subset of other rows
Global Row Aggregation A row that aggregates all other numerical rows in the table (excluding local-group aggregation rows).
Local Row-Group Aggregation A row that aggregates a subset of numerical rows in the table.
Explicit Row Aggregation Terms A row where some textual cell value clearly states it’s an aggregation (e.g., total, sum, percentage).
Implicit Row Aggregation Terms A row where all textual cell values are ambiguous about whether it’s an aggregation.
Row Group Label A row that acts as a heading/title for either all or a subset of rows below or above it.
Split Header Cell One cell that is split diagonally (visually) and contains two values.
Standard Relational Table Table displays none of the other HCT properties (nesting, aggregations, labels).
Balanced Column Nesting: Column nesting can be likened to a tree-like data structure. Similar to the definition of a balanced tree, we can see that each branch of column names (from the top most parent till each leaf) has the same depth. In case of all the table shown in the Figure 3, the depth is 2.
Figure 3: Example of Balanced Column Nesting
Unbalanced Column Nesting: In this the depth of all branches in the nested column tree do not have the same depth. In Figure 4, the first table has ”Population characteristics, 2003” as the top most parent, and has branches with depth of 2; the second table has the parent ”Cultivable Land” with branches of depth 2 and 3 as well.
Figure 4: Example of Unbalanced Column Nesting
Symmetric Column Nesting: Column nesting is labeled as symmetric when the children of all the cells in highest level of nesting that has more than 1 cell are all identical. In the case of the first table in Figure 5, the top two cells have identical child branches (Total, Injuries-Slight, Injuries-Sever, Death). Whereas the second table has the parent cells ”Urban” and ”Rural” both with identical children values (”Women” and ”Men”). We do not consider column values such as ”Month” (in the first table) and ”Area/Age Group” in the second table as they do not exhibit any nesting. The first table also shows an example of where column nesting is unbalanced but is still symmetrical.
Figure 5: Example of Symmetric Column Nesting
Asymmetric Column Nesting: Asymmetric column nesting is when the conditions stated above from symetric column nesting are not met. This entails parent cells having different children cells (branches). In the first table in Figure 6 we see the the parent cell ”Total” has the children (”G.Total”, ”%Female”, ”F”, ”%Male”, ”M”) whereas the other 2 parent cells (”Non-Qataris” and ”Qataris” have a different set of children (”%”, ”Total”, ”F”, ”M”). Similarly, in second table, we see that the parent cells ”Age (percent)” and ”One race (percent)” have different children cells, thus making this column nesting assymetric.
Figure 6: Example of Asymmetric Column Nesting
Balanced Row Nesting: Very similar to column nesting, balanced row nesting entails that all row nesting branches have the same depth (in this case depth is measured left-to-right rather than top-down as with columns). The 3 tables in Figure 7 show this property as the children values (the second column from the left in each table) are identical across their parent cells (left most column in each table).
Figure 7: Example of Balanced Row Nesting
Unbalanced Row Nesting: This is when the conditions for balanced row nesting are unmet and branches of nesting in rows have different depths. In the table shown in Figure 8 we see that the branches ”Orca - Sys. message - Synthetic” (depth of 3) and ”Orca - Original model” (depth of 2) have different depths, thus they are unbalanced.
Figure 8: Example of Unbalanced Row Nesting
Symmetric Row Nesting: Similar to symmetric column nesting, this property entails each parent cell in the first column (left most) that exhibits nesting properties having identical children cells (branches). We can see in the first table in Figure 9 that all the parent cells (”HDFS”, ”Abase”, ”Faas”, ”Alibaba”) have the same children cells (”RMSE”, ”MAE”, ”p-risk”). Similarly, all the parent cells in the second table have the same children cells (”No.of Hotels”, ”No.of Rooms”, ”No.of Beds”).
Figure 9: Example of Symmetric Row Nesting
Asymmetric Row Nesting: Antithetical to symetric row nesting, this property entails the parent cells having different children cells. The table on the left in Figure 10 shows different children values for each of the parent cells (”Weight-based” with 5 children cells, ”Activation-based” with 2 children cells, ”Passport-based” with 2 children cells that have different values from all the others). The same can be seen in the table on the right as the children values are different for both parent cells.
Figure 10: Example of Asymmetric Row Nesting
Global Column Aggregation: This property entails that there is a column in the table that meets one of the following criteria: (1) is the aggregate of all columns in the table including or excluding local column-group aggregate columns or (2) is the aggregate of all local column-group aggregate columns present in the table. In this case an aggregation can mean any operation or function (e.g sum, max, min, average, different, or any other equation applied consistently across the column that uses values from other columns). For the first table in Figure 11, the column ”Total - Total” is a global aggregate as it meets both the conditions: it is the sum of all columns excluding ”Total - NonQatari” and ”Total - Qatari” which are local-group aggregate columns, and it is the sum of the aforementioned local-group aggregate columns (which are the only such columns in the table). Similarly, the ”Total” column in the second table and the ”Avg” column in the third table perform the sum and arithmetic mean (average) operations on all the other columns in their tables respectively.
Figure 11: Example of Global Column Aggregation
Local Column-Group Aggregation: A column that aggregates only a subset of numerical columns (including other local column-group aggregation columns) within the table. The first table in Figure 12 has two such columns. ”Cultivable Land - Cultivated Land - Total” is the sum of all the other children columns under the parent cell ”Cultivable Land - Cultivated Land”. ”Cultivable Land - Total” is the sum of the aforementioned local-group aggregate and the ”Cultivable Land - UnCultivated Lands” column. Similarly, in second table the ”Tonne kilometres performed - Total” column sums all the children columns of the parent cell ”Tonne kilometres performed” and thus is a local column-group aggregate column.
Figure 12: Example of Local Column-Group Aggregation
Explicit Column Aggregation Terms: A column where the cell value clearly indicates aggregation (e.g., “Total”, ”Percentage”, ”Sum”). For nested columns, if even one of the cells in the branch has an explicit cell value then the aggregation is considered as explicit. Figure 13 shows three tables that show columns having values like ”Change”, ”Average”, and ”Total”.
Figure 13: Example of Explicit Column Aggregation Terms
Implicit Column Aggregation Terms: A column that aggregates data, but the column name does not make this explicit. Figure 14 illustrates this ambiguity as the column ”Yield” in the first table is the ratio of the other two columns. Although arguable that the name ”yield” is indicative of a aggregation, we do not consider such terms that require some degree of domain familiarity to be understood. Thus, it is considered as an implicit aggregation term. The second table presents a more definitive example of this property as the column ”Pakistan” is the summation of all the other columns (which represent different parts of Pakistan). Since it is difficult to understand that this is an aggregation without having prior knowledge about Pakistani geography, we consider this an implicit aggregation term.
Figure 14: Example of Implicit Column Aggregation Terms
Global Row Aggregation: A row that aggregates (using any function) all other rows in the table. Figure 15 presents examples of such aggregate rows that use the sum and arithmetic mean aggregations.
Figure 15: Example of Global Row Aggregation
Local Row-Group Aggregation: A row that aggregates a specific subset of rows (including other row-group aggregate rows) but not the entire table. Examples are shown in Figure 16 where the highlighted rows aggregate (sum in this case) subsets of rows. This property in particular has various visual manifestations: indentation in the first table, numbered rows in the second, explicit ”total” cell in third.
Figure 16: Example of Local Row-Group Aggregation
Explicit Row Aggregation Terms: A row where at least one cell explicitly indicates aggregation (e.g., “Total”, “Average”, “Sum”). This is clearly shown in Figure 17.
Figure 17: Example of Explicit Row Aggregation Terms
Implicit Row Aggregation Terms: A row that likely aggregates data, but none of its textual cells explicitly say so. Figure 18 illustrates this. As stated earlier, this property is often shown using numbered cells, indentation, domain specific terms (such as ”Gross Domestic Product”, font stylization, and other visual characteristics.
Figure 18: Example of Implicit Row Aggregation Terms
Row Group Label: A row that acts as a heading or title for a set of rows above or below it. See Figure 19 for examples. The first table has the row group labels ”Pakistan” and ”Punjab” respectively. The second table has ”C. Oil and Non-Oil”.
Figure 19: Example of Row Group Label
Split Header Cell: A single header cell split diagonally and containing two distinct values. This visual property is shown in Figure 20. In the first table the term ”Sector” describes the column names and ”Occupation” the values in the column below it. Similarly, in the second table, ”Year” describes the other column names, and ”Type of License” is the column name for that particular column.
Figure 20: Example of Split Header Cell
Standard Relational Table: A table that does not exhibit any of the defined HCT properties — no nesting, no aggregation, no labels. Figure 21 shows examples of this simple table form.
Figure 21: Example of Standard Relational Table
A.2 Table Annotation Tool
Figure 22: A user-friendly tool for annotating/viewing HCTs and their properties.
We developed a user-friendly tool for annotating tables and setting their properties as illustrated in Figure 22. The left side visualizes one table at a time, while the right side lists the possible allowed properties.
As an example, the illustrated table in Figure 22 contains “Column Nesting”. And since the child column names (e.g., {“Grand Total”, “Females”, “Males”}) across all groups are at the same level (same depth from the top parent cell), it is considered a “Balanced Column Nesting”. Similarly, since all branches of the parents have identical structure and cell values, the column nesting is considered “Symmetric”. Similar rules apply to the “Row Nesting” property. For example, the table in Figure 22 does not containing any row nesting, thus that property is not selected. The table also contains column aggregation at both levels, local (e.g., “Total” under “Non-Qataris”, “Males” under “Grand Total”), and global (e.g., “Total” under “Grand Total”). At the row level, this table contains only global aggregation (the last row). We also keep track of whether the aggregation columns carry explicit aggregate-meaningful labels (e.g., SUM, TOTAL, AVERAGE, MIN, MAX, etc.) as in the table illustrated in Figure 22.
Certainly, given the compound complexity among the different properties, we could have captured the properties at a more coarse or grained granularity. Nevertheless, we opt for this level of details because it sufficiently captures the underlying complexity of the HCT objects, and would provide good insights on how these properties correlate to QA performance without being dragged into too many features which requires high dimensional analysis. the annotation tool was developed in-house and we plan to release it on GitHub [^1] https://github.com/shahmeer99/HCT-QA-Benchmark/ page soon.
A.3 Understanding the distribution of table properties
The Figure 23 and Figure 24 show the distribution of the top-$5$ main and detailed real table properties per dataset. The property distribution is very different between datasets. For instance, column and row nesting are present in QNPC and ArXiv datasets (Figure 23(a) and (d)) but no row nesting appears in US C. or Pak C. datasets. Column and row aggregation appear in all datasets except in ArXiv with only a small proportion of row aggregation. QNPC datasets have the most various and complex combinations of main properties, while ArXiv have the least various.
Regarding detailed table properties in Figure 24, QNPC and US C. show quite complex and various combinations compared to Pak. C and ArXiv. While balanced column nesting is a typical property for all the tables, it is the major one in Pak. C and ArXiv while it is more balanced in the other two datasets.
Refer to caption Refer to caption
(a) QNPC (b) Pak. C
Refer to caption Refer to caption
(c) US C. (d) ArXiv
Figure 23: Distribution of main table properties for each real dataset.
Refer to caption Refer to caption
(a) QNPC (b) Pak. C
Refer to caption Refer to caption
(c) US C. (d) ArXiv
Figure 24: Distribution of detailed table properties for each real dataset.
Appendix B Question properties
B.1 Query Properties Definitions
Definitions of our query properties are in Table 7.
Table 7: Definitions of query-related properties used in question analysis.
Property Name Definition
Row Filter (RF) Whether the query filters rows based on some condition.
RF Condition Type Lookup The condition filters based on exact matches to cell values.
RF Condition Type Expression The condition uses numeric or logical expressions (e.g., >>, <<).
RF Involved Columns Single Only one column is involved in the filtering condition.
RF Involved Columns Multiple More than one column is involved in filtering.
RF Max Depth Of Involved Columns Maximum header depth of columns involved in filtering.
RF Retained Rows Single Only one row satisfies the filtering condition.
RF Retained Rows Multiple More than one row satisfies the filtering condition.
RF Num Of Conditions Number of conditions used to filter rows.
Returned Columns (RC) Whether the query returns specific columns.
RC Project On Plain Columns are returned without applying expressions.
RC Project On Expression Columns are returned with computed expressions.
RC Max Depth Of Involved Columns Maximum header depth among returned columns.
RC Expression In Table Present Returned expression is already present within the table.
RC Expression In Table Not Present Expression used in the result does not directly appear in the table.
RC Num Of Output Columns Number of columns returned in the query answer.
Yes/No Whether the question expects a yes/no answer.
Yes/No Scope ’Single’ if the yes/no judgment is based on one row or column, ’Multiple’ if it depends on multiple rows/columns.
Aggregation Whether the query includes any aggregation operation.
Aggregation On Plain Aggregation is applied on column values as they are.
Aggregation On Expression Aggregation is applied on computed expressions.
Aggregation Details Specifies the type of aggregation (e.g., min, max, count, sum, average).
Aggregation Grouping Local Aggregation is grouped by some column values.
Aggregation Grouping Global Aggregation is done over all table rows (no grouping).
Aggregation In Table Present The aggregated result already exists within the table.
Aggregation In Table Not Present The aggregation result is not directly present in the table.
Aggregation Num Of Aggregations Number of different aggregations needed to answer the query.
Rank Whether the question involves ranking or ordering.
Rank On ’Plain’ if ranking is performed directly on a column, else ’Expression’ if ranking is based on computed values or expressions
B.2 Human Verification Process
To ensure the accuracy and clarity of model-generated queries, we implemented a rigorous human verification and correction process. This process was designed to both evaluate the quality of the generated queries and to produce a high-quality dataset of verified query-answer pairs.
B.2.1 Annotation Protocol
Each model-generated query was evaluated by two independent annotators. Annotators were presented with the query, the associated table, and the model-provided answer. They were asked to complete two tasks:
  1. (i)
    Query Quality Rating: Annotators rated the query as one of the following:
    • Good: The query was well-formed, unambiguous, and meaningful with respect to the table.
    • Ambiguous: The query could be interpreted in multiple ways or was under-specified.
    • Nonsensical: The query did not make logical sense given the table content or was completely unrelated.
  2. (ii)
    Answer Correction: Annotators were instructed to check if the provided answer matched the correct answer, based solely on the table. If the answer was incorrect, they provided the corrected version.
If annotators disagreed on the quality label or the answer correction, a third annotator adjudicated the final decision.
B.2.2 Annotator Quality Control
To validate the consistency and attentiveness of annotators, we included a set of human-generated queries whose correctness was manually verified beforehand. Annotators who incorrectly marked more than one of these as ”nonsensical” or flagged correct answers as incorrect were removed from the pool. This vetting procedure ensured a reliable annotation process.
B.2.3 Query Outcome Statistics
After annotation:
  • 4% of model-generated queries were marked as nonsensical.
  • 24% were marked as ambiguous and subsequently rewritten.
  • 72% were rated as good and used as-is.
  • Approximately 50% of the model-generated queries required answer correction.
B.2.4 Examples of Query Types
Good Query:
What was the total number of passengers in 2019?
(Clearly refers to a valid column and year in the table; unambiguous and directly answerable)
Ambiguous Query:
How many were there in 2020?
(Unclear what ”many” refers to—passengers, flights, delays? Multiple possible interpretations)
Rewritten Version of Ambiguous Query:
How many total passengers were recorded in 2020?
Nonsensical Query:
what was the year for passengers and plane and its year ?
(Combines unrelated concepts not present in the table; not meaningful)
B.3 Query Generation and Annotation Query Tool
Figure 25: A user-friendly tool for query generation and their properties.
We developed a tool for QA generation shown in Figure 25. The left side displays the HCT table, while the right side enables the annotator to create a question and identify its properties. As shown in the figure, we capture the “Row Filter” property, which corresponds to the WHERE Clause in the SQL statement, where we capture the condition type (“Lookup” or “Expression”), the number of columns involved in the conditions, if the table involves multi-level column names, we capture the max level depth in the conditions, and the number of retained rows after all conditions are applied. The “Returned Columns” property corresponds to the SELECT clause in the SQL statement, where we capture the number of columns produced in the output, whether or not expressions are involved, and the max level depth among the output columns.
More complex properties involve “Aggregation”, “Rank”, and “Yes/No”, where we capture more details for each property as illustrated in Figure 25. For example, for “Aggregation”, we capture whether it is global or local, the number of involved aggregations, and the type of the aggregation. On the other hand, for the “Yes/No” query, which could involve many of the other properties (e.g., filtering, aggregation, and rank), we additionally capture whether the final Boolean expression is evaluated on a “Single” row or across “multiple” rows.
B.4 Understanding the distribution of query properties
The Figure 26 and Figure 27 show the distribution of the top-$5$ main and detailed real question properties per dataset.
The main question property distributions in Figure 26 are quite similar across datasets, with similar proportions of Returned.Columns, Row.Filter, and Aggregation properties. The proportion of the Returned.Columns and Row.Filter combination is more prominent in QNPC and Pak. C than in the other two datasets.
Regarding the detailed question properties in Figure 27, the distribution of top-$5$ properties is slightly decreasing for all datasets, while there are a few but frequent complex patterns in all datasets, and the remaining patterns are rare.
Refer to caption Refer to caption
(a) QNPC (b) Pak. C
Refer to caption Refer to caption
(c) US Census. (d) ArXiv
Figure 26: Distribution of main question properties for each real dataset.
Refer to caption Refer to caption
(a) QNPC (b) Pak. C
Refer to caption Refer to caption
(c) US Census. (d) ArXiv
Figure 27: Distribution of detailed question properties for each real dataset.
Appendix C Synthetic Data Generation
Refer to caption Refer to caption
(a) (b)
Figure 28: Overview of synthetic tables generation. (a) We generated $4,679$ unique synthetic HCTs from 77 semantics with up to $15$ questions each, leading to $67,747$ benchmark data triplets $(T_{HCT},Q_{NL},A)$. (b) These triplets derive from ground-truth data triplets $(T_{REL},Q_{SQL},A)$: Tables $T_{HCT}$ are obtained by pivoting relational tables $T_{REL}$; Natural language questions $Q_{NL}$ are obtained from templates of SQL queries $Q_{SQL}$; and the ground-truth answer AA is obtained by running $Q_{SQL}$ on $T_{REL}$.
Figure 29: Details of synthetic tables generation. User-defined semantic data, table and question templates are used to generate instances of pairs $(T_{REL},T_{HCT})$ of relational tables and pivoted human-centric variants as well as corresponding pairs $(Q_{SQL},Q_{NL})$ of templated SQL queries and NL questions. All the values (dashed color frames) of columns A, B, C, and D of the relational table are factorized as row and column headers in the HCTs (HCT1, HCT2). In contrast, the whole value column E (blue frame) is reproduced in the core area of the HCT (HCT3). In this instance, the ground-truth answer $A$ to the question "What is the amount of Export of Milk in 2017?" ($Q_{NL}$) related to $T_{HCT}$ is "57" (Dashed blue frame). This answer AA is obtained by running the SQL query $Q_{SQL}$ paired with $Q_{NL}$ onto $T_{REL}$. See Section C.1 for more details.
The statistics of the semantics of the synthetic tables and questions we generated are listed in Figure 28a. A summary view of the synthetic generation process is displayed in Figure 28b.
In this section, we explain the synthetic table and QA generation process in Section C.1. Then we detail the configuration templates for semantics (Section C.2), tables (Section C.3), and questions (Section C.4).
C.1 Generating process
The first step of the process is to manually define three different sets of templates (Symbols refer to Figure 29):
Semantic data
are textual items used as attributes and values to form realistic relational tables (e.g., ”Year”: ”2015”, ”2016”, ”2017”; ”Food category”: ”Beverage”, ”Dairy”, ”Cereals”;…). Details of this data structure are given in Section C.2.
Table templates
group several of the semantic data to form a family of pairs ($T_{HCT},T_{REL}$) of HCTs and corresponding relational tables. For instance, to generate a ”Food import-export” HCT, we first indicate the use of semantic data ”Year”, ”Import-Export”, food ”Category” and ”Item” (Columns A, B, C, D of $T_{REL}$, HCT1 and HCT2 headers in $T_{HCT}$), and then specify the number and complexity of the tables to generate in terms of (i) the range of numerical values to add in the core of the table (HCT3, ”Value” column E of $T_{REL}$), (ii) the depth of the column and row headers (HCT1, HCT2) and values therein, (iii) the presence and type of aggregation in $T_{HCT}$ (”Average” in HCT1 and HCT2) and (iv) other visual aspects (e.g., indented headers). Several format parameter values (ii-iv) are randomly generated from the same relational table instance $T_{REL}$ to vary the appearance of the HCTs without changing its content. Details of these table templates are given in Section C.3.
NL++SQL templates
are natural language question templates paired with SQL query templates, both related to a specific table template. There are $15$ different templates with increasing levels of complexity. For instance, the template:
What is the amount of_\$Import_Export of__\$Item__\$Category in__\$Quarter==of==\$Year? will generate the following natural language question based on the items selected from the current table instance:
What is the amount of Export of Milk in 2017?” that will be used as a prompt for the evaluated LLM along with the $T_{HCT}$ instance. The corresponding SQL query is:
$Q_{SQL}$: "SELECT Value FROM DBdata WHERE((Import_Export = ’Export’ AND Year = ’2017’))
AND ((Category = ’Dairy’ AND Item = ’Milk’))"
and will be run on the $T_{REL}$ instance paired with $T_{HCT}$ to get the ground-truth answer A=57A="57".
Details of these question templates are given in Appendix C.4.
Afterward, referring to Figure 29, we automatically process these template files to generate a large number of HCT and question instances: table templates use (A) semantic data to generate table instances (B) varying randomly or systematically the template parameters to specify the semantic data (C) to be used in rows (HCT1) or columns (HCT2), the aggregates, and the values (HCT3) of that table instance. From each table instance, we generate a relational table $T_{REL}$~(D) and one of its randomized HCT representations $T_{HCT}$ (E) using the pivottabler [9] R package. We also use its related set of question templates (F) to generate both SQL queries $Q_{SQL}$ and their natural language templated versions $Q_{NL}$ (G). We run the SQL query on the relational table instance $T_{REL}$ (H) to get the ground-truth answer $A$ (I), valid by design for the natural language question. There are $15$ question templates for each table template, leading to up to $15$ types of questions for each HCT. The generated table values (Column E ”Value” in $T_{REL}$, HCT3 in $T_{HCT}$) have no duplicate to avoid ambiguities when evaluating the matching between the ground-truth and LLM-predicted answers. Some semantic data are hierarchical, such as food categories and items (HCT2, columns C,D); we ensure they cannot be split across row and column headers of the HCT.
Here are the $15$ types of natural language questions $Q_{NL}$ generated for the table $T_{HCT}$ in Figure 29 based on templates (”,” and ”;” in answers, are column and row separators respectively):
  • $\mathbf{Q_{NL}1}$: one column, one row selection: ”What is the amount of Export of Milk in 2017?” $\rightarrow$ ”57”
  • $\mathbf{Q_{NL}2}$: one column, many rows selection: ”What is the amount of Export of Butter, Coffee, Milk, or Tea in 2017?” $\rightarrow$ ”57; 146 ; 980; 749”
  • $\mathbf{Q_{NL}3}$: many columns, one row selection: ”What is the amount of Export of Tea in 2017 or 2018?” $\rightarrow$ ”749; 622”
  • $\mathbf{Q_{NL}4}$: $Q_{NL}3$ + expression not in the table: ”What are the total and minimum amounts of Export or Import of Milk in 2017 or of Export of Milk in 2018?” $\rightarrow$ ”1442,57”
  • $\mathbf{Q_{NL}5}$ (opt): $Q_{NL}4$ + expression from the table: (only if an aggregate exists in the HCT) ”What is the average amount of Export or Import of Coffee?” $\rightarrow$ ”533.75”
  • $\mathbf{Q_{NL}6}$: many rows, many columns selection: ”What is the amount of Export or Import of Butter, Cream, or Tea in 2017 or of Export of Butter, Cream, or Tea in 2018?” $\rightarrow$ ”444; 146; 443; 736; 44; 821; 751; 749; 622”
  • $\mathbf{Q_{NL}7}$: $Q_{NL}2$ + aggregation not in the table: ”What are the total and minimum amounts of Import of Butter, Coffee, Cream, Milk, or Tea in 2017?” $\rightarrow$ ”2996,266”
  • $\mathbf{Q_{NL}8}$: $Q_{NL}6$ + grouping per column + reporting: ”What are the minimum and maximum amounts for each Import-Export and Year, of Export or Import of Butter, Coffee, Cream, Milk, or Tea in 2017 or 2018? Please, report the corresponding Import-Export and Year.” $\rightarrow$ ”Export,2017,44,980; Export,2018,121,821; Import,2017,266,799; Import,2018,179,879”
  • $\mathbf{Q_{NL}9}$: $Q_{NL}2$ + grouping per row: ”What is the minimum amount for each Category, of Import of Beverage or Dairy in 2018?” $\rightarrow$ ”673;179”
  • $\mathbf{Q_{NL}10}$: $Q_{NL}6$ + grouping per rows + reporting: ”What is the minimum amount for each Category, of Import of Beverage or Dairy in 2018? Please, report the corresponding Category.” $\rightarrow$ ”Beverage,673; Dairy,179”
  • $\mathbf{Q_{NL}11}$: $Q_{NL}6$ + grouping per rows and columns + reporting: ”What is the minimum amount for each Category, Import-Export, and Year, of Import of Beverage or Dairy in 2017 or of Export or Import of Beverage or Dairy in 2018? Please, report the corresponding Category, Import-Export, and Year.” $\rightarrow$ ”Beverage,Export,2018,121; Beverage,Import,2017,266; Beverage,Import,2018,673; Dairy,Export,2018,443; Dairy,Import,2017,444; Dairy,Import,2018,179”
  • $\mathbf{Q_{NL}12}$: $Q_{NL}2$ + top-K + row filter: ”What are the bottom 5 amounts of Export of Beverage or Dairy in 2018?” $\rightarrow$ ”121; 443; 586; 622; 821”
  • $\mathbf{Q_{NL}13}$: $Q_{NL}2$ + row filter + ordering: ”What are the amounts ordered by increasing values of Import of Beverage or Dairy in 2017?” $\rightarrow$ ”266; 444; 736; 751; 799”
  • $\mathbf{Q_{NL}14}$: $Q_{NL}2$ + operation on column: ”What are the Category and Item for which the amount of Import in 2018 is greater than 513.3?” $\rightarrow$ ”Dairy,Milk; Beverage,Coffee; Beverage,Tea”
  • $\mathbf{Q_{NL}15}$: $Q_{NL}14$ with condition + reporting: ”What is the amount of Export in 2018 of Category and of Item for which the amount of Import in 2018 is greater than 513.3? Please, report the corresponding Category, Item, and amount of import-export.” $\rightarrow$ ”Dairy,Milk,586; Beverage,Coffee,121; Beverage,Tea,622”
C.2 Semantic templates
Figure 30: Semantics for generating synthetic tables and questions. Examples of $T_{REL}$ (a) and a corresponding $T_{HCT}$ (b) related to the amount of food import-export in some place. (c) JSON format to describe Food Import-Export semantic terms used to generate these tables (see Section C.3) and the related questions (see Section C.4). Not all semantics terms are used in $T_{REL}$ and $T_{HCT}$.
In this section, all terms refer to Figure 30.
Our synthetic data relies on semantic terms naming the dataset’s attributes and values. A synthetic $T_{HCT}$ (Figure 30b) is obtained by pivoting a relational table $T_{REL}$ (Figure 30a) composed of at least two nominal attributes and a single numerical attribute always named Value.
There are two types of nominal attributes:
  • Independent attributes like Import-Export, Years, or Quarters can be organized as row or column headers at any nesting level in $T_{HCT}$.
  • Hierarchical attributes like food Category (e.g.Meat, Cereals, Dairy, Beverage) and food Item (e.g.Cream, Soda, Milk) are dependent on each other. If both attributes appear in the HCT, the food Item must be displayed at a lower level than the food Category, i.e. closer to the inner part of the HCT, and all must appear together either in row headers or column headers.
Semantics are encoded in a JSON object (Figure 30c) with two main keys: data and values.
  • data refers to a list of attributes (e.g.ImportExport, Years, Quarters) or groups of attributes (e.g.Food), each described with three keys:
    • code is the primary key used in table and question templates to identify the independent attributes (e.g.ImportExport or Years) and the hierarchical groups of attributes (e.g.Food).
    • names is the list of attribute names displayed in $T_{REL}$ column headers but typically hidden (implicit) in $T_{HCT}$ and $Q_{NL}$. It is a list with a single name for independent attributes, or with as many names as levels for hierarchical groups of attributes (e.g.Food), ordered from top (outer) to bottom (inner) levels (e.g.Category, Item).
    • values is the list of values (e.g.2015, 2016... for the attribute Year). It can also be the hierarchy of values of a group of attributes. For instance, in Food group, Meat is a value of the attribute Category, and Beef, Chicken…, are values of the attribute Item, whose parents are Meat in the Food hierarchy. Attribute values are always displayed as row or column headers in $T_{HCT}$.
  • values refers to the numerical values to be generated in both the core part of $T_{HCT}$ and the Value column of $T_{REL}$. This object contains shortcut keys (e.g.realUnit or intSigned) to generate specific ranges and types of numbers when called in table templates. The value associated with such shortcuts must be a list of two numbers $[m,M]$ representing the minimum $m$ and maximum $M$ possible values. If any of these min and max is a decimal number, the generated value will be a real number with two decimal digits drawn uniformly in this interval. If both min and max are integers, the generated value will be an integer drawn uniformly in this interval. The generator will ensure no replica in the generated values, reals or integers, possibly overflowing the max value if asked to generate more numeric values than the range allows.
Customization of semantic data
All terms in semantic data are customizable except the JSON keys data, code, names, and values, following the rules above. The number of decimals of generated real numbers is specified by NUM_DECIMAL_DIGITS_REAL_FORMAT in config.R. These semantic data are stored in the file PARAM_semantics.json.
C.3 Table templates
Figure 31: Example of a synthetic table template JSON file described in Section C.3, used to generate sets of Food Import/Export HCT and relational tables. This template is stored in the file PARAM_tableTemplates.json
The generic table templates described in the file PARAM_tableTemplate.json are used to generate a (large) set of $T_{REL}$ and $T_{HCT}$ tables from the semantic data (See Section C.2). All terms below refer to Figure 31.
This template contains the following keys:
  • replica is an integer setting how many random variants of each initial table template from the list tables will be generated.
  • shuffle is a list of column and row shuffling commands that will be applied to all the initial table templates listed in tables. It is used to vary the structure of the initial table. none is the default; row mixes only row attributes; cols mixes only column attributes; rowscols mixes rows and columns, keeping them as rows or columns, respectively; all mixes across rows and columns (some columns can become rows and vice-versa, except for hierarchical attributes which stay grouped in rows or columns). The shuffling process is a random permutation of the rows or columns to be mixed.
  • col_row_levels is a list of authorized depths for nested rows and columns. It must be a list of strings ”dcdc_drdr” indicating the depth of the column header ($dc$), and of the row header ($dr$), with $dc$ and $dr$ in $\{1,2,3\}$.
  • col_row_name_pos is a list of authorized displays for row and column attribute names. It must be a list of strings ”C_R” indicating where the column (C) or row (R) attribute shall be displayed. R and C can be none to hide the attribute name, left to show it on the left side of the header, or top to show it at the top of the header. (Place-holder for a future version, ignored by the generator).
  • col_row_agg_pos is a list of authorized positions for aggregate rows or columns. It must be a list of strings ”C_R” indicating on which side of the header the aggregate column (C) or row (R) shall be displayed: R can be set to top or bottom, C can be set to left or right, and both can be set to none (Default) to hide the corresponding aggregate.
  • row_format can be set to new to use one column for each level of row header, or indent to create an empty row for all level $L$ and indent all rows at level $L+1$. The header of the empty row is the attribute value for that row. An aggregate can fill this row if col_row_agg_pos is set to *_top.
  • tables is a list of table content specifications (values, headers, and aggregates) with coherent semantics. Each initial table template in the list has the following keys:
    • valueName is the title displayed at the top of the $T_{HCT}$ table (e.g., ”Food import-export in tons”)
    • values is either a shortcut name referring to one of the keys (e.g.realUnit) stored in values in the semantic data (see Figure 30), or directly an interval of the form $[m,M]$ following the description of the main key values in Section C.2
[…]
    • rowCodes is a list of one or more of the data code keys (e.g.Food or Quarters) listed in the semantic data. The generator will pick among these codes to generate the rows of $T_{HCT}$, and one part of the nominal attributes of $T_{REL}$.
    • rowSamples is a list containing as many intervals $[m,M]$ as there are codes listed in rowCodes, in the same order. Each interval specifies the minimum $m$ and maximum $M$ number $n$ of values to select among the list $L$ of values of the specified attribute. The generator first picks $n$ at random in $[m,M]$, then selects $n$ consecutive values in $L$ starting at a random position in $L$ but ensuring there are always $n$ values in the result if the list $L$ is long enough, or return $L$ itself otherwise. For instance, let’s consider an independent attribute Years with a list of values: $L=[2001,2002,2003,2004,2005,2006]$, if we set rowSamples to $[1,5]$, and the generator randomly picks $n=3$, it then picks a random index $i$ between $1$ and $4$, e.g.i=2i=2, to generate the final subset $[2002,2003,2004]$ of years to use in the table. For hierarchical attributes, the same process applies recursively to all levels of the hierarchy, starting from the top level. If the interval is set to $[0]$ instead, all the values of the corresponding independent attribute are generated, or all the values at all hierarchy levels are generated for hierarchical attributes. Note that the col_row_levels parameter limits the depth of row and column headers, so the attributes finally displayed in $T_{HCT}$.
    • colCodes is the same as rowCodes but for columns. The generator will pick among these codes to generate the columns of $T_{HCT}$, and the other part of the nominal attributes of $T_{REL}$.
    • colSamples is the same as rowSamples but for columns.
    • agg_name1 is the name of the aggregate to display in the table.
    • agg_fun1 is the aggregate function to use among sum (default), avg (average), min, max.
These generic table templates PARAM_tableTemplate.json are used with the semantic data PARAM_semantics.json to generate individual templates using the script 1_geneAllTablesJSONfromPatterns.R. The individual templates generated and described below are stored in the file PARAM_tablesToGenerate.json.
Customization of generic table templates
These generic table templates are customizable except for all the JSON keys and following the rules above.
Figure 32: Two examples of individual table templates (a,b), each with two random samples (1, 2) of $T_{REL}$ (a’, b’) and $T_{HCT}$ (a”, b”). Templates a and b differ by the depth of column headers (col_row_levels), row indentation (row_format), presence and position of aggregates (agg_pos1). The sampled $T_{HCT}$ (1, 2) of a given template (a or b), further differs by the randomized order of the attributes (shuffle), number and names of the attribute values (sample), core numerical values (values), and presence of table cell borders (random).
An individual table template is a JSON object (Figure 32) instantiating one combination of the parameters shuffle, col_row_levels, col_row_name_pos, col_row_agg_pos, and row_format common to all initial table templates, and set in the generic table template. It contains the following keys:
  • name is the identifier of the family of $T_{HCT}$ and $T_{REL}$ tables specified by this individual table template.
  • replica is identical to replica set in the generic table template.
  • agg_fun1 is identical to agg_fun1 set in the generic table template.
  • agg_name1 is identical to agg_name1 set in the generic table template.
  • values is identical to values set in the generic table template.
  • valueName is identical to valueName set in the generic table template.
  • shuffle is one of the shuffling commands listed in shuffle of the generic table template.
  • row_format is one of the row format commands listed in row_format of the generic table template.
  • columns and rows are lists of attribute codes obtained after filtering the first $dr$ row attributes and $dc$ column attributes and setting the aggregate position using the col_row_levels and agg_pos1 parameters from the parameter combination of this instance. sample are identical to colSamples and rowSamples of the corresponding attribute set in the generic table template.
Finally, replica instances of $T_{HCT}$ are generated using the script 2_geneHCTfromJSONtables.R which takes as input the files PARAM_semantics.json, PARAM_tableTemplate.json, and PARAM_tablesToGenerate.json. It returns HTML files named [name]_n_DB.html ($T_{REL}$) and [name]_n_HCT.html ($T_{HCT}$) with $n$ the replica number.
This process draws samples from each individual template, randomizing all the remaining parameters: the shuffle (stochastic) command is applied to rows and columns attributes, then the values of each attribute are sampled using the sample parameter, and the values in the core of the table are sampled from the values parameter. The absence or presence of cell borders in the $T_{HCT}$ layout is randomized for each final sample.
Two examples of individual templates with two of the $5$ sampled $T_{REL}$ and corresponding $T_{HCT}$ are shown in Figure 32.
Customization of individual table templates
There is no possible customization of these individual table templates automatically generated from the generic table templates.
C.4 Question templates
Figure 33: Synthetic question templates for the Food Import/Export (a) and Number of Students (b) table semantics.
Figure 33 shows the natural language question templates related to ”Food import-export in tons” (a) and ”Number of students” (b) semantics and tables, as a JSON object with the following keys:
  • tableName is identical to the valueName key in the generic table template (see Section C.3).
  • rowColCodes gather the rowCodes and colCodes lists of attribute codes used in rows and columns of the generic table template.
  • valueMeaning is a text characterizing a value in the table continuing the sentence: ”A value in this table represents the…” (e.g., ”amount of import-export”).
  • simplifyNested lists all the names group of hierarchical attributes used in the generic table template.
  • template_report is a sentence to report a specific list of attributes \$REPORTATTR selected for the query result in templates $8,10,11,15$.
  • template_N a list of templates for questions of type N (See the list of $15$ types in Section C.1).
    • A template is a natural language sentence with some place-holder variables to be filled by the generator depending on the table instance content and the type of question. All the templates must show all the possible attributes set in the generic table template, but the question generator will randomly pick the ones it will use, depending on the template type; the others will be absent.
    • We use \$attribute_name to specify variables that will be automatically replaced by the generator with values from that attribute present in the table instance (e.g.\$Category could be replaced by Meat, Dairy, Cereals, or Beverage). These attribute_name must appear in the rowColCodes list for independent attributes, and in the simplifyNested list for groups of hierarchical attributes.
    • All words in capital letters \$RESERVED and their connected words should not be modified, and must appear as-is in their respective template.
    • In templates $4,5,7,8,9,10,11$, an expression \$EXPR will be replaced by one of the aggregation names agg_name1 (min, max...) in the generic table template. Some question can ask several expressions at once, like the min and the max, calling for a plural version of the sentence. We use // to separate the singular and plural versions of a word that the generator will pick automatically depending on the number of expression terms generated. For instance, the template $7$: What is/are_the_\$EXPR_amount/amounts of_\$Import_Export of... can generate ”What is the minimum amount of Export of…” or ”What are the minimum and maximum amounts of Export of…”
    • In templates $8,9,10,11$, \$GROUPBY will be replaced by for each followed by a list of attribute names. For instance, the template $9$: What is/are_the_\$EXPR_amount/amounts \$GROUPBY of_\$Import_Export of__\$Item__\$Category in__\$Quarter==of==\$Year ? can generate the question ”What is the minimum amount for each Category, of Import of Beverage or Dairy in 2018?”
    • In template $12$, top/bottom_\$TOPK will be replaced by an integer and either top or bottom automatically selected, e.g.”What are the top 5 amounts of Import of…
    • In template $13$, the group \$ORDERBYDESC_decreasing/increasing_values must be used as-is and can generate the question ”What are the amounts ordered by increasing values of Import of Beverage or Dairy in 2017?” or ”What are the amounts ordered by decreasing values of…”
    • In template $14$, the word \$OPPATTR will be replaced by a list of attribute names, and the word \$OPANDVAL by either greater than or lower than and a certain value, leading to a question like: ”What are the Category and Item for which the amount of Import in 2018 is greater than 513.3?”
    • We use _ to link words like of, in to their attribute. If an attribute is not present in an instance of a question, all the words connected to it with _\_ will be absent too (e.g.of_\$Import_Export).
    • We use __ to link several attributes together. For instance, we use __\_\_ in of__\$Item__\$Category so if \$Item is absent, the of word stays connected to \$Category like in of Meat. This pattern can generate of Beef and of Meat, but will not generate of Beef Meat because these attributes are part of the same hierarchical group.
    • We use == to link several attributes separated by a word that only exist if both surrounding attributes are present. For instance: we can say in Q2 of 2022 or in 2022 or in Q2. Thus, we tell that of is only present if both \$Quarter and \$Year are present together, by coding in__\$Quarter==of==\$Year.
    • We use ((not_an_attribute)) to ensure the presence of a word which is not an attribute. For instance, we want to enable the generation of the number of English male students but also the number of students if the \$Nationality and \$Gender attributes are absent from the generated question. As students is not an attribute value, we use the template of_((students)) of__\$Nationality__\$Gender__students to enable this option. The generator will only produce the first part of_((students)) if the word students does not appear in the rest of the generated sentence (See Figure 33b).
    • The order of the attributes used in the template is always preserved.
    • Additional rules, a template shall never start with a \$ and a space must be set before the final question mark. Attribute names composed of several words separated by - should replace it by _.
All $14$ templates are repeated with variations for the $7$ semantics in the PARAM_NLquestionTemplates.json file, guiding the creation of the same templates for new semantics. Note that the question $Q_{NL}15$ is automatically generated from templates $13$ and $14$.
The script 3_geneSQLandNLQfromJSONtemplates.R reads the question template file PARAM_NLquestionTemplates.json and the other parameter files, to generate the natural language questions $Q_{NL}$ and corresponding SQL queries $Q_{SQL}$ for each question template as well as their ground truth answers $A$ from the tables generated from PARAM_tablesToGenerate.json.
Customization of question templates
These question templates are customizable except for all the JSON keys and following the rules above.
The Figure 34 shows the matching between the synthetic question types and the question properties described in Section B.
Figure 34: Correspondence between synthetic question templates (types 11 to $15$ in columns) and question properties (rows). Cells are green when the property is present in the template type. $n$ is a numeric value that can vary for a given type.
C.5 Synthetic generation code
The list of codes and the instructions for installing and running them are given in https://github.com/shahmeer99/HCT-QA-Benchmark/blob/main/synthetic_data_generator/README_SYNTHETIC_GENERATOR.md.
Appendix D Additional Experiments and Details
D.1 Model Details
Table 8 contains all the information about the models used in our experiments.
Table 8: Summary of Evaluated Language Models
Model Name Provider Size
TEXT-ONLY LLMs
Large Models – 100B+ (Closed)
gpt-o1-mini OpenAI 100B+
gpt-4o OpenAI 100B+
gpt-35 OpenAI 100B+
Medium Models – 27B-70B (Open)
Qwen2.5-72B-Instruct Alibaba 72B
Llama-3.1-70B-Instruct Meta 70B
Llama-3.3-70B-Instruct Meta 70B
gemma-2-27b-it Google 27B
gemma-3-27b-it Google 27B
Mixtral-8x7B-Instruct-v0.1 MistralAI 56B
Small Models – 3B-12B (Open)
Meta-Llama-3-8B-Instruct Meta 8B
Llama-3.1-8B-Instruct Meta 8B
Qwen1.5-7B-Chat Alibaba 7B
Qwen2-7B-Instruct Alibaba 7B
Qwen2.5-7B-Instruct Alibaba 7B
gemma-2-9b-it Google 9B
gemma-3-12b-it Google 12B
Mistral-7B-Instruct-v0.3 MistralAI 7B
Mathstral-7B-v0.1 MistralAI 7B
FanarC QCRI 9B
aya-23-8B CohereForAI 8B
Phi-3.5-mini-instruct Microsoft 3.8B
Model Name Provider Size
VISION-TEXT VLMs/MLMs
Large Models – 100B+ (Closed)
gpt-4o-vision OpenAI 100B+
Small Models – 3B-12B (Open)
Pixtral-12B-2409 MistralAI 12B
InternVL2-4B SenseTime 4B
InternVL2_5-8B-MPO SenseTime 8B
Phi-3-vision-128k-instruct Microsoft 4.2B
Phi-3.5-vision-instruct Microsoft 4.2B
Idefics3-8B-Llama3 HuggingFace/Inference.ai 8B
Qwen2-VL-7B-Instruct Alibaba 7B
llava-1.5-7b-hf LMSYS Org (LLaVA) 7B
llava-v1.6-mistral-7b-hf MistralAI (LLaVA) 7B
llava-v1.6-vicuna-7b-hf LMSYS Org (LLaVA) 7B
Molmo-7B-D-0924 ImitateML 7B
paligemma2-10b-ft-docci-448 Google 10B
D.2 Additional Trend Analysis & Observations
(RQ6) Does model verbosity affect the output correctness?
Our initial hypothesis was that if models simply take the kitchen-sink approach and output more values in their answers then they would have a higher probability of matching values in the ground truth and thereby inflate their recall scores. As shown in Figure 35, there is no clear correlation between answer length and recall. Most models, including the best performing ones, seem to be clustered around the middle on the verbosity scale. The average number of values in the ground-truth answers for all queries is  4, yet most models tend to provide between 2-3 values in their answers. This suggests that models may in fact be conservative in their answer verbosity as which is antithetical to the overly verbose nature of these generative models.
Figure 35: Recall Score vs. Verbosity for Text-Only Models on Real World HCTs
(RQ7) How do question properties affect model performance?
The synthetic queries we generated have 15 types including various types of selection, projection, expression, and aggregation. The model performance per template type is shown in Figure 36. As expected, models tend to do best on selection queries be it single-cell or multi-cell selection. This is clear as templates 1, 2, and 3 have the highest average recall across all models. Models tend to do better on queries involving expressions on columns (templates 5, 14) than queries involving aggregations (template 7). In fact, models struggle the most with template 7 which requires doing an actual aggregation (rather than retrieving an existing one from the HCT). Models like LLaMA-3.3-70B significantly outperform even models like GPT-4o and GPT-35 on this template. It is also evident that models clearly have their strengths and weaknesses in terms of templates. GPT-4o gets 58% on template 11, whereas other models like LLaMA and Mistral do not even cross 10%. However, this dominance completely disappears in Templates 4, 7, and 9 where the GPT-4o fails to score more than 8% (compared to other models that get up to 42%). There are also unexpected trends across model generations. Gemma-2-27B scores a 52% recall on template 14 which is only topped by GPT-4o and Qwen2.5 72B, both of which are relatively larger models. However, on the exact same template (14), Gemma-3-27B scores exactly half of its predecessor with a recall score of 26%. This shows that newer generations of models are not necessarily improving their capabilities on this task across the board.
Figure 36: Recall Score for Text-Only Models on all Synthetic HCTs per Question Template
(RQ8) How are model size and performance correlated?
It is clear that large closed-weight models like GPT are far superior to their smaller open-weight competitors. However, a fairer comparison to answer this question would be between the same model version but in different sizes. Figure 37 shows exactly that and we can see that larger models tend to do considerably better than their smaller versions. LLAMA-3.1 improved its recall score by nearly 25% with a 9x increase in parameter size. Similarly, Gemma-2 improves its recall score by 11% with a 3x parameter size increase. The improvement between Mistral-7B and Mixtral-8x7B is only 8% but this is most likely due to the fact that the latter is a Mixture-of-Experts model where not all experts (active parameters) are being used for this task. Overall, increase in parameter size does improve performance to a good degree. On average when we consider LLaMA3.1, Gemma-2, Gemma-3, and Qwen2.5 we see an approximate 4% increase in recall score per every 10B additional parameters.
Figure 37: Recall Score Across Size (Number of Parameters) within Model Families
(RQ9) In what cases do VLMs perform better than LLMs?
Recent advances in the image understanding and reasoning capabilities of VLMs make them a suitable candidate for the task at hand. VLMs also allow to skip the extraction and conversion to CSV process as they can directly consume the images of the tables from their source documents. On average, VLMs score 25.5% recall whereas text-only LLMs score 25% as shown in (Table 1). GPT4o when prompted using the text-modality achieves an average recall of 61.6%, compared to 62.3% when using the image-modality. This illustrates the progress vision models have made in terms of document understanding and reasoning tasks such as this one. We have established that HCTs inherently have visual characteristics that are crucial to understand the content in the table. A lot of those visual clues are lost when the HCT is converted to text format (regardless of which text format its represened in). VLMs are able to overcome this information loss by taking advantage of those visual cues. This is further shown by the top performing small sized VLM model, Pixtral-12B with a recall score of 45.9% which is nearly 8% better than the best performing text-only LLM of that size range (Qwen2.5-7B with recall 38%). Furthermore, Microsoft’s Phi-3-Vision-128k (image-modality), which is only a 4.2B parameter model outperforms every text-only LLM in the small size category (many of which are 2x in size). These results clearly highlight the potential of VLMs in visual-reasoning tasks such as HCT-QA. Another interesting observation is that VLMs struggle considerably more than LLMs with larger HCTs. GPT4o performs almost identically in both the text and vision modalities except in the US CENSUS dataset which on average has 437 cells per table (nearly 3x more compared to the other datasets, see Table 5). This trend can also be seen in the smaller VLMs as most top performing models drop in performance when dealing with US CENSUS HCTs.
Figure 38: LLMs vs. VLMs on all Real World HCTs - Recall Score
Regression analysis of table properties and mean recall score
(across all models) is shown in Table 9.
Table 9: Effect of table properties on Recall. Only properties with significant effect ($p<0.05$) are shown
Table Property Coef. (p-value)
Symmetric Row Nesting -0.103 (*0.016)
Asymmetric Row Nesting -0.142 (*0.001)
Symmetric Col. Nesting -0.095 (0.148)
Asymmetric Col. Nesting -0.127 (0.061)
Balanced Row Nesting -0.251 (0.073)
Balanced Col. Nesting -0.270 (*0.022)
Unbalanced Col. Nesting -0.315 (*0.009)
Global Column Aggreg. -0.035 (*0.028)
Implicit Col. Aggreg. 0.081 (*0.001)
Implicit Row Aggreg. -0.065 (*0.007)
Row Group Label -0.069 (*0.000)
D.3 Prompting Details & Experiments
For certain models (like Gemma-2) that do not accept a system prompt we add the system prompt as a prefix to the user prompt and send it as one prompt with the ”user” role.
D.3.1 Prompts
System Prompt for Inference You are a table question answering assistant. You are given a table and a question. Your task is to answer the question based on the information in the table. The table structure may be complex and not a standard relational table, so try to understand the structure of the table when answering the question. If the question cannot be answered using information from the table, return ‘No Answer‘. Do not provide any explanations, equations, code, or text explaining intermediate steps in figuring out the answer. Return only the final answer itself. If there are multiple values, separate each value with the ‘ || ‘ token.
User Prompt for Inference Answer the question given the input table. Do not provide any explanations, equations, code, or text explaining intermediate steps in figuring out the answer. Return only the final answer itself. # INPUTS: Table: ‘{table}‘ Question: ‘{quesiton}‘ # OUTPUT: Answer:
D.3.2 0-shot vs. 1-shot prompting
The majority of the models we used in our experiments can handle context lengths of 8,000+ tokens. This means that it is possible to do 1-shot prompting with a full table-question-answer triplet as the shot. This approach could help leverage in-context learning to improve the model performance. We experimented on a small set of 1,000 queries using Real World HCTs from all four data sources. Our 1-shot prompt consisted of a fixed example for all the questions in the test set. We avoided testing the models given an example question on the same HCT the test question was on as that reflects a very unrealistic scenario. The results are shown in Table 10. We can see that adding the 1-shot example does not significantly affect the recall score as for gemma-2-9b-it it only increases by 1.3% and for Llama-3.1-8B-Instruct it actually decreases by less than 1%. Upon manual inspection, we found that the models were able to follow the output format stated in the prompt (values split by ‘ || ‘ and no explanations) in the 0-shot scenario almost perfectly. As the 1-shot example was on a different table, it provided no ”learning” for the model to answer the test question it was being asked. Due to these results, our remaining experiments use 0-shot prompting.
Table 10: 0-shot vs 1-shot Prompting - Recall Score
Model 0-Shot Recall 1-Shot Recall
gemma-2-9b-it 0.36 0.373
Llama-3.1-8B-Instruct 0.247 0.239
D.4 Evaluation Metrics
We experimented using Meta-Llama-3-70B as judge model to gauge the correctness of a model’s answer. The prompt given to the judge model outlined a clear set of instructions: (1) minor spelling mistakes or typos in textual values can be ignored, (2) rounding numerical values is fine if done reasonably, (3) question must be answered fully in order to be ”correct”. We compared this model’s ”correct” labels to our CC Score (Complete Containment, which we defined as 1 when Recall Score == 1, else 0). The comparison is shown in Figure 39. It is evident that the model’s judgment and the CC score closely align with each other. Since running an LLM-as-a-Judge is expensive in terms of computational resources, we opt to use Recall, Precision, and CC Score as our evaluation metrics.
Figure 39: CC Score vs. LLM-as-a-Judge Score on Real World HCTs
D.5 VLMs vs LLMs
We can see from the results in Table 1 that all VLMs are sensitive to table size. HCTs arXiv and Pak have relatively lower number of cells and number of words on average and vision models tend to do the best on these tables. It is evident that all vision models struggle with US Census tables (which are the largest tables on average with good margin). Gpt-4o scores 39.6% recall on US Census which is less than majority of medium sized text-only models which are considerably smaller in size.
This is further clear by the fact the gpt-4o-vision which outperforms its text-only version (gpt-4o) in all 3 datasets (QNPC, Pak, arXiv) but only scores a recall of 39.6% for US Census which is 20% less than that of gpt-4o (text-only) at 59%. Some examples of cases where gpt-4o gets answers a question correctly in one modality but gets it wrong in the other are provided in Figure 40 and Figure 41.
Figure 40: gpt-4o-vision is correct but gpt-4o text-only is wrong
Figure 41: gpt-4o-vision is wrong but gpt-4o text-only is correct
D.6 Finetuning Details
We fine-tuned 2 models (Meta-Llama-3-8B-Instruct and FanarC-9B) using LLaMA Factory [^2] https://github.com/hiyouga/LLaMA-Factory using the following hyperparameters:
Finetuning Configuration Summary Finetuning Type: Full-parameter finetuning Max Sequence Length: 4096 tokens Precision: bfloat16 (bf16) Training Epochs: 1 Batch Size: 2 per device Gradient Accumulation: 2 steps Learning Rate: 5e-7 Learning Rate Schedule: Cosine with min LR = 5e-8
Details on the training and test data can be found in Table 11. The dataset itself, with the exact train + validation + test splits used in our finetuning experiments, can be found on HuggingFace [^3] https://huggingface.co/datasets/qcri-ai/HCTQA.
Table 11: Number of Questions and Tables by Data Source (Train vs. Test)
Split Num Questions Num Tables
arXiv US CENSUS PAK QNPC arXiv US CENSUS PAK QNPC
Train 1104 304 774 5685 468 171 101 764
Test 121 39 100 732 55 22 14 97
D.7 Confidence Intervals (Error Bars) for Main Experiments
In the appendix section we present the confidence intervals (CI @ 95%) for all the experiments presented in Section 4.
Table 12: Precision and Recall with 95% confidence intervals for results in Table 2
Model Format Precision [95% CI] Recall [95% CI]
gpt-4o Nonsensical-CSV 0.378 [0.370, 0.386] 0.264 [0.257, 0.270]
gpt-4o Regular-CSV 0.476 [0.468, 0.485] 0.407 [0.399, 0.415]
gpt-4o HTML 0.559 [0.550, 0.567] 0.491 [0.483, 0.499]
Llama-3.3-70B-Instruct Nonsensical-CSV 0.216 [0.210, 0.223] 0.226 [0.220, 0.233]
Llama-3.3-70B-Instruct Regular-CSV 0.380 [0.372, 0.387] 0.407 [0.399, 0.414]
Llama-3.3-70B-Instruct HTML 0.456 [0.448, 0.464] 0.497 [0.489, 0.505]
gemma-3-27b-it Nonsensical-CSV 0.211 [0.204, 0.218] 0.169 [0.164, 0.175]
gemma-3-27b-it Regular-CSV 0.344 [0.337, 0.352] 0.305 [0.298, 0.312]
gemma-3-27b-it HTML 0.457 [0.450, 0.465] 0.429 [0.421, 0.437]
gemma-3-12b-it Nonsensical-CSV 0.218 [0.211, 0.225] 0.136 [0.131, 0.141]
gemma-3-12b-it Regular-CSV 0.417 [0.409, 0.425] 0.310 [0.304, 0.317]
gemma-3-12b-it HTML 0.483 [0.475, 0.491] 0.399 [0.392, 0.406]
Qwen2.5-7B-Instruct Nonsensical-CSV 0.257 [0.250, 0.263] 0.183 [0.178, 0.188]
Qwen2.5-7B-Instruct Regular-CSV 0.339 [0.331, 0.346] 0.285 [0.279, 0.292]
Qwen2.5-7B-Instruct HTML 0.431 [0.424, 0.439] 0.380 [0.374, 0.387]
Meta-Llama-3-8B-Instruct Nonsensical-CSV 0.207 [0.200, 0.213] 0.149 [0.144, 0.154]
Meta-Llama-3-8B-Instruct Regular-CSV 0.243 [0.236, 0.250] 0.193 [0.187, 0.199]
Meta-Llama-3-8B-Instruct HTML 0.322 [0.316, 0.328] 0.28 [0.274, 0.286]
Mathstral-7B-v0.1 Nonsensical-CSV 0.160 [0.155, 0.166] 0.130 [0.125, 0.135]
Mathstral-7B-v0.1 Regular-CSV 0.200 [0.194, 0.207] 0.188 [0.182, 0.194]
Mathstral-7B-v0.1 HTML 0.271 [0.264, 0.278] 0.276 [0.269, 0.283]
Llama-3.1-8B-Instruct Nonsensical-CSV 0.062 [0.058, 0.066] 0.051 [0.048, 0.054]
Llama-3.1-8B-Instruct Regular-CSV 0.155 [0.150, 0.161] 0.139 [0.134, 0.144]
Llama-3.1-8B-Instruct HTML 0.274 [0.267, 0.281] 0.242 [0.235, 0.248]
Table 13: Precision and Recall with 95% confidence intervals for results in Table 3
Model Finetuning Set Precision [95% CI] Recall [95% CI] CC [95% CI]
FanarC-9B No FT 0.262 [0.236, 0.287] 0.234 [0.210, 0.258] 0.171 [0.148, 0.195]
FanarC-9B Real 0.638 [0.611, 0.665] 0.646 [0.619, 0.673] 0.552 [0.521, 0.583]
FanarC-9B Real + Synth 0.864 [0.859, 0.869] 0.867 [0.862, 0.872] 0.734 [0.726, 0.744]
Llama-3-8B-Instruct No FT 0.304 [0.278, 0.329] 0.279 [0.255, 0.303] 0.183 [0.158, 0.207]
Llama-3-8B-Instruct Real 0.600 [0.572, 0.628] 0.617 [0.589, 0.645] 0.525 [0.494, 0.556]
Llama-3-8B-Instruct Real + Synth 0.841 [0.834, 0.847] 0.849 [0.843, 0.855] 0.706 [0.696, 0.716]
Table 14: Precision, Recall, and CC with 95% confidence intervals for Large Text-Only and Medium Text-Only Models on Real World HCTs only from Table 1
Model Dataset Precision [95% CI] Recall [95% CI] CC [95% CI]
gpt-o1 ALL 0.691 [0.682, 0.700] 0.685 [0.676, 0.694] 0.641 [0.632, 0.651]
gpt-o1 arxiv 0.740 [0.717, 0.763] 0.732 [0.710, 0.755] 0.699 [0.674, 0.723]
gpt-o1 qnpc 0.670 [0.659, 0.680] 0.665 [0.654, 0.675] 0.617 [0.606, 0.628]
gpt-o1 us census 0.775 [0.735, 0.815] 0.756 [0.716, 0.797] 0.716 [0.671, 0.761]
gpt-o1 pak 0.741 [0.715, 0.768] 0.742 [0.715, 0.768] 0.708 [0.679, 0.736]
gpt-4o ALL 0.635 [0.626, 0.644] 0.616 [0.607, 0.625] 0.531 [0.521, 0.541]
gpt-4o qnpc 0.627 [0.617, 0.638] 0.606 [0.596, 0.617] 0.515 [0.504, 0.527]
gpt-4o arxiv 0.646 [0.622, 0.670] 0.630 [0.606, 0.655] 0.565 [0.538, 0.591]
gpt-4o us census 0.601 [0.555, 0.648] 0.591 [0.544, 0.637] 0.535 [0.485, 0.585]
gpt-4o pak 0.691 [0.663, 0.718] 0.677 [0.650, 0.704] 0.600 [0.569, 0.630]
gpt-35 ALL 0.419 [0.410, 0.428] 0.425 [0.416, 0.434] 0.329 [0.320, 0.339]
gpt-35 qnpc 0.412 [0.402, 0.422] 0.415 [0.405, 0.426] 0.315 [0.304, 0.326]
gpt-35 arxiv 0.450 [0.426, 0.474] 0.468 [0.444, 0.493] 0.386 [0.360, 0.412]
gpt-35 us census 0.375 [0.330, 0.420] 0.395 [0.349, 0.441] 0.333 [0.286, 0.381]
gpt-35 pak 0.443 [0.415, 0.471] 0.447 [0.419, 0.475] 0.356 [0.326, 0.387]
Qwen2.5-72B-Instruct ALL 0.579 [0.570, 0.588] 0.586 [0.577, 0.595] 0.490 [0.480, 0.500]
Qwen2.5-72B-Instruct qnpc 0.574 [0.564, 0.585] 0.582 [0.572, 0.593] 0.479 [0.467, 0.490]
Qwen2.5-72B-Instruct arxiv 0.594 [0.570, 0.619] 0.595 [0.570, 0.619] 0.521 [0.494, 0.547]
Qwen2.5-72B-Instruct us census 0.505 [0.458, 0.552] 0.519 [0.472, 0.566] 0.465 [0.415, 0.515]
Qwen2.5-72B-Instruct pak 0.622 [0.594, 0.650] 0.629 [0.601, 0.657] 0.541 [0.510, 0.573]
Llama-3.1-70B-Instruct ALL 0.490 [0.481, 0.499] 0.475 [0.466, 0.484] 0.381 [0.371, 0.390]
Llama-3.1-70B-Instruct qnpc 0.489 [0.478, 0.500] 0.474 [0.464, 0.485] 0.374 [0.363, 0.385]
Llama-3.1-70B-Instruct arxiv 0.469 [0.444, 0.494] 0.454 [0.429, 0.479] 0.386 [0.359, 0.412]
Llama-3.1-70B-Instruct us census 0.477 [0.429, 0.524] 0.473 [0.426, 0.520] 0.411 [0.362, 0.460]
Llama-3.1-70B-Instruct pak 0.531 [0.502, 0.559] 0.515 [0.486, 0.543] 0.413 [0.382, 0.444]
gemma-2-27b-it ALL 0.477 [0.468, 0.486] 0.469 [0.460, 0.478] 0.365 [0.356, 0.375]
gemma-2-27b-it qnpc 0.473 [0.462, 0.483] 0.465 [0.455, 0.476] 0.356 [0.345, 0.368]
gemma-2-27b-it arxiv 0.479 [0.454, 0.504] 0.464 [0.439, 0.488] 0.379 [0.353, 0.405]
gemma-2-27b-it us census 0.396 [0.350, 0.442] 0.406 [0.360, 0.453] 0.351 [0.304, 0.399]
gemma-2-27b-it pak 0.536 [0.507, 0.564] 0.525 [0.497, 0.553] 0.415 [0.384, 0.446]
Mixtral-8x7B-Instruct-v0.1 ALL 0.284 [0.276, 0.292] 0.286 [0.278, 0.295] 0.209 [0.201, 0.217]
Mixtral-8x7B-Instruct-v0.1 qnpc 0.279 [0.270, 0.288] 0.278 [0.268, 0.287] 0.197 [0.187, 0.206]
Mixtral-8x7B-Instruct-v0.1 arxiv 0.306 [0.284, 0.328] 0.322 [0.299, 0.346] 0.259 [0.235, 0.282]
Mixtral-8x7B-Instruct-v0.1 us census 0.248 [0.208, 0.288] 0.274 [0.232, 0.316] 0.227 [0.185, 0.269]
Mixtral-8x7B-Instruct-v0.1 pak 0.307 [0.281, 0.333] 0.307 [0.280, 0.333] 0.229 [0.202, 0.255]
gemma-3-27b-it ALL 0.448 [0.439, 0.457] 0.431 [0.422, 0.440] 0.350 [0.341, 0.360]
gemma-3-27b-it qnpc 0.441 [0.430, 0.451] 0.424 [0.413, 0.434] 0.338 [0.327, 0.349]
gemma-3-27b-it arxiv 0.460 [0.435, 0.485] 0.440 [0.415, 0.465] 0.377 [0.351, 0.403]
gemma-3-27b-it us census 0.419 [0.373, 0.465] 0.417 [0.371, 0.464] 0.359 [0.311, 0.407]
gemma-3-27b-it pak 0.500 [0.471, 0.530] 0.480 [0.450, 0.509] 0.403 [0.373, 0.434]
Llama-3.3-70B-Instruct ALL 0.471 [0.462, 0.480] 0.497 [0.488, 0.507] 0.411 [0.402, 0.421]
Llama-3.3-70B-Instruct qnpc 0.467 [0.456, 0.477] 0.494 [0.483, 0.505] 0.401 [0.389, 0.412]
Llama-3.3-70B-Instruct arxiv 0.453 [0.429, 0.478] 0.476 [0.451, 0.502] 0.415 [0.389, 0.442]
Llama-3.3-70B-Instruct us census 0.456 [0.410, 0.502] 0.489 [0.442, 0.536] 0.432 [0.382, 0.481]
Llama-3.3-70B-Instruct pak 0.531 [0.502, 0.559] 0.555 [0.526, 0.584] 0.475 [0.444, 0.506]
Table 15: Precision, Recall, and CC with 95% confidence intervals for Small Text-Only models on Real World HCTs only from Table 1
Model Dataset Precision [95% CI] Recall [95% CI] CC [95% CI]
Qwen1.5-7B-Chat ALL 0.183 [0.175, 0.190] 0.170 [0.163, 0.177] 0.144 [0.138, 0.151]
Qwen1.5-7B-Chat qnpc 0.181 [0.173, 0.190] 0.168 [0.160, 0.176] 0.140 [0.132, 0.148]
Qwen1.5-7B-Chat arxiv 0.194 [0.174, 0.215] 0.185 [0.164, 0.205] 0.169 [0.149, 0.189]
Qwen1.5-7B-Chat us census 0.156 [0.121, 0.191] 0.157 [0.122, 0.192] 0.140 [0.105, 0.174]
Qwen1.5-7B-Chat pak 0.187 [0.164, 0.211] 0.172 [0.149, 0.194] 0.145 [0.123, 0.167]
Qwen2-7B-Instruct ALL 0.242 [0.235, 0.249] 0.299 [0.290, 0.307] 0.216 [0.208, 0.224]
Qwen2-7B-Instruct qnpc 0.240 [0.232, 0.248] 0.297 [0.288, 0.307] 0.208 [0.198, 0.217]
Qwen2-7B-Instruct arxiv 0.235 [0.216, 0.255] 0.302 [0.279, 0.325] 0.251 [0.227, 0.274]
Qwen2-7B-Instruct us census 0.227 [0.190, 0.264] 0.289 [0.246, 0.331] 0.243 [0.200, 0.286]
Qwen2-7B-Instruct pak 0.274 [0.250, 0.298] 0.310 [0.284, 0.336] 0.215 [0.189, 0.240]
Qwen2.5-7B-Instruct ALL 0.350 [0.342, 0.358] 0.382 [0.374, 0.391] 0.273 [0.265, 0.282]
Qwen2.5-7B-Instruct qnpc 0.350 [0.341, 0.359] 0.380 [0.370, 0.390] 0.264 [0.254, 0.275]
Qwen2.5-7B-Instruct arxiv 0.337 [0.315, 0.359] 0.373 [0.349, 0.396] 0.291 [0.266, 0.315]
Qwen2.5-7B-Instruct us census 0.270 [0.231, 0.310] 0.336 [0.291, 0.381] 0.284 [0.239, 0.329]
Qwen2.5-7B-Instruct pak 0.396 [0.370, 0.422] 0.429 [0.401, 0.457] 0.312 [0.282, 0.341]
Mathstral-7B-v0.1 ALL 0.278 [0.270, 0.286] 0.267 [0.259, 0.275] 0.202 [0.194, 0.210]
Mathstral-7B-v0.1 qnpc 0.274 [0.264, 0.283] 0.263 [0.254, 0.272] 0.192 [0.183, 0.201]
Mathstral-7B-v0.1 arxiv 0.290 [0.267, 0.313] 0.278 [0.256, 0.301] 0.237 [0.214, 0.260]
Mathstral-7B-v0.1 us census 0.266 [0.224, 0.309] 0.268 [0.226, 0.310] 0.233 [0.190, 0.275]
Mathstral-7B-v0.1 pak 0.301 [0.274, 0.328] 0.284 [0.258, 0.310] 0.213 [0.188, 0.239]
gemma-3-12b-it ALL 0.418 [0.410, 0.427] 0.372 [0.364, 0.381] 0.268 [0.259, 0.277]
gemma-3-12b-it qnpc 0.421 [0.410, 0.431] 0.372 [0.363, 0.382] 0.261 [0.250, 0.271]
gemma-3-12b-it arxiv 0.386 [0.362, 0.411] 0.352 [0.328, 0.375] 0.280 [0.256, 0.304]
gemma-3-12b-it us census 0.355 [0.310, 0.400] 0.352 [0.307, 0.397] 0.295 [0.249, 0.340]
gemma-3-12b-it pak 0.471 [0.442, 0.500] 0.410 [0.382, 0.437] 0.297 [0.269, 0.326]
gemma-2-9b-it ALL 0.373 [0.364, 0.381] 0.351 [0.343, 0.360] 0.247 [0.239, 0.256]
gemma-2-9b-it qnpc 0.377 [0.367, 0.387] 0.353 [0.343, 0.363] 0.243 [0.233, 0.253]
gemma-2-9b-it arxiv 0.348 [0.325, 0.370] 0.345 [0.322, 0.368] 0.265 [0.241, 0.288]
gemma-2-9b-it us census 0.287 [0.244, 0.329] 0.299 [0.256, 0.343] 0.253 [0.210, 0.297]
gemma-2-9b-it pak 0.409 [0.381, 0.437] 0.368 [0.341, 0.394] 0.254 [0.227, 0.282]
Mistral-7B-Instruct-v0.3 ALL 0.209 [0.201, 0.216] 0.206 [0.199, 0.213] 0.159 [0.152, 0.166]
Mistral-7B-Instruct-v0.3 qnpc 0.208 [0.199, 0.216] 0.201 [0.193, 0.210] 0.151 [0.143, 0.160]
Mistral-7B-Instruct-v0.3 arxiv 0.215 [0.194, 0.235] 0.222 [0.201, 0.243] 0.185 [0.164, 0.206]
Mistral-7B-Instruct-v0.3 us census 0.188 [0.151, 0.224] 0.205 [0.166, 0.244] 0.176 [0.138, 0.214]
Mistral-7B-Instruct-v0.3 pak 0.216 [0.193, 0.239] 0.221 [0.197, 0.245] 0.174 [0.150, 0.197]
Meta-Llama-3-8B-Instruct ALL 0.302 [0.294, 0.310] 0.280 [0.273, 0.288] 0.184 [0.176, 0.191]
Meta-Llama-3-8B-Instruct qnpc 0.309 [0.300, 0.318] 0.285 [0.276, 0.294] 0.181 [0.172, 0.190]
Meta-Llama-3-8B-Instruct arxiv 0.269 [0.247, 0.290] 0.255 [0.234, 0.275] 0.184 [0.163, 0.205]
Meta-Llama-3-8B-Instruct us census 0.237 [0.198, 0.276] 0.245 [0.205, 0.285] 0.191 [0.152, 0.231]
Meta-Llama-3-8B-Instruct pak 0.324 [0.298, 0.351] 0.293 [0.268, 0.318] 0.201 [0.176, 0.226]
FanarC ALL 0.267 [0.259, 0.275] 0.243 [0.235, 0.250] 0.182 [0.175, 0.190]
FanarC qnpc 0.270 [0.261, 0.280] 0.243 [0.234, 0.252] 0.178 [0.169, 0.186]
FanarC arxiv 0.273 [0.250, 0.295] 0.258 [0.236, 0.280] 0.215 [0.193, 0.237]
FanarC us census 0.271 [0.228, 0.314] 0.260 [0.218, 0.303] 0.235 [0.193, 0.278]
FanarC pak 0.238 [0.214, 0.263] 0.210 [0.187, 0.233] 0.151 [0.129, 0.174]
Phi-3.5-mini-instruct ALL 0.122 [0.115, 0.128] 0.108 [0.103, 0.114] 0.084 [0.079, 0.089]
Phi-3.5-mini-instruct qnpc 0.114 [0.107, 0.121] 0.102 [0.095, 0.108] 0.078 [0.072, 0.085]
Phi-3.5-mini-instruct arxiv 0.161 [0.142, 0.180] 0.145 [0.127, 0.162] 0.115 [0.098, 0.132]
Phi-3.5-mini-instruct us census 0.081 [0.055, 0.107] 0.076 [0.051, 0.102] 0.062 [0.038, 0.086]
Phi-3.5-mini-instruct pak 0.137 [0.116, 0.157] 0.118 [0.099, 0.137] 0.092 [0.074, 0.110]
aya-23-8B ALL 0.115 [0.109, 0.121] 0.103 [0.097, 0.108] 0.074 [0.069, 0.079]
aya-23-8B qnpc 0.113 [0.106, 0.120] 0.098 [0.092, 0.104] 0.069 [0.063, 0.074]
aya-23-8B arxiv 0.134 [0.117, 0.151] 0.135 [0.118, 0.152] 0.111 [0.094, 0.128]
aya-23-8B us census 0.110 [0.080, 0.140] 0.101 [0.073, 0.129] 0.078 [0.051, 0.104]
aya-23-8B pak 0.111 [0.092, 0.129] 0.095 [0.078, 0.112] 0.063 [0.048, 0.079]
Llama-3.1-8B-Instruct ALL 0.227 [0.220, 0.234] 0.218 [0.210, 0.225] 0.143 [0.136, 0.150]
Llama-3.1-8B-Instruct qnpc 0.223 [0.215, 0.232] 0.215 [0.207, 0.223] 0.137 [0.129, 0.145]
Llama-3.1-8B-Instruct arxiv 0.248 [0.227, 0.269] 0.237 [0.216, 0.258] 0.177 [0.156, 0.197]
Llama-3.1-8B-Instruct us census 0.204 [0.166, 0.242] 0.204 [0.166, 0.242] 0.165 [0.128, 0.203]
Llama-3.1-8B-Instruct pak 0.235 [0.211, 0.258] 0.214 [0.192, 0.236] 0.127 [0.106, 0.148]
Table 16: Precision, Recall, and CC with 95% confidence intervals for all Text-Only models on Synthetic HCTs only from Table 1
Model Dataset Precision [95% CI] Recall [95% CI] CC [95% CI]
FanarC Synthetic 0.179 [0.176, 0.182] 0.145 [0.143, 0.148] 0.074 [0.072, 0.076]
Mixtral-8x7B-Instruct-v0.1 Synthetic 0.170 [0.168, 0.172] 0.159 [0.157, 0.161] 0.070 [0.068, 0.072]
gemma-3-27b-it Synthetic 0.340 [0.337, 0.343] 0.297 [0.295, 0.300] 0.156 [0.153, 0.158]
gpt-4o Synthetic 0.521 [0.518, 0.524] 0.439 [0.436, 0.442] 0.249 [0.246, 0.252]
Qwen1.5-7B-Chat Synthetic 0.138 [0.136, 0.141] 0.104 [0.102, 0.106] 0.047 [0.045, 0.049]
Mistral-7B-Instruct-v0.3 Synthetic 0.144 [0.142, 0.146] 0.115 [0.113, 0.117] 0.051 [0.049, 0.052]
gemma-3-12b-it Synthetic 0.425 [0.422, 0.428] 0.290 [0.287, 0.292] 0.104 [0.102, 0.106]
Phi-4-mini-reasoning Synthetic 0.000 [0.000, 0.000] 0.000 [0.000, 0.000] 0.000 [-0.000, 0.000]
Phi-3.5-mini-instruct Synthetic 0.162 [0.159, 0.164] 0.125 [0.123, 0.127] 0.056 [0.054, 0.058]
aya-23-8B Synthetic 0.068 [0.066, 0.070] 0.058 [0.057, 0.060] 0.034 [0.032, 0.035]
Llama-3.1-70B-Instruct Synthetic 0.451 [0.448, 0.454] 0.427 [0.424, 0.431] 0.257 [0.254, 0.260]
gemma-2-9b-it Synthetic 0.349 [0.346, 0.352] 0.275 [0.272, 0.277] 0.107 [0.105, 0.109]
Llama-3.3-70B-Instruct Synthetic 0.402 [0.399, 0.406] 0.436 [0.433, 0.439] 0.274 [0.270, 0.277]
gemma-2-27b-it Synthetic 0.378 [0.375, 0.381] 0.322 [0.320, 0.325] 0.131 [0.129, 0.134]
gpt-35 Synthetic 0.391 [0.388, 0.394] 0.309 [0.307, 0.312] 0.127 [0.124, 0.129]
gemma-7b-it Synthetic 0.003 [0.003, 0.003] 0.003 [0.002, 0.003] 0.000 [0.000, 0.000]
Mathstral-7B-v0.1 Synthetic 0.211 [0.209, 0.214] 0.190 [0.188, 0.193] 0.091 [0.089, 0.093]
Qwen2-7B-Instruct Synthetic 0.278 [0.275, 0.281] 0.266 [0.263, 0.268] 0.104 [0.102, 0.107]
Meta-Llama-3-8B-Instruct Synthetic 0.235 [0.232, 0.237] 0.181 [0.179, 0.183] 0.067 [0.065, 0.069]
Llama-3.1-8B-Instruct Synthetic 0.153 [0.150, 0.155] 0.132 [0.130, 0.134] 0.058 [0.056, 0.060]
Qwen2.5-72B-Instruct Synthetic 0.502 [0.500, 0.505] 0.491 [0.488, 0.493] 0.244 [0.240, 0.247]
Table 17: Precision, Recall, and CC with 95% confidence intervals for half of Vision-Text models on Real World HCTs only from Table 1 - Part 1
Model Dataset Recall [95% CI] CC [95% CI]
gpt-4o ALL 0.623 [0.614, 0.632] 0.527 [0.517, 0.537]
gpt-4o pak 0.704 [0.678, 0.730] 0.613 [0.582, 0.643]
gpt-4o qnpc 0.615 [0.604, 0.625] 0.513 [0.501, 0.525]
gpt-4o us census 0.396 [0.350, 0.441] 0.328 [0.281, 0.375]
gpt-4o arxiv 0.675 [0.652, 0.698] 0.597 [0.570, 0.623]
InternVL2-4B ALL 0.309 [0.301, 0.317] 0.215 [0.207, 0.223]
InternVL2-4B pak 0.372 [0.346, 0.399] 0.261 [0.234, 0.289]
InternVL2-4B qnpc 0.290 [0.280, 0.299] 0.193 [0.184, 0.202]
InternVL2-4B us census 0.254 [0.213, 0.294] 0.202 [0.161, 0.242]
InternVL2-4B arxiv 0.385 [0.362, 0.408] 0.298 [0.274, 0.323]
InternVL2_5-8B-MPO ALL 0.264 [0.256, 0.271] 0.178 [0.171, 0.186]
InternVL2_5-8B-MPO pak 0.302 [0.277, 0.327] 0.203 [0.178, 0.229]
InternVL2_5-8B-MPO qnpc 0.239 [0.231, 0.248] 0.152 [0.144, 0.160]
InternVL2_5-8B-MPO us census 0.228 [0.189, 0.267] 0.181 [0.142, 0.219]
InternVL2_5-8B-MPO arxiv 0.375 [0.352, 0.399] 0.298 [0.274, 0.323]
Phi-3-vision-128k-instruct ALL 0.411 [0.402, 0.420] 0.309 [0.300, 0.318]
Phi-3-vision-128k-instruct pak 0.445 [0.417, 0.473] 0.338 [0.308, 0.368]
Phi-3-vision-128k-instruct qnpc 0.408 [0.398, 0.418] 0.300 [0.289, 0.311]
Phi-3-vision-128k-instruct us census 0.273 [0.231, 0.315] 0.230 [0.188, 0.272]
Phi-3-vision-128k-instruct arxiv 0.442 [0.418, 0.466] 0.357 [0.332, 0.383]
Idefics3-8B-Llama3 ALL 0.396 [0.388, 0.405] 0.288 [0.279, 0.297]
Idefics3-8B-Llama3 pak 0.458 [0.431, 0.486] 0.337 [0.307, 0.367]
Idefics3-8B-Llama3 qnpc 0.382 [0.372, 0.392] 0.269 [0.258, 0.279]
Idefics3-8B-Llama3 us census 0.351 [0.307, 0.395] 0.284 [0.239, 0.329]
Idefics3-8B-Llama3 arxiv 0.438 [0.413, 0.462] 0.354 [0.329, 0.380]
Qwen2-VL-7B-Instruct ALL 0.059 [0.055, 0.063] 0.046 [0.042, 0.050]
Qwen2-VL-7B-Instruct pak 0.066 [0.052, 0.081] 0.057 [0.043, 0.072]
Qwen2-VL-7B-Instruct qnpc 0.060 [0.055, 0.065] 0.045 [0.040, 0.050]
Qwen2-VL-7B-Instruct us census 0.071 [0.047, 0.094] 0.052 [0.030, 0.074]
Qwen2-VL-7B-Instruct arxiv 0.045 [0.034, 0.055] 0.038 [0.028, 0.048]
llava-v1.6-mistral-7b-hf ALL 0.143 [0.137, 0.149] 0.097 [0.092, 0.103]
llava-v1.6-mistral-7b-hf pak 0.174 [0.153, 0.196] 0.120 [0.099, 0.140]
llava-v1.6-mistral-7b-hf qnpc 0.139 [0.132, 0.146] 0.092 [0.085, 0.099]
llava-v1.6-mistral-7b-hf us census 0.105 [0.076, 0.135] 0.088 [0.060, 0.116]
llava-v1.6-mistral-7b-hf arxiv 0.150 [0.132, 0.167] 0.113 [0.096, 0.130]
Table 18: Precision, Recall, and CC with 95% confidence intervals for half of Vision-Text models on Real World HCTs only from Table 1 - Part 2
Model Dataset Recall [95% CI] CC [95% CI]
llava-1.5-7b-hf ALL 0.091 [0.086, 0.096] 0.069 [0.064, 0.074]
llava-1.5-7b-hf pak 0.109 [0.090, 0.128] 0.095 [0.077, 0.113]
llava-1.5-7b-hf qnpc 0.094 [0.088, 0.100] 0.068 [0.062, 0.074]
llava-1.5-7b-hf us census 0.085 [0.059, 0.110] 0.062 [0.038, 0.086]
llava-1.5-7b-hf arxiv 0.067 [0.054, 0.080] 0.056 [0.044, 0.068]
Molmo-7B-D-0924 ALL 0.394 [0.385, 0.403] 0.298 [0.288, 0.307]
Molmo-7B-D-0924 pak 0.472 [0.444, 0.500] 0.368 [0.337, 0.398]
Molmo-7B-D-0924 qnpc 0.376 [0.366, 0.386] 0.276 [0.265, 0.286]
Molmo-7B-D-0924 us census 0.282 [0.241, 0.324] 0.225 [0.183, 0.267]
Molmo-7B-D-0924 arxiv 0.465 [0.440, 0.489] 0.383 [0.357, 0.409]
paligemma2-10b-ft-docci-448 ALL 0.098 [0.092, 0.103] 0.065 [0.060, 0.070]
paligemma2-10b-ft-docci-448 pak 0.111 [0.093, 0.129] 0.081 [0.064, 0.098]
paligemma2-10b-ft-docci-448 qnpc 0.102 [0.096, 0.108] 0.066 [0.060, 0.072]
paligemma2-10b-ft-docci-448 us census 0.093 [0.067, 0.119] 0.062 [0.038, 0.086]
paligemma2-10b-ft-docci-448 arxiv 0.065 [0.053, 0.077] 0.048 [0.037, 0.060]
Pixtral-12B-2409 ALL 0.459 [0.450, 0.467] 0.343 [0.334, 0.353]
Pixtral-12B-2409 pak 0.532 [0.504, 0.559] 0.404 [0.374, 0.435]
Pixtral-12B-2409 qnpc 0.435 [0.425, 0.445] 0.312 [0.301, 0.323]
Pixtral-12B-2409 us census 0.400 [0.354, 0.446] 0.344 [0.296, 0.391]
Pixtral-12B-2409 arxiv 0.547 [0.523, 0.571] 0.465 [0.439, 0.492]
llava-v1.6-vicuna-7b-hf ALL 0.163 [0.157, 0.170] 0.111 [0.105, 0.118]
llava-v1.6-vicuna-7b-hf pak 0.204 [0.181, 0.226] 0.136 [0.114, 0.157]
llava-v1.6-vicuna-7b-hf qnpc 0.158 [0.151, 0.166] 0.106 [0.099, 0.113]
llava-v1.6-vicuna-7b-hf us census 0.132 [0.100, 0.163] 0.098 [0.068, 0.128]
llava-v1.6-vicuna-7b-hf arxiv 0.169 [0.150, 0.187] 0.125 [0.108, 0.143]
Phi-3.5-vision-instruct ALL 0.272 [0.264, 0.279] 0.186 [0.178, 0.194]
Phi-3.5-vision-instruct pak 0.365 [0.339, 0.392] 0.261 [0.234, 0.289]
Phi-3.5-vision-instruct qnpc 0.244 [0.236, 0.253] 0.159 [0.150, 0.167]
Phi-3.5-vision-instruct us census 0.178 [0.142, 0.213] 0.140 [0.105, 0.174]
Phi-3.5-vision-instruct arxiv 0.375 [0.351, 0.398] 0.290 [0.266, 0.314]
For confidence intervals of the results shown in Table 1: (i) large and medium text-only LLMs on Real World HCTs are in Table 14, (ii) small text-only LLMs on Real World HCTs in Table 15, (iii) all text-only LLMs on synthetic HCTs in Table 16, (iv) all vision models on Real World HCTs in Tables 17 and  18.
For confidence intervals of the results shown in the format variation experiment (Table 2) and the finetuning experiment (Table 3) please see Table 12 and Table 13 respectively.
Appendix E Extraction Techniques for Real World HCTs
E.1 Canonical Representation
As described in Section 3.1 we define a canonical CSV representation of HCTs that appropriately handle every property. Column Nesting is handled by ensuring the hierarchy is maintained and that each parent is above the left most (for English) child. Row Nesting is handled similarly by ensuring that the parent is there in the row of its top most child. Row group labels should be at the left most of the row and the remaining cell values should be empty. Aggregations and all other columns should be parsed and represented properly without error. An example of our canonical CSV representation can be found in Figure 42.
Figure 42: A HCT along with its canonical CSV representation
E.2 Azure Document Intelligence AI Performance
The Azure Document Intelligence tool extracts the canonical CSV representation from the HCT nearly perfectly. Here we give some examples of its output on our 4 real world data sources.
Figure 43: Azure Document Intelligence extraction on QNPC Table
Figure 44: Azure Document Intelligence extraction on arXiv Table
Figure 45: Azure Document Intelligence extraction on US CENSUS Table
Figure 46: Azure Document Intelligence extraction on PAK Table
It is important to note that there are rare cases where Azure Document Intelligence does have some issues with the extracted CSV. Figure 47 shows such a case where the tool adds an extra column on the left side of the table. Upon manually eyeballing random subsets of the extracted CSVs we found issues such as this to be very rare.
Figure 47: Imperfect extraction Azure Document Intelligence on QNPC Table