### v3

# Mohammad S. Ahmad    Zan Naeem*    Michael Aupetit    Ahmed Elmagarmid    Mohamed Eltabakh    Xiaosong Ma* * Work done while at QCRI.    Mourad Ouzzani    Chaoyi Ruan*    Hani Al-Sayeh

Abstract
Tabular data embedded in PDF files, web pages, and other types of documents is prevalent in various domains. These tables, which we call human-centric tables (HCTs for short), are dense in information but often exhibit complex structural and semantic layouts. To query these HCTs, some existing solutions focus on transforming them into relational formats. However, they fail to handle the diverse and complex layouts of HCTs, making them not amenable to easy querying with SQL-based approaches. Another emerging option is to use Large Language Models (LLMs) and Vision Language Models (VLMs). However, there is a lack of standard evaluation benchmarks to measure and compare the performance of models to query HCTs using natural language. To address this gap, we propose the Human-Centric Tables Question-Answering extensive benchmark (HCT-QA) consisting of thousands of HCTs with several thousands of natural language questions with their respective answers. More specifically, HCT-QA includes 1,880 real-world HCTs with 9,835 QA pairs in addition to 4,679 synthetic HCTs with 67.7K QA pairs. Also, we show through extensive experiments the performance of 25 and 9 different LLMS and VLMs, respectively, in an answering HCT-QA’s questions. In addition, we show how finetuning an LLM on HCT-QA improves F1 scores by up to 25 percentage points compared to the off-the-shelf model. Compared to existing benchmarks, HCT-QA stands out for its broad complexity and diversity of covered HCTs and generated questions, its comprehensive metadata enabling deeper insight and analysis, and its novel synthetic data and QA generator.
I Introduction
Answering natural language (NL) questions over structured and semi-structured data is a fundamental problem [62, 56, 1, 49]. The task of Table QA focuses on retrieving or computing answers to NL questions grounded in tabular data [21, 65, 6, 16, 61, 17]. Early studies used primarily relational tables from Wikipedia and widely used databases [65, 33, 23]. Over time, research has evolved in two directions: (1) Direct question answering using neural models [21, 37, 25], and (2) translating NL questions into SQL queries (NL-to-SQL) [32, 6, 16, 61, 17]. Despite increasing attention to this area, existing efforts remain concentrated on relational or relational-like table structures.
Tabular data embedded in PDFs files, web pages, and other types of documents is widespread across many domains. Figure 1 illustrates examples of such data obtained from the Qatar National Planning Council [51], scientific publication archives [2], the US Census [58], and the Pakistan Bureau of Statistics [48]. These tables employ intricate structural and visual layouts to convey large volumes of information in a concise representation to support strategic decision-making.
Figure 1: Examples of Human-Centric Tables (HCTs) from different real-world data sources exhibiting various levels of structural and semantic complexity including column nesting (L1), row nesting (L2), row group labels (L3), row aggregation (L4), and column aggregation (L5)
Unlike traditional relational tables, which have a simple layout, document-embedded tables are primarily designed for human readability rather than automated processing. We refer to this form of tabular data as Human-Centric Tables (HCTs). HCTs represent a distinct data modality, characterized by intricate layouts, rich semantic content, and significant practical value. However, systematically processing HCTs for tasks such as information extraction, interpretation, and question answering remains a major challenge.
With the rapid advances in multimodal document understanding [55, 36, 39, 35], the scope of Table QA is progressively extending beyond relational tables [17, 16, 42]. This evolution warrants the need for not only systems capable of robust question answering over HCTs, but also methodologies and benchmarks that enable rigorous evaluation and identification of existing limitations of these systems while answering questions on HCTs. Although a few benchmarks have been recently introduced, they suffer from key shortcomings—most notably, the absence of metadata-driven analyses (i.e., analyses incorporating HCTs and question characteristics), and limited scalability due to the reliance on extensive manual effort.
Properties and Challenges of HCTs. HCTs exhibit highly diverse structural features designed for human interpretation. Figure 1 showcases common structural properties observed in HCTs including column nesting (L1), row nesting (L2), and row group labels (L3), as well as embedded aggregations that may appear row-wise (L4) or column-wise (L5). Visual cues-such as variations in text or row color (e.g., L3 in Table (d) and L4 in Table (c)), bold or emphasized text (e.g., L4 in Table (a) and Table (c)), and the positioning of elements within cells (e.g., L2 in Table (b)) convey additional semantic information.
These structural and visual complexities make direct question answering with neural QA models challenging. Also, as HCTs are not inherently relational, SQL queries cannot be executed on them directly. Existing efforts, such as AutoTables [34], have attempted to transform HCT-like tables into relational formats. However, such approaches are limited to relatively simple layouts, often struggle on complex real-world HCTs. Consequently, the pipeline that first converts HCTs to relational tables and then applies NL to SQL transformations remains fragile and error-prone [29, 17, 16, 42].
New Opportunities with LLMs and VLMs. Recent advances in Large Language Models (LLMs) and Vision Language Models (VLMs) pave the way for a new LLM-oriented approach for querying HCTs. This approach can (1) efficiently handle HCTs alongside other human-centric data formats (e.g., charts and images), (2) avoid the complex and error-prone NL-to-SQL pipeline, and (3) directly understand NL questions. However, a systematic evaluation to identify the strengths, weaknesses, and potential improvements of LLM-oriented approaches remains largely an uncharted area. To address this gap, we propose HCT-QA, a novel benchmark designed to assess LLMs and VLMs as the engines for querying HCTs. HCT-QA can be also used to evaluated other types of data processing engines designed to answer NL questions on HCTs.
TABLE I: Comparison of major Table QA benchmarks with HCT-QA.
Benchmark Name Data Type Num HCT Tables Table Modalities Num Sources Num QA Pairs Table Metadata QA Metadata Avg. Questions Per Table Question Syntactic Similarity
HiTab [13] Tables 3,597 JSON 3 10,672 3 0.45
MultiHiertt [69] Tables + Text 9,842 HTML 2 10,440 N/A 0.23
TableVQA-Bench [30] Tables 205 HTML, Image 1 250 1.2 0.186
TAT-QA [71] Tables + Text 2,757 2D List 1 12,650 N/A 0.2879
HCT-QA (Ours) Tables 6,559 HTML, CSV, Image, MD 4 + Synthetic 77,582 11 0.325
Challenges and Contributions. Devising the proposed benchmark involves several challenges: (1) Collecting and generating HCTs with an extensive diversity of complex layouts (Figure 1). (2) Constructing an extensive variety of NL questions ranging from simple selection to complex aggregation and ranking questions. (3) Obtaining accurate ground truth answers, and developing meaningful evaluation metrics to compare them to LLMs’ answers. (4) Exploring the impact of LLM-specific strategies such as finetuning on answering questions on HCTs. The key contributions of HCT-QA are:
  • A large set of HCTs and question-answer pairs (QA) consisting of 1,880 HCTs from four real-world data sources spanning diverse domains. HCTs are provided in various formats including images, CSV, HTML, and Markdown. We also provide 9,835 QA pairs crafted by experts and verified by annotators (Section III-A).
  • Extensive manually-obtained metadata of the HCTs and QA pairs to capture the most common properties of HCTs and questions. This metadata allows for a deeper understanding of LLM/VLM performance and a better identification of performance gaps (Section III-B).
  • A novel synthetic HCT generator to complement the real-world datasets. We used the generator to produce 4,679 HCTs and 67,747 template-based QA pairs across seven semantic domains. The generator offers flexible configuration making it a valuable and practical resource for the research community (Section IV).
  • An extensive evaluation of $25$ LLMs and $9$ VLMs of different sizes ranging from $3B$ to $100B+$ parameters. Our evaluation reveals that large closed-weight models perform the best overall, the promise of processing HCT images directly using VLMs, and the effectiveness of finetuning some models using HCT-QA (Section V).
Our benchmark is publicly available on HuggingFace [19]. All experimentation code and the synthetic HCT generator are publicly available on our GitHub repository [18].
II Related work
Several research and commercial efforts addressed document understanding and information extraction from complex documents [55, 36, 35]. However, none of these fully address the core question What are the opportunities and challenges of using LLM-based approaches to query HCTs?
Table extraction techniques in several commercial tools [3, 14, 22, 15] and recent research [7, 44, 40, 59, 45, 28, 47, 52] focus on extracting tables similar to HCTs from PDFs into machine-readable formats such as CSV and Markdown. However, they do not provide end-users any systematic support for querying, analyzing, or discovering the extracted tables. Also, our analysis shows that VLMs can bypass this extraction step by operating directly on the image version of HCTs.
Table QA approaches focus on relational tables as they use NL-to-SQL to answer the questions [63, 4, 60, 5, 54, 38]. However, SQL queries cannot be executed on tables with complex layouts like HCTs. As mentioned above, AutoTables [34] fail to achieve high accuracy on complex real-world HCTs. Other approaches tried to perform repairs and transformations on non-relational tables but still do not present a means to convert a HCT into a relational table [20, 26]. Such a multi-step, database-oriented pipeline-spanning HCTs extraction, schema discovery, and NL-to-SQL transformation introduces several error-prone phases and could be an overkill for practical applications. Other non-NL-to-SQL related approaches such as direct inference using deep learning models have again been restricted to handling relational tables [66, 21, 37, 8].
Table QA benchmarks mostly focus on flat Wikipedia tables which, when compared to HCTs, are significantly simpler [64, 23, 43, 10, 9, 11, 68, 72]. FinQA  [12] consist of tables from the finance domain but specifically exclude tables with complex nested structures. There are also several benchmarks that cover only one source or domain and have a very small number of QA pairs and tables such as AIT-QA[27] with only 515 QA pairs on 116 tables and TableVQA-Bench [30] with only 250 QA pairs on 205 HCTs. Some benchmarks focus on multi-modal QA where questions are on text, tables, images or other modalities [50, 70, 69, 67, 57, 41]. Compared to these benchmarks, HCT-QA focuses on complex-layout HCTs from diverse real-world sources, provides tens of thousands of QA pairs on thousands of HCTs without requiring metadata, and provides a synthetic HCT and QA pair generator for scalability.
In Table I, we present the key distinguishing dimensions of the related benchmarks. It can be observed that each of them suffers from key limitations like a limited number of HCTs or QA pairs, a small number of questions per table, etc. Also, as the table shows, the most distinctive advantage of HCT-QA lies in its comprehensive table- and QA-level metadata, which enables fine-grained and metadata-driven analysis.
In addition, to assess the diversity of questions, we introduced two statistics: the average number of questions per table and the question syntactic similarity. The syntactic similarity is computed at the table level (for tables with more than one question) by calculating the pairwise word-level Jaccard similarity between all question pairs, followed by averaging the results. As shown, HCT-QA exhibits a higher average number of questions per table and greater lexical diversity.
Zooming-in on HiTab [13] (the closest to HCT-QA), we see that it includes HCTs from three sources and provides only 10,672 QA pairs. In contrast, HCT-QA encompasses nearly twice as many HCTs drawn from a broader range of sources and offers over seven times more QA pairs. Also, HCT-QA includes the visual representations (images) of the HCTs, thereby enabling evaluation of VLMs. Moreover, a closer examination of samples from existing benchmarks revealed several data quality issues. For instance, after manually inspecting 206 randomly selected QA samples from the HiTab benchmark, we found that 22% of the ground-truth answers were incorrect, 8% of the questions were irrelevant to their associated tables, 37% contained minor grammatical errors, and 12% included typographical mistakes. In developing HCT-QA, we placed a strong emphasis on ensuring data quality through a comprehensive and rigorous verification process.
III Real-world datasets
In this section, we present the real-world dataset. Each of its element consists of a quadruplet in the form of $(T_{HCT},Q_{NL},A,M)$, where $T_{HCT}$ is an HCT object, $Q_{NL}$ is a NL question, $A$ is the ground truth answer, and $M$ is rich metadata about $T_{HCT}$ and $Q_{NL}$. In the following, we first present the details of the HCTs collected from thousands of PDFs across diverse domains (Section III-A), and then present the details of the QA collection process (Section III-B). Figure 2 illustrates the dataset preparation process.
TABLE II: Definitions of HCT (table) properties.
Property Name Definition
Column Nesting Headers/column names are spread across more than one level (row).
  Balanced * The depth of each branch within the nesting is the same throughout the table.
  Unbalanced The depth of each branch within the nesting is not the same.
  Symmetric* Child column names repeat uniformly across table columns where nesting is present.
  Asymmetric* Child column names are different across the table in columns where nesting is present.
Row Nesting A cell spans multiple rows but is displayed as one merged cell.
  Balanced* The depth of each row branch is the same.
  Unbalanced The depth of each row branch is not the same.
  Symmetric* Child cell values repeat across the table uniformly in rows where nesting is present.
  Asymmetric* Child cell values are different across the table in rows where nesting is present.
Col. Aggregation Columns that aggregate all or a subset of other columns.
  Global* A column that aggregates all other numerical columns in the table (excluding local-group aggregation columns).
  Local* A column that aggregates a subset of numerical columns in the table.
  Explicit* An aggregation column where its name clearly states it’s an aggregation (e.g., total).
  Implicit* An aggregation column where its name is ambiguous about whether it’s an aggregation.
Row Aggregation Rows that aggregate all or a subset of other rows.
  Global* A row aggregating all other numerical rows in the table (excluding local-group aggregation rows).
  Local* A row that aggregates a subset of numerical rows in the table.
  Explicit* A row where some textual cell value clearly states it’s an aggregation (e.g., total)
  Implicit* A row where all textual cell values are ambiguous about whether it’s an aggregation.
Others
  Row Group Label A row that acts as a heading/title for either all or a subset of rows below or above it.
  Split Header Cell One cell that is split diagonally (visually) and contains two values.
  Stand. Rel. Table* Table displays none of the other HCT properties (nesting, aggregations, labels).
III-A Real-world HCTs
Data Sources. We collected thousands of PDFs from diverse sources, including Qatar National Planning Council (QNPC)  [51] Research Papers from ArXiv.org, biorxiv.org, and medarch.org, US Census covering PDF reports from the US Census Bureau [58] and Pakistan Census from the official website of the Pakistan government [48]. These sources cover various domains including education, land usage, machine learning, data management, finance, astrophysics, materials science, cryptography, astrophysics, medicine, and security. In total, we extracted 1,880 HCTs.
Extraction and Transformation. We employed the Landing AI tool [31] to generate an HTML representation of the HCTs, from which we further derived the corresponding CSV and Markdown formats. As part of the extraction pipeline, we conducted manual verification on over 10% of the extracted HCTs to ensure the tool’s ability to capture the structural complexities such as column nesting, row nesting, row groupings, and embedded aggregations. The verified sample demonstrated 100% accuracy. Also, we utilized Azure AI Document Intelligence [14] to extract HCT images. A similar 10% manual verification confirmed the absence of errors. Figure 3(a) presents statistics on each source.
Figure 2: Preparing the real-world dataset. (a) Extracting HCTs. (b) Annotating HCTs with table properties. (c) Generating QA pairs with question properties.
Characterization of HCT Properties. Across the broad range of HCTs we investigated, we observed that certain structural layouts are highly prevalent such as column nesting, row nesting, and embedded aggregations whereas others, like nested tables, occur less frequently and are considered outliers. In this study, we focus on common structural patterns and exclude HCTs exhibiting rare or irregular layouts. The dominant layout patterns (referred to as properties) are summarized in Table II. The top-level properties are highlighted in green, while additional sub-properties are specified for cases where finer distinctions are expected to influence the LLM/VLM’s comprehension of HCTs, and hence impact the QA accuracy. For instance, if an HCT exhibits column nesting (L1 in Figure 1), we further specify whether the nesting is balanced—meaning all child nodes occur at the same hierarchical depth—or unbalanced, where nesting levels vary. Referring to Figure 1, Tables (b), (c), and (d) demonstrate balanced column nesting, whereas Table (a) illustrates an unbalanced column nesting. We also record whether the nesting is symmetric—where column names and layouts are consistent across parent nodes—or asymmetric, where they differ. A similar level of detail is captured for the row nesting property (L2 in Figure 1). Embedded aggregation columns and rows also represent a common structural characteristic across HCTs. For these cases, we track whether the aggregation is global (spanning all columns or rows) or local (restricted to a subset), and whether it is explicit, e.g., “Total,” “Sum,” and “Avg.”, or implicit, e.g., “Fruits,” denoting a collection of fruit categories.
Manually creating fine-grained metadata for each HCT object represents a distinctive feature HCT-QA, enabling more in-depth analyses of LLM and VLM performance. Figure 3(a) presents aggregated statistics illustrating the distribution of tables across various property combinations. As shown, the combination of {column nesting, row aggregation, column aggregation} is the most frequent, appearing in 396 HCTs and column nesting is the most prevalent structural property overall, occurring in more than 900 HCTs. To support the systematic annotation of this metadata, we developed a custom annotation interface [18] that allows human experts to visually inspect each HCT and select the corresponding layout properties. Each annotation is subsequently reviewed and verified by a second annotator to ensure accuracy and consistency.
Refer to caption Refer to caption
(a) Statistics of HCTs and QA pairs (b) Main properties of real-world HCTs
Refer to caption Refer to caption
(c) Main properties of real-world QA pairs (d) Main properties of synthetic HCTs
Figure 3: Main properties of the HCTs and QA pairs in HCT-QA.
III-B Questions and Ground-Truth Answers Generation
A question over an HCT object refers to a NL question aimed at retrieving specific information—either directly (e.g., selecting particular value(s)) or through computation (e.g., expressions or aggregations). Our focus is on single-HCT questions. For the real-world dataset, we adopt two approaches for question generation: (1) Manually-generated questions – these are produced by the authors of this paper who are database experts with several years of combined experience in academia and industry, and (2) model-generated questions – these are produced by an LLM. All QA pairs, regardless of their source, are subsequently subject to independent human verification to maintain high dataset quality.
Characterization of Question Properties. We designed a set of properties, inspired by common SQL operations, to systematically characterize each question. These properties capture both the type and complexity of operations—such as filtering, aggregation, rank, and projection—required to derive the answer. For each operation, additional sub-properties are maintained to capture more details on the operation’s complexity. For example, the question properties include the number of filtering conditions, whether the projected columns include expressions or not, the number and type of aggregators (e.g., min, max, and avg.) as well as if they are local (a specific subset of values in the HCT) or global (values across the HCT), and the ranking on plain column or expression. These properties provide valuable insights into the types of questions that models tend to find most challenging. To facilitate capturing these properties, we designed a custom interface as depicted in Figure 4. These properties are only provided for the manually-generated questions.
Figure 4: Question interface for the properties of manually-generated questions
Manually-Generated QA Pairs. To generate a question and its answer, we use the annotation interface shown in Figure 4. This interface enables the user to visualize the table image, formulate a question with its associated answer, and annotate the relevant question properties. Also, the interface presents real-time statistics on the types of questions generated, encouraging the user to create a balanced distribution of questions involving diverse operators, expressions, and aggregations. Each question undergoes verification by two independent persons to ensure quality and consistency. We guarantee that every HCT includes at least one manually-generated question. In total, we produced 3,248 such questions with their properties based on real-world HCTs, with each question requiring approximately 1.5 to 3 minutes to complete.
Model-Generated QA Pairs. To expand our question set beyond, we use OpenAI’s GPT [46] to automatically generate additional questions. The model is provided with the HCT image, a sample manually-generated QA pair from the previous step (using one-shot prompting), and detailed instructions specifying the required constraints, such as the target operations and the desired level of question complexity. The model then produces a set of eight QA pairs per HCT, enabling efficient large-scale question generation. Using this approach, we generated 6,587 additional QA pairs on real-world HCTs in approximately 30 minutes.
To ensure the accuracy and clarity of model-generated questions, we established a rigorous human verification and correction pipeline. Each model-generated QA pair is reviewed by a team of annotators who (i) assess the quality of the question as good, ambiguous or nonsensical and (ii) correct the corresponding answer if it is incorrect or poorly formatted. Questions that are rated as nonsensical are excluded from the dataset, while those marked as ambiguous are re-written. Our annotation team comprised of 14 undergrad students majoring in Computer Science who had experience with databases. Annotators were given a training session by the authors on the task description and did supervised mock exercises of the task. Each QA pair is evaluated by two annotators independently. In cases of disagreement, a third annotator makes the final decision. To monitor the consistency and attentiveness of annotators, we embedded a set of pre-validated, manually-generated questions within their evaluation batches. Annotators who misclassified more than one of these verified questions (e.g., labeling them as ambiguous or nonsensical or incorrectly flagging their answers) were excluded from the annotation pool. Only a single annotator failed our hidden test questions and was replaced.
Our model-based pipeline for QA generation and verification was highly effective and scalable. Overall, only 4% of model-generated questions were marked as nonsensical, 24% as ambiguous, and 72% as good. Moreover, the model’s answers required minor corrections for nearly 50% of the questions. Combining the manually-generated and model-generated questions, we ended up with a total of 9,835 QA pairs over the 1,880 HCTs. In Figure 3(a), we present the distribution of the questions across the different datasets. Moreover, in Figure 3(b), we present aggregated distribution of the question properties over all questions. For example, roughly 2,000 questions involve aggregations, while 1,000 questions involve ranking, and there are 594 yes/no questions. Overall, Figure 3(b) shows a good distribution of NL questions on real-world HCTs ranging from simple yes/no questions to complex combination of aggregation, ranking, column expression, etc.
Manual Effort. Referring to Figure 2, to prepare 1,880 real-world HCTs, a combined 150 manual work hours (MWH) were spent as follows: 15 MWH for HCTs filtering to meet certain standards (e.g., visual quality and no out-of-scope properties), 15 MWH for sample verification (10% from text and image modalities), 60 MWH for annotating HCT properties, and 80 MWH for 2-way verification of the annotated HCTs and their properties. To prepare 9,835 QA pairs, 1020 MWH were spent as follows: 150 MWH for manually creating and annotating QA pairs and their properties, 270 MWH for 2-way verification of the manually created and annotated QA pairs, and 600 MWH to verify model generated QA pairs. The creation and verification of the manual QA pairs was done by the authors, whereas the model generated QA pair verification was done by a team of 14 annotators. Combined, this entire effort took around 1,170 MWH.
IV HCT-QA Synthetic Generator
Figure 5: HCT-QA Synthetic Data Generator workflow
To enable large-scale generation of HCTs with their corresponding QA pairs, we built a synthetic generator comprising two components: a HCT Data Generator (Section IV-A) and a QA Generator (Section IV-B). This generator serves as a valuable resource for the research community, facilitating the exploration of research challenges on HCTs. The generator is highly configurable, allowing it to capture nearly all common HCT and question properties defined earlier. It can also be configured to create QA pairs with multiple operators, which would be difficult to create manually at scale.
The main goal of the HCT-QA Synthetic Generator is to reduce the time and labor cost required to create HCTs and QA pairs on them. The entire process of manually extracting real-world HCTs, creating manually generated or human-verified model-generated QA pairs, and labeling the HCT and question properties takes around 4 minutes per QA pair. In contrast, creating the domain vocabulary for an entire family of synthetic HCTs, specifying the related table and question templates takes about 1 hour in total with the generation of a synthetic HCT instance and a related QA pair from that family taking only less than a second. Thus, the synthetic generator enables scaling up the generation process to several thousands HCTs and QA pairs in a few hours, while still controlling their diversity and complexity through the template design. As we will show throughout this section, the HCT-QA Synthetic Generator has the following four properties:
  1. 1.
    Scalability: It can be easily extended to cover new domains and generate thousands of HCTs and QA pairs with minimal efforts.
  2. 2.
    Guaranteed semantic correctness: Based on a template-based approach, semantic fidelity can be ensured across the constructed HCTs and their generated NL questions.
  3. 3.
    Broad coverage of question complexity: NL questions are generated using various SQL templates spanning different levels of sophistication and operator complexities.
  4. 4.
    Disambiguation: The generated NL questions have a clear interpretation in terms of what they are referring to in the table regardless of their complexity.
The synthetic HCT and QA generation process (Figure 5) begins with domain specifications (a) to create custom table templates (b) and question templates (d) from which, inspired by SQL, sample instances of relational tables $T_{REL}$ and SQL queries $Q_{SQL}$ are generated, respectively. Executing $Q_{SQL}$ on $T_{REL}$ gives ground-truth results $A$ (e). We derive HCT templates (c) from relational table templates to control how to pivot and style the same $T_{REL}$ instance into different variants of HCTs instances $T_{HCT}$. We derive NL templates (f) from SQL templates that shape the transcription of $Q_{SQL}$ queries into NL questions $Q_{NL}$. Results $A$ (e) to $(T_{REL},Q_{SQL})$ instances are used as answers to $(T_{HCT},Q_{NL})$ instances as they are identical by design. Metadata $M$ characterizing HCTs and question properties (Section III) are computed too, forming the $(T_{HCT},Q_{NL},A,M)$ benchmark quadruplets.
Figure 6: (a) Statistics of the synthetic HCTs and QAs for 77 domains. An example of a relational table $T_{REL}$ (b) and a generated $T_{HCT}$ from it (c). (d) Examples of synthetic QAs, where $Q_{SQL}$ is transformed into $Q_{NL}$ using custom templates, and executed on $T_{REL}$ to get the ground-truth answer $A$.
IV-A Synthetic HCT Data Generator
The HCT Data Generator generates HCTs in multiple formats such as CSV, HTML, PDF, and PNG, and supports various stylistic features such as row header indentation, various background colors and fonts, and line separations. Through a combination of configuration files and templates, the generator can reproduce $15$ (marked with “*” in Table II) of the $19$ HCT properties. The remaining 4 properties are not supported in this version due to their complex irregularities.
Domain Specifications (Figure  5(a)). We begin by defining a set of realistic application domains, such as import–export statistics across countries and months over several years. Each domain is associated with a domain vocabulary consisting of (1) a JSON configuration file listing attribute names and their set of possible values and (2) catalog tables listing a comprehensive set of attributes from the vocabulary, their data types, and possible value ranges. Various HCTs are then created by different combinations of attributes and value ranges. For example, in a “Food import–export” domain (Figure 6(b)), user-defined attributes such as Year: {2015,...,2023}, ImportExport Op.: {Import, Export}, Food Category: {Meat, Dairy}, and Food Item: {Chicken, Beef, Cream, Milk, Yogurt} are used to generate HCTs describing food import-export activities over years.
Types of Attributes (Figure  5(a)). We distinguish between two types of attributes, independent attributes such as Year and Import-Export for which all pairs of values are possible, e.g., (2022, Import), and hierarchical attributes such as (Country, State, City) and (Food category, Food item), where attribute values follow an inherent order and only certain combinations are meaningful. For example, combinations such as (California, Boston) are invalid. Independent attributes can appear as either row or column headers at any nesting level. In contrast, hierarchical attributes must respect logical order — parent attributes, such as Country, must precede child ones, such as City, from left to right or top to bottom within headers. Moreover, all attributes belonging to the same hierarchy must be placed together, either entirely in the row headers or entirely in the column headers, e.g., an HCT cannot have Country as a row header and City as a column header. More details are provided in [18].
Relational Table Templates (Figure 5(b)). We manually define Table templates in JSON format that give details about which attributes and values to pick for a given scenario to form a family of relational tables. For example, for the import-export scenario, the data generator uses the table template to create the relational table ($T_{REL}$) depicted in Figure 6(b). The template defines a target field, e.g., the Value numerical attribute in Figure 6(b), to contain the fine-grained value (automatically generated) for each combination over the rest of the attributes. The generator enforces uniqueness constraints on the target field to prevent ambiguity when evaluating model-predicted answers against ground-truth values. Relational tables serve two essential purposes: (1) They enable systematic generation of HCTs through pivoting operations [24, 53]; and (2) they are used to construct the ground truth answers by executing SQL queries on them without human intervention.
HCT Templates (Figure 5(c)). These templates define the final construction rules for producing HCTs instances from the domain specifications and structural combinations set in the relational table template from which they are automatically generated. They are highly configurable, for example, one can specify the depth of row and column headers, the assignment of attributes to each header, and the inclusion and type of aggregation functions and whether the aggregation occurs at a global or local level. They also capture layout specifications such as nesting balance and symmetry, as well as stylistic elements, e.g., line separations, and indentation.
As shown in Figure 6(c), an example HCT instance ($T_{HCT}$) is generated from its corresponding relational table ($T_{REL}$). In this case, attributes (Import_Export, Year) form nested column headers, while (Category, Items) define nested row headers. The column nesting is balanced and symmetric, whereas the row nesting is balanced but asymmetric. This instance includes both local and global aggregation rows and columns (average). Food Category row headers (Beverage and Dairy) are indented and their row contains implicit local average aggregation of food Items (e.g., Coffee and Tea for Beverage), while global aggregation of food Category and column aggregations are all explicit (”Average” appears in the header). These properties are automatically recorded in the metadata $M$. It is worth mentioning that with the relational table and the HCT templates, the Synthetic HCT Data Generator only requires the domain vocabulary to generate HCTs for new domains.
Statistics on Synthetic HCTs. In our benchmark and as summarized in Figure 6(a), we covered seven distinct domains and generated hundreds of HCTs per domain with a total of 4,679 synthetic HCTs. These HCTs exhibit different complexities and layouts as presented in Figure 3(c).
IV-B Synthetic Question & Answer Generator
Our synthetic QA Generator coupled with the HCT Generator, allows us to scale HCT-QA to produce tens of thousands of QA on thousands of synthetic HCTs with $100\%$ accuracy.
Figure 7: Example of a SQL query template, a basic NL template, and its instantiation on the HCT (top) for synthetic query generation.
SQL Query Templates (Figure 5(d). The QA Generator builds upon the same domain specifications and table templates described in Section IV-A. Table III lists 15 SQL query templates, each representing a distinct level of complexity. These templates are used to instantiate SQL queries with varying structures and operations from simple or multiple rows and columns selection (SQL1,2,3,6), to multiple aggregations over one row or one column (SQL4,7), when the same aggregation is already displayed in the HCT (SQL5), or grouped by any attribute levels or only by the top level one, and with or without reporting the attributes (SQL8,9,10,11). We also generate sorting and top-k queries (SQL12,13) as well as selections with operations (SQL14) and a selection conditioned on another selection (SQL15).
As an example, Figure 7 presents a SQL query template (SQL7 in Table III) to generate global aggregation queries (i.e., without sub-groups) on the relational representation $T_{REL}$ (not shown) of an HCT. The template applies one equality condition to each attribute involved in column nesting, allowing it to naturally extend to multi-level nesting scenarios. When applied to the presented HCT, these conditions trace through the structure, one level at a time, to reach a specific column (green & blue colors of the WHERE clauses correspond to the same color boxes in the HCT column headers). Additionally, the template introduces IN conditions on attributes associated with row nesting. The HCT in Figure 7 contains no row nesting, and hence a maximum of one condition is possible (i.e., on the Month column). In contrast, the HCT in Figure 6(c) contains row nesting, and the template allows putting conditions on the corresponding columns (i.e., on the Category and Item columns in $T_{REL}$). These conditions contribute to increasing the query’s complexity. Finally, the template applies one or more aggregation functions over the selected output values.
Ground Truth Generation (Figure  5(e). The relational table $T_{REL}$ and the generated SQL queries are imported into a relational DBMS for execution and the generation of the ground-truth answer $A$, thus forming the triplet ($T_{REL},Q_{SQL},A$). In Figure 6(d), we illustrate three SQL query instances generated from different SQL query templates (SQL9, SQL12, and SQL15 in Table III).
Natural Language Templates (Figures 5(d-f)). For each SQL query template, we manually define a NL template to transcribe $Q_{SQL}$ into $Q_{NL}$ specific to the HCT domain. To illustrate the need for these custom NL templates, let’s consider the SQL query template in Figure 7 that would be transcribed into a (basic) NL template as:
"What is the \$AGG1 and \$AGG2 of Values for \$Col_Level1 equals \$v1 and \$Col_level2 equals \$v2 and \$RowNest_Level1 IN <list>"
Then, a (basic) NL question instance, given some values from the HCT in Figure 7, would be:
"What is the SUM and AVG of Values for Airlines equals ‘Other Airlines’ and Direction equals ‘Arrival’ and Month IN [‘January’, ‘March’, ‘May’]"
Obviously, this basic direct transcription does not form a seamless NL question. Hence, we customize the NL templates for each of the seven domains covered in our generator (Figure 6(a)), generating more natural questions. We illustrate a few examples in Figure 6(d). Similar to the construction of synthetic HCTs, the Synthetic QA generator is domain-agnostic. It relies solely on specifying a domain by its vocabulary describing attributes (e.g. Month) and values (May, June…) in a JSON format, and the manual design of $15$ $Q_{NL}$ templates based on this vocabulary and simple rules to ensure English syntax. Each $Q_{NL}$ template transcribes in natural language one of the $15$ $Q_{SQL}$ templates shown in Table III, ensuring non-repetitive queries. The full list of $15$ $Q_{NL}$ templates for each of the $7$ domains and the rules to build them, is available in Appendix D in the extended report [18]. As each $Q_{NL}$ template corresponds to a $Q_{SQL}$ template with specific query properties, we automatically record them in the metadata $M$.
TABLE III: SQL templates of Synthetic HCT-QA. $r|c$: depth of row||column (R||C) headers of the HCT. $Attr_{C}^{*}$: all attributes of C. $Attr_{R}^{1}$: left-most attribute of R. NumNum: avg. of selected column. $Dir\stackrel{{\scriptstyle\text{rand}}}{{\sim}}\{$ASC, DESC}\}, $k\stackrel{{\scriptstyle\text{rand}}}{{\sim}}\{2,3,4,5\}$, and $Op\stackrel{{\scriptstyle\text{rand}}}{{\sim}}\{<,>\}$.
SQL templates
SQL1 SELECT Value FROM Data WHERE (AttrR = ValR)r AND (AttrC = ValC)c
SQL2 SELECT Value FROM Data WHERE (AttrR IN ListValR)r AND (AttrC = ValC)c
SQL3 SELECT Value FROM Data WHERE (AttrR = ValR)r AND (AttrC IN ListValC)c
SQL4 SELECT AGG1(Value), AGG2(Value) FROM Data WHERE (AttrR = ValR) AND (AttrC IN ListValC)
SQL5 SELECT AGG(Value) FROM Data WHERE (AttrR = ValR)r AND (AttrC IN ListValC)c
SQL6 SELECT Value FROM Data WHERE (AttrR IN ListValR)r AND (AttrC IN ListValC)c
SQL7 SELECT AGG1(Value), AGG2(Value) FROM Data WHERE (AttrR IN ListValR)r AND (AttrC = ValC)c
[…]
SQL11 SELECT $_{R}^{1}, Attr$$_{C}^{*}$, AGG1(Value) FROM Data WHERE (AttrR IN ListValR)r AND (AttrC IN ListValC)c GROUP BY $_{R}^{1}, Attr$$_{C}^{*}$
SQL12 SELECT Value FROM Data WHERE (AttrR IN ListValR)1 AND (AttrC = ValC)c ORDER BY VALUE Dir LIMIT k
SQL13 SELECT Value FROM Data WHERE (AttrR IN ListValR)1 AND (AttrC = ValC)c ORDER BY VALUE Dir
[…]
Statistics on Synthetic Questions. We generated 67,747 QA pairs on the $4,679$ synthetic HCTs across seven domains.
Scalability. Thanks to manually defined JSON templates, our Synthetic Data Generator can be augmented with minimal effort. The domain specifications can be augmented at will, adding attribute names and their list of values, to generate HCTs in any domain of interest. Also, the relational table templates allow controlling the diversity and complexity of the generated HCTs. SQL query templates can be easily configurable due to their computational nature. The corresponding NL templates require minimum effort to adapt to a new domain, using a small set of syntax rules. This template-based automation allows unlimited HCT-QA accurate data generation in minimal time without the need for costly verification.
Manual Effort. A key advantage of the synthetic data generator is that, once the tool is implemented—a process that required several hundred hours of development—adding new SQL or natural language templates for a specific domain typically takes no more than two to three hours.
V Experimental evaluation
While HCT-QA is designed to evaluate data processing engines during answering NL questions on HCTs, our evaluation focus on LLMs and VLMs. To this end, we explain our experimental setup in Section V-A and then answer seven research questions (RQs) experimentally in Section V-B.
V-A Experimental setup
Models. We tested a large variety of open- and closed-weight LLMs and VLMs (Table IV). All open-weight models were downloaded from HuggingFace.
Datasets. There are 1,880 real-world HCTs from four data sources along with 9,835 QA pairs and 4,679 synthetic HCTs with 67,747 QA pairs (total of 6,559 HCTs with 77,582 QA pairs). For the results in Table IV, the text modality HCTs were represented in HTML format and the vision-text modality HCTs were images encoded in base64.
Prompting Approach and Hyperparameters. We experimented with zero-shot and one-shot prompting with no clear indication that the latter improved model performance. Thus, all subsequent experiments use zero-shot prompting. We ask the model to ”only answer the question using information from the HCT ” and allow it to refuse to answer the question if it deems it unanswerable. For inference, the temperature was set to 0, and the maximum output tokens to 256. For finetuning, the learning rate was 5.0e-07 and the number of epochs was 2. The finetuning config file can be found in [18]. We post-process the model responses using a rule-based approach to extract the actual answer if one was provided. All inference and finetuning experiments were run on 8 NVIDIA A100 GPUs. Inference which required around 77k prompts took a total of 129 hours, averaging 1 hour, 1.5 hours, and 22 hours per small, medium, and cloud model, respectively. Additionally, our finetuning experiments took 30 hours.
Evaluation Metrics. We use two metrics for model accuracy: (i) F1 Score–commonly adopted in several table QA benchmarks, e.g.,  [65, 11, 71, 69, 13], and (ii) Complete Containment (CC Score)–a binary score that evaluates to 1 iff the recall is 100%, otherwise it is 0. The two metrics are complementary: F1 rewards partial correctness of the answer while CC rewards an answer only if it contains the ground truth in its entirety. For example, assume the ground truth is [“Apple”, “300”] and three distinct models output: M1 [“Apple”, “300”], M2 [“Apple”], and M3 [“Apple”, “300”, “Orange”, “10”], then their scores will be: M1 (F1= 1.0, CC= 1), M2 (F1= 0.66, CC= 0), and M3 (F1= 0.66, CC= 1).
TABLE IV: Overall performance of models on HCT-QA
Model Name Model Param Size QNPC PAK US CENSUS arXiv Synthetic Tables Average
F1 CC Score F1 CC Score F1 CC Score F1 CC Score F1 CC Score F1 CC Score
TEXT-ONLY LLMs
Large Models – 100B+ (Closed Weight)
ChatGPT 3.5 175B 0.413 0.315 0.445 0.356 0.385 0.333 0.459 0.386 0.345 0.279 0.409 0.334
ChatGPT 4o Unknown 0.663 0.571 0.720 0.639 0.650 0.583 0.719 0.654 0.563 0.365 0.663 0.562
ChatGPT 4.1 Unknown 0.621 0.545 0.640 0.576 0.650 0.600 0.667 0.605 0.543 0.415 0.624 0.548
Average - 0.566 0.477 0.602 0.524 0.562 0.505 0.615 0.548 0.484 0.353 0.566 0.481
Medium Models – 27B–72B (Open Weight)
Llama-3.1-70B-Instruct 70B 0.586 0.484 0.639 0.551 0.585 0.519 0.632 0.555 0.512 0.323 0.591 0.486
Llama-3.3-70B-Instruct 70B 0.573 0.497 0.636 0.571 0.548 0.530 0.645 0.595 0.499 0.361 0.580 0.511
Qwen2.5-72B-Instruct 72B 0.621 0.532 0.673 0.593 0.589 0.540 0.678 0.612 0.582 0.339 0.629 0.523
Qwen3-32B 32B 0.523 0.435 0.578 0.491 0.538 0.479 0.576 0.503 0.458 0.227 0.535 0.427
gemma-2-27b-it 27B 0.549 0.444 0.617 0.514 0.509 0.466 0.631 0.552 0.433 0.207 0.548 0.437
gemma-3-27b-it 27B 0.535 0.445 0.598 0.514 0.527 0.469 0.622 0.560 0.476 0.275 0.552 0.453
Average (Medium) - 0.565 0.473 0.624 0.539 0.550 0.500 0.631 0.563 0.493 0.289 0.572 0.473
Small Models – 4B–20B (Open Weight)
gemma-2-9b-it 9B 0.475 0.360 0.537 0.413 0.409 0.357 0.497 0.422 0.383 0.171 0.460 0.345
gemma-3-12b-it 12B 0.498 0.371 0.545 0.412 0.465 0.401 0.543 0.443 0.449 0.195 0.500 0.364
Qwen1.5-7B-Chat 7B 0.220 0.173 0.268 0.220 0.194 0.182 0.314 0.268 0.124 0.059 0.224 0.180
Qwen2-7B-Instruct 7B 0.333 0.273 0.356 0.305 0.283 0.289 0.375 0.341 0.295 0.107 0.328 0.263
Qwen2.5-7B-Instruct 7B 0.417 0.343 0.454 0.384 0.361 0.383 0.483 0.445 0.334 0.124 0.410 0.336
Qwen3-14B 14B 0.490 0.430 0.535 0.482 0.458 0.454 0.577 0.547 0.394 0.156 0.491 0.414
Mistral-7B-Instruct-v0.1 7B 0.071 0.062 0.077 0.065 0.032 0.027 0.143 0.129 0.062 0.040 0.077 0.065
Mistral-7B-Instruct-v0.2 7B 0.119 0.108 0.120 0.123 0.137 0.124 0.169 0.173 0.059 0.035 0.121 0.112
Mistral-7B-Instruct-v0.3 7B 0.295 0.219 0.321 0.255 0.253 0.222 0.342 0.298 0.151 0.068 0.273 0.213
Mathstral-7B-v0.1 7B 0.326 0.266 0.387 0.320 0.184 0.169 0.362 0.329 0.215 0.097 0.295 0.236
Meta-Llama-3-8B-Instruct 8B 0.344 0.241 0.374 0.268 0.334 0.302 0.376 0.301 0.262 0.100 0.338 0.242
Llama-3.1-8B-Instruct 8B 0.363 0.257 0.393 0.278 0.327 0.290 0.416 0.342 0.310 0.119 0.362 0.257
Phi-3-mini-4k-instruct 4B 0.233 0.167 0.265 0.185 0.223 0.198 0.339 0.267 0.211 0.107 0.254 0.185
Phi-3.5-vision-instruct 4B 0.274 0.193 0.235 0.170 0.190 0.143 0.245 0.189 0.108 0.059 0.211 0.151
Phi-4-mini-instruct 4B 0.225 0.156 0.131 0.086 0.044 0.037 0.154 0.116 0.062 0.021 0.123 0.083
Fanar-C-9B-Instruct 9B 0.431 0.322 0.463 0.351 0.473 0.416 0.522 0.451 0.437 0.217 0.465 0.351
Average (Small) - 0.320 0.246 0.341 0.270 0.273 0.249 0.366 0.316 0.241 0.105 0.308 0.237
VISION-TEXT VLMs
Large Models – 100B+ (Closed Weight)
ChatGPT 4o Vision Unknown 0.543 0.513 0.638 0.613 0.348 0.328 0.618 0.597 - - 0.537 0.513
Small Models – 3B–12B (Open Weight)
Pixtral-12B-2409 12B 0.378 0.289 0.471 0.372 0.300 0.273 0.457 0.411 0.364 0.161 0.394 0.301
InternVL2-4B 4B 0.236 0.181 0.281 0.199 0.148 0.162 0.305 0.297 0.265 0.102 0.247 0.188
InternVL2_5-8B-MPO 8B 0.193 0.129 0.270 0.196 0.177 0.170 0.349 0.288 0.208 0.076 0.240 0.172
Phi-3-vision-128k-instruct 4B 0.214 0.187 0.280 0.252 0.210 0.199 0.339 0.317 0.241 0.135 0.257 0.218
Phi-3.5-vision-instruct 4B 0.208 0.138 0.314 0.227 0.148 0.120 0.332 0.269 0.265 0.133 0.253 0.177
llava-1.5-7b-hf 7B 0.067 0.061 0.096 0.092 0.067 0.057 0.042 0.039 0.030 0.006 0.061 0.051
llava-v1.6-vicuna-7b-hf 7B 0.123 0.091 0.157 0.118 0.095 0.077 0.133 0.107 0.115 0.032 0.125 0.085
Molmo-7B-D-0924 7B 0.130 0.067 0.164 0.100 0.092 0.075 0.147 0.103 0.156 0.027 0.138 0.074
Average (Small) - 0.194 0.143 0.254 0.194 0.155 0.142 0.263 0.229 0.205 0.084 0.214 0.158
Note: ChatGPT-4o Vision was not run on synthetic HCT images due to API constraints
V-B Analysis and findings
(RQ1) How do language models perform overall on HCT-QA? As shown in Table IV, the relatively recent versions of large closed-weight LLMs like ChatGPT 4o and ChatGPT 4.1 on the text modality perform reasonably well, achieving F1 scores of 66% and 62% respectively. This shows that even for state-of-the-art LLMs, there is ample room for improvement on this task. Despite ChatGPT 4o being the best performing model overall, we see that its predecessor ChatGPT 3.5 scores an average F1 of 40% which is less than all of the medium sized open-weight LLMs. This highlights the progress more recent LLMs have made in their reasoning abilities despite their much smaller size. Furthermore, Qwen2.5-72B-Instruct achieves an average F1 of 62.9%, which is only 3.4 percentage points less than ChatGPT 4o’s score, despite its size of 72B parameters being considerably smaller than the OpenAI model. Similarly, Qwen2.5-72B-Instruct’s CC Score of 52% is only 4 percentage points less that ChatGPT 4o CC Score of 56%. On average, large LLMs achieve an F1 of 56% compared to 57% for medium-sized LLMs. Even, when excluding the lagging ChatGPT 3.5, large LLMs have an average F1 score of 64% which is only 7 percentage points higher than the medium-sized LLMs’ average. This showcases the potential for medium-sized LLMs on this task.
With an average F1 of 30%, small LLMs lag behind their larger competitors. Even among small LLMs, gemma-3-12b-it’s average F1 score of 50% outperforms ChatGPT 3.5 by nearly 10 percentage points. However, we see a considerable drop in gemma-3-12b-it’s CC Score; only 36.4% which is far below the average for medium and large models. This indicates that it does get a lot of the questions partially correct but struggles to provide the fully correct answer. Moreover, small VLMs with an average F1 of 21.4% are far worse off than small LLMs. But, among the VLMs, Pixtral-12B outperforms many LLMs such as Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Qwen2-7B-Instruct. This indicates that VLMs do have the potential to compete with similarly sized LLMs on HCT-QA, possibly by utilizing the visual features of HCTs that may be lost in a textual representation. Overall, there remains clear room for improvement for both LLMs and VLMs. Furthermore, the F1 scores of models on synthetic HCTs is, on average, slightly less than on real-world HCTs. This shows that our synthetic generator produces HCTs and QA pairs that are challenging for all models. However, there is a considerable drop in the CC Scores on synthetic HCTs. This could be due to many of the synthetic QA pairs requiring answers with many more values than the QA pairs on real-world HCTs–as the number of values in the answer increases, the models’ ability to generate a completely correct answer significantly degrades. For example, the Qwen2.5-72B-Instruct model under the synthetic HCT tables has the highest F1 score while its CC score is relatively low.
In Figure 8 we also compare how models perform on QA pairs created using our three methods. Synthetic questions appear to be the most challenging which is expected as the majority require Row Filtering, Expression on Columns, and Aggregation together. On average, models of all sizes perform best on the manually-generated QA pairs. This indicates that our model-generated and synthetic questions provide a slightly harder challenge for the evaluated models.
(RQ2) Are models improving over time? There appears to be no consistent trend with regards to newer generations of models outperforming their predecessors. The Qwen family of small LLMs shows continuous improvement as Qwen1.5-7B, Qwen2-7B, Qwen2.5-7B, and Qwen3-14B all outperform their predecessor with average F1 scores of 22.4%, 32.8%, 41%, and 49.1% respectively. This shows that the Qwen family is improving the performance of their small LLMs by an average of 9% per generation. The same can be seen with their CC Scores as Qwen3-14B is 23 percentage points better than Qwen1.5-7B-Chat. This shows that the models are getting much better at giving the complete correct answer. We see similar trends with the Mistral 7B LLMs which also improve by nearly 9% per generation and Gemma LLMs that improve by 4%. Conversely, the Phi family of small LLMs shows consistent deterioration in each successive generation with Phi-3-mini-4k-instruct achieving an average F1 score of 25.4% which is nearly double that of Phi-4-mini-instruct, despite the latter being released 8 months later. The ChatGPT family shows a mixed trend with a significant improvement of 26 percentage points between ChatGPT 3.5 and ChatGPT 4.0, but then ChatGPT 4.1 under-performs its predecessor by 4 percentage points. There are also cases like gemma-2-27b-it and gemma-3-27b-it which shows the newer generation improving by a mere 0.4 percentage points in terms of average F1. We also see a similar trend with small VLMs as Phi-3-vision-128k-instruct outperforms its successor, Phi-3.5-vision-instruct, by 0.4 percentage points. Thus, we cannot say that newer versions of models outperform their older versions across the board.
Figure 8: Model performance by size across different QA creation types
TABLE V: Model F1 scores across different HCTs input formats
c Model CSV Markdown HTML Text Obfuscated HTML ChatGPT 4.1 0.484 0.529 0.543 0.499 ChatGPT 4o 0.541 0.567 0.563 0.531 Qwen2.5-72B-Instruct 0.422 0.453 0.582 0.508 Llama-3.3-70B-Instruct 0.431 0.443 0.499 0.398 Qwen2.5-7B-Instruct 0.301 0.311 0.334 0.358 Llama-3.1-8B-Instruct 0.191 0.262 0.311 0.277 Average 0.395 0.426 0.472 0.42
TABLE VI: Finetuned model performances across different test sets.
Model Name Finetuning Set Test Sets (F1 Score)
Real HCTs Synthetic HCTs HiTab [13]
Llama-3.1-8B-Instruct None 0.351 0.333 0.538
Llama-3.1-8B-Instruct Real HCTs 0.593 0.534 0.606
Llama-3.1-8B-Instruct Synth. HCTs 0.471 0.879 0.561
Llama-3.1-8B-Instruct Real & Synth. HCTs 0.595 0.886 0.616
(RQ3) Does the input format of the HCT matter in text-modality? For the text modality LLMs, we tested the effect of representing the HCT in CSV and Markdown formats instead of HTML. The results are shown in Table V. The HTML format outperforms CSV and Markdown formats by an average F1 score difference of 7.7 and 4.6 percentage points, respectively. This result is intuitive as HTML is a more expressive format that can better retain the complex structure of the HCTs. Some models are more sensitive than others to the table format. Qwen2.5-72B-Instruct performs 13 percentage points better with the HTML format than with Markdown and shows a similar difference between the HTML and CSV formats as well. Whereas, ChatGPT 4o shows a relative indifference to the table format as there is only a 0.2 percentage point difference in F1 scores between the Markdown and HTML formats, with the former performing minutely better. Our results show that while large LLMs like ChatGPT can perform well with both Markdown and HTML, medium and small sized LLMs perform much better with an HTML representation of the HCT.
We also explore the effect of semantically coherent and meaningful text within the HCTs. We created Text Obfuscated-HTML versions of the HCT HTMLs where we (i) replace all column names with ambiguous terms (ii) replace all text cell values in the HCTs with similar ambiguous terms, and (iii) replace the corresponding QA pairs with the respective ambiguous terms from the obfuscated table. Although the complexity of the questions and table structure remains unchanged, we see that the text obfuscation negatively impacts model performance. LLMs on average, experience a decrease in F1 score of 5 percentage points on Text Obfuscated-HTML format when compared to the original HTMLs. Even a large model like ChatGPT 4.1 cannot compensate for the lack of semantics as its F1 score drops by 4.4 percentage points.
(RQ4) Does finetuning models improve performance? We finetuned the Llama-3.1-8B-Instruct model on three FT sets: (i) Real HCTs which has 7,867 (80% of total) HCT-QA triplets on real world HCTs only, (ii) Synth. HCTs which has 54,186 (80% of total) HCT-QA triplets on synthetic HCTs only, (iii) Real + Synth. HCTs which combines both the  Real HCTs and  Synth. HCTs sets (62,053 in total). We then evaluate the models on the respective Real HCTs and Synth. HCTs test sets. We ensure that no HCT in the test set is seen during finetuning. Furthermore, to test the generalization of finetuning in HCT-QA we test on the HiTab dataset [13]. The results are shown in Table VI. Finetuning on the Real + Synth. HCTs set drastically increases the performance of the model across the board. The F1 score improvement, compared to the non-finetuned model is 24.1 percentage points on real HCTs from HCT-QA, 55 percentage points on synthetic HCTs from HCT-QA, and nearly 8 percentage points on the out of domain HiTab dataset. Furthermore, finetuning on the Synth. HCTs set yields an F1 score improvement of 12 percentage points on the real HCTs test set. This reinforces the usefulness of our synthetic data generator since training only on synthetic HCTs generalizes well to real-world HCTs.
TABLE VII: Effect of table properties on F1. Only properties with significant effect ($p<0.05$) are shown
Table Properties Coef. (p-value) Table Properties Coef. (p-value)
Symmetric Row Nesting -0.103 (0.016) Global Column Aggreg. -0.035 (0.028)
Asymmetric Row Nesting -0.142 (0.001) Implicit Col. Aggreg. 0.081 (0.001)
Balanced Col. Nesting -0.270 (0.022) Implicit Row Aggreg. -0.065 (0.007)
Unbalanced Col. Nesting -0.315 (0.009) Row Group Label -0.069 (0.000)
(a) Refer to caption
(b) Refer to caption
Figure 9: F1 scores for models across a subset of question properties (a) and aggregation types (b). Models are ordered by decreasing size from top to bottom, and by decreasing average scores from left to right.
(RQ5) How do HCT properties affect model performance? Table VII shows the regression coefficient of various HCT properties against F1 scores. We can see that the presence of most of these properties, except Implicit Column Aggregation Terms, decreases the F1 score on average. We also see a clear trend in models being able to better understand symmetrical and balanced nesting patterns compared to asymmetrical and unbalanced ones: Symmetric Row Nesting decreases F1 by 10 percentage points whereas Asymmetric Row Nesting decreases it by 14 percentage points. The same trend is seen when comparing Balanced Column Nesting with Unbalanced Column Nesting with the former having a decrease of 27 percentage points on F1 against 31.5 percentage points for the latter. Column and row nesting increases the HCT difficulty more than any type of aggregate rows or columns, having the largest negative impact on F1 scores.
(RQ6) How do question properties affect model performance? Our question metadata allows for fine-grained analysis to understand what types of questions different models excel at and struggle with. Figure 9 shows such an analysis on a subset of models and metadata. Heatmap (A) shows that all models struggle with questions requiring aggregations or expressions on columns. We can also see that Qwen3-14B, despite being better than models like Fanar-C-9B and gemma-3-27b at questions involving ranking and row filtering, pales in comparison to both on yes/no questions. Furthermore, Qwen2.5-72B shows good performance on aggregation questions compared to other models. In general, models perform better with yes/no and row filter questions as this type of question is answered directly from the data cells, similarly to scan, filter and project stateless operators in databases (i.e., a single row/cell is enough to answer the question), while ranking is more challenging for the models because, similar to the ORDER BY stateful operator in databases, it requires a full view of the data. Aggregation is even more challenging because it operates on multiple cells, which can be in different locations of the HCT. In Figure 9), we observe that, for example, end users would be better off using gemma-3-27b than the larger Llama-3.3-70B and Llama-3.1-70B models for questions requiring aggregation, row filtering, and yes/no answers, as their performance is nearly the same, while gemma-3-27b is much more efficient due to its smaller size.
We go one step deeper and analyze what types of aggregations these models struggle with in particular. Heatmap (B) in Figure 9 shows the F1 scores of the models on the five different types of aggregation. We can clearly see that all models struggle significantly with averaging, presumably because this involves a multi-step calculation. We can also see that both Llama-3.1-70B and Llama-3.3-70B are particularly good at min and max questions but struggle more with sum questions. Whereas, Qwen2.5-72B is the only model that performs reasonably well on sum questions which would explain how it outperforms the other models on aggregation questions in general. As expected, all models struggle with questions involving aggregation and particularly those that require averaging values. The metadata provided in HCT-QA allows for such granular analysis to be done for each individual model so that specific areas for improvement can be identified.
(RQ7) How are model size and performance correlated? Figure 10 shows how F1 score changes between different size variants of the same model. As expected, larger models consistently outperform their smaller counterparts. The gemma-2 models have the sharpest increase with $0.40\%$ F1 per 1B parameters (F1//B). Even when scaling to larger sizes like with Qwen2.5 and Llama-3.1 we see an increase of $0.34\%$ F1//B and $0.37\%$ F1//B , respectively. Hence, it is worth exploring fine-tuning smaller models for this task rather than simply relying on larger, resource-hungry off-the-shelf models.
Figure 10: F1 score across model size (parameters) within model variants
VI Conclusion and future work
In this paper, we presented HCT-QA, a comprehensive benchmark designed to evaluate language models on question answering over human-centric tables (HCTs), typically embedded within documents and web pages.
Our benchmark reveals several key insights. First, both LLMs and VLMs exhibit strong potential for question answering on HCTs. Second, finetuning LLMs substantially enhances performance, even when finetuned on only synthetic data, models demonstrated strong generalization to real-world HCTs. Finally, even the best-performing models (e.g., ChatGPT 4o) continue to face challenges with complex queries (e.g., involving aggregations) on complex HCT structures. Futhermore, our benchmark can be beneficial to other researchers in at least two ways: (i) Using the synthetic data generator to create training and test HCTs and QA pairs for new thematic domains in a cheap and quick manner. This is especially useful for training, finetuning, and evaluating new models. (ii) Adding question templates to the generator to cover more types of questions.
Future Work. Our benchmark can be extended along several complementary directions. First, we plan to broaden the scope of supported queries by enabling cross-HCT question answering, e.g., queries involving joins or unions across multiple HCTs. We also aim to incorporate queries that require more advanced OLAP operations, including CUBE and PIVOT. Initial experiments with two medium-sized LLMs indicate that these models are currently unable to handle such operators, highlighting substantial room for improvement.
Second, we intend to conduct deeper studies on fine-tuning strategies for both LLMs and VLMs to improve their structural understanding of HCTs. In addition, we will evaluate emerging systems that go beyond current LLM/VLM architectures, as these models may offer improved reasoning capabilities for complex tabular and visual data.
Third, we plan to enhance the linguistic diversity of the synthetic NL questions. One direction is use question paraphrasing capabilities using LLMs, which would increase lexical variation in the generated QA pairs. This extension will require robust validation strategies to ensure semantic equivalence between the original and paraphrased questions. Furthermore, the generator can be enriched with additional NL templates that capture implicit comparative terms such as most, least, highest, and lowest, instead of relying solely on explicit SQL-style operators such as MIN and MAX.
VII AI-Generated Content Acknowledgment
Here are the different places where we used AI models in our work:
  • LLMS were used to generate natural language questions and answers for the benchmark dataset as explained in Sections III-A.
  • The LandingLens from LandingAI tool was used for automated extraction and conversion of tabular data from PDF documents into HTML format (Section III-A).
  • Azure AI Document Intelligence was used to extract table images from PDF documents (Section III-A).
Appendix A HCT properties
A-A HCT Property definitions and examples
Definitions of HCT specific table properties are given in Table II and expanded upon with examples in this section.
Balanced Column Nesting: Column nesting can be likened to a tree-like data structure. Similar to the definition of a balanced tree, we can see that each branch of column names (from the top most parent till each leaf) has the same depth. In case of all the table shown in the Figure 14, the depth is 2.
Unbalanced Column Nesting: In this the depth of all branches in the nested column tree do not have the same depth. In Figure 13, the first table has ”Population characteristics, 2003” as the top most parent, and has branches with depth of 2; the second table has the parent ”Cultivable Land” with branches of depth 2 and 3 as well.
Symmetric Column Nesting: Column nesting is labeled as symmetric when the children of all the cells in highest level of nesting that has more than 1 cell are all identical. In the case of the first table in Figure 11, the top two cells have identical child branches (Total, Injuries-Slight, Injuries-Sever, Death). Whereas the second table has the parent cells ”Urban” and ”Rural” both with identical children values (”Women” and ”Men”). We do not consider column values such as ”Month” (in the first table) and ”Area/Age Group” in the second table as they do not exhibit any nesting. The first table also shows an example of where column nesting is unbalanced but is still symmetrical.
Figure 11: Example of Symmetric Column Nesting
Asymmetric Column Nesting: Asymmetric column nesting is when the conditions stated above from symetric column nesting are not met. This entails parent cells having different children cells (branches). In the first table in Figure 12 we see the the parent cell ”Total” has the children (”G.Total”, ”%Female”, ”F”, ”%Male”, ”M”) whereas the other 2 parent cells (”Non-Qataris” and ”Qataris” have a different set of children (”%”, ”Total”, ”F”, ”M”). Similarly, in second table, we see that the parent cells ”Age (percent)” and ”One race (percent)” have different children cells, thus making this column nesting assymetric.
Figure 12: Example of Asymmetric Column Nesting
Balanced Row Nesting: Very similar to column nesting, balanced row nesting entails that all row nesting branches have the same depth (in this case depth is measured left-to-right rather than top-down as with columns). The 3 tables in Figure 15 show this property as the children values (the second column from the left in each table) are identical across their parent cells (left most column in each table).
Figure 13: Example of Unbalanced Column Nesting
Figure 14: Example of Balanced Column Nesting
Figure 15: Example of Balanced Row Nesting
Unbalanced Row Nesting: This is when the conditions for balanced row nesting are unmet and branches of nesting in rows have different depths. In the table shown in Figure 16 we see that the branches ”Orca - Sys. message - Synthetic” (depth of 3) and ”Orca - Original model” (depth of 2) have different depths, thus they are unbalanced.
Figure 16: Example of Unbalanced Row Nesting
Symmetric Row Nesting: Similar to symmetric column nesting, this property entails each parent cell in the first column (left most) that exhibits nesting properties having identical children cells (branches). We can see in the first table in Figure 17 that all the parent cells (”HDFS”, ”Abase”, ”Faas”, ”Alibaba”) have the same children cells (”RMSE”, ”MAE”, ”p-risk”). Similarly, all the parent cells in the second table have the same children cells (”No.of Hotels”, ”No.of Rooms”, ”No.of Beds”).
Figure 17: Example of Symmetric Row Nesting
Asymmetric Row Nesting: Antithetical to symetric row nesting, this property entails the parent cells having different children cells. The table on the left in Figure 18 shows different children values for each of the parent cells (”Weight-based” with 5 children cells, ”Activation-based” with 2 children cells, ”Passport-based” with 2 children cells that have different values from all the others). The same can be seen in the table on the right as the children values are different for both parent cells.
Figure 18: Example of Asymmetric Row Nesting
Global Column Aggregation: This property entails that there is a column in the table that meets one of the following criteria: (1) is the aggregate of all columns in the table including or excluding local column-group aggregate columns or (2) is the aggregate of all local column-group aggregate columns present in the table. In this case an aggregation can mean any operation or function (e.g sum, max, min, average, different, or any other equation applied consistently across the column that uses values from other columns). For the first table in Figure 19, the column ”Total - Total” is a global aggregate as it meets both the conditions: it is the sum of all columns excluding ”Total - NonQatari” and ”Total - Qatari” which are local-group aggregate columns, and it is the sum of the aforementioned local-group aggregate columns (which are the only such columns in the table). Similarly, the ”Total” column in the second table and the ”Avg” column in the third table perform the sum and arithmetic mean (average) operations on all the other columns in their tables respectively.
Figure 19: Example of Global Column Aggregation
Local Column-Group Aggregation: A column that aggregates only a subset of numerical columns (including other local column-group aggregation columns) within the table. The first table in Figure 20 has two such columns. ”Cultivable Land - Cultivated Land - Total” is the sum of all the other children columns under the parent cell ”Cultivable Land - Cultivated Land”. ”Cultivable Land - Total” is the sum of the aforementioned local-group aggregate and the ”Cultivable Land - UnCultivated Lands” column. Similarly, in second table the ”Tonne kilometres performed - Total” column sums all the children columns of the parent cell ”Tonne kilometres performed” and thus is a local column-group aggregate column.
Figure 20: Example of Local Column-Group Aggregation
Explicit Column Aggregation Terms: A column where the cell value clearly indicates aggregation (e.g., “Total”, ”Percentage”, ”Sum”). For nested columns, if even one of the cells in the branch has an explicit cell value then the aggregation is considered as explicit. Figure 21 shows three tables that show columns having values like ”Change”, ”Average”, and ”Total”.
Figure 21: Example of Explicit Column Aggregation Terms
Implicit Column Aggregation Terms: A column that aggregates data, but the column name does not make this explicit. Figure 22 illustrates this ambiguity as the column ”Yield” in the first table is the ratio of the other two columns. Although arguable that the name ”yield” is indicative of a aggregation, we do not consider such terms that require some degree of domain familiarity to be understood. Thus, it is considered as an implicit aggregation term. The second table presents a more definitive example of this property as the column ”Pakistan” is the summation of all the other columns (which represent different parts of Pakistan). Since it is difficult to understand that this is an aggregation without having prior knowledge about Pakistani geography, we consider this an implicit aggregation term.
Figure 22: Example of Implicit Column Aggregation Terms
Global Row Aggregation: A row that aggregates (using any function) all other rows in the table. Figure 23 presents examples of such aggregate rows that use the sum and arithmetic mean aggregations.
Figure 23: Example of Global Row Aggregation
Local Row-Group Aggregation: A row that aggregates a specific subset of rows (including other row-group aggregate rows) but not the entire table. Examples are shown in Figure 24 where the highlighted rows aggregate (sum in this case) subsets of rows. This property in particular has various visual manifestations: indentation in the first table, numbered rows in the second, explicit ”total” cell in third.
Figure 24: Example of Local Row-Group Aggregation
Explicit Row Aggregation Terms: A row where at least one cell explicitly indicates aggregation (e.g., “Total”, “Average”, “Sum”). This is clearly shown in Figure 25.
Figure 25: Example of Explicit Row Aggregation Terms
Implicit Row Aggregation Terms: A row that likely aggregates data, but none of its textual cells explicitly say so. Figure 26 illustrates this. As stated earlier, this property is often shown using numbered cells, indentation, domain specific terms (such as ”Gross Domestic Product”, font stylization, and other visual characteristics.
Figure 26: Example of Implicit Row Aggregation Terms
Row Group Label: A row that acts as a heading or title for a set of rows above or below it. See Figure 27 for examples. The first table has the row group labels ”Pakistan” and ”Punjab” respectively. The second table has ”C. Oil and Non-Oil”.
Figure 27: Example of Row Group Label
Split Header Cell: A single header cell split diagonally and containing two distinct values. This visual property is shown in Figure 28. In the first table the term ”Sector” describes the column names and ”Occupation” the values in the column below it. Similarly, in the second table, ”Year” describes the other column names, and ”Type of License” is the column name for that particular column.
Figure 28: Example of Split Header Cell
Standard Relational Table: A table that does not exhibit any of the defined HCT properties — no nesting, no aggregation, no labels. Figure 29 shows examples of this simple table form.
Figure 29: Example of Standard Relational Table
A-B Table Annotation Tool
Figure 30: A user-friendly tool for annotating/viewing HCTs and their properties.
We developed a user-friendly tool for annotating tables and setting their properties as illustrated in Figure 30. The left side visualizes one table at a time, while the right side lists the possible allowed properties.
As an example, the illustrated table in Figure 30 contains “Column Nesting”. And since the child column names (e.g., {“Grand Total”, “Females”, “Males”}) across all groups are at the same level (same depth from the top parent cell), it is considered a “Balanced Column Nesting”. Similarly, since all branches of the parents have identical structure and cell values, the column nesting is considered “Symmetric”. Similar rules apply to the “Row Nesting” property. For example, the table in Figure 30 does not containing any row nesting, thus that property is not selected. The table also contains column aggregation at both levels, local (e.g., “Total” under “Non-Qataris”, “Males” under “Grand Total”), and global (e.g., “Total” under “Grand Total”). At the row level, this table contains only global aggregation (the last row). We also keep track of whether the aggregation columns carry explicit aggregate-meaningful labels (e.g., SUM, TOTAL, AVERAGE, MIN, MAX, etc.) as in the table illustrated in Figure 30.
Certainly, given the compound complexity among the different properties, we could have captured the properties at a more coarse or grained granularity. Nevertheless, we opt for this level of details because it sufficiently captures the underlying complexity of the HCT objects, and would provide good insights on how these properties correlate to QA performance without being dragged into too many features which requires high dimensional analysis. the annotation tool was developed in-house and we plan to release it on GitHub [^1] https://github.com/shahmeer99/HCT-QA-Benchmark/ page soon.
A-C Understanding the distribution of table properties
The Figure 31 and Figure 32 show the distribution of the top-$5$ main and detailed real table properties per dataset. The property distribution is very different between datasets. For instance, column and row nesting are present in QNPC and ArXiv datasets (Figure 31(a) and (d)) but no row nesting appears in US C. or Pak C. datasets. Column and row aggregation appear in all datasets except in ArXiv with only a small proportion of row aggregation. QNPC datasets have the most various and complex combinations of main properties, while ArXiv have the least various.
Regarding detailed table properties in Figure 32, QNPC and US C. show quite complex and various combinations compared to Pak. C and ArXiv. While balanced column nesting is a typical property for all the tables, it is the major one in Pak. C and ArXiv while it is more balanced in the other two datasets.
Refer to caption Refer to caption
(a) QNPC (b) Pak. C
Refer to caption Refer to caption
(c) US C. (d) ArXiv
Figure 31: Distribution of main table properties for each real dataset.
Refer to caption Refer to caption
(a) QNPC (b) Pak. C
Refer to caption Refer to caption
(c) US C. (d) ArXiv
Figure 32: Distribution of detailed table properties for each real dataset.
Appendix B Synthetic Data Generation
B-A Synthetic Generator
In the following subsections, we detail the configuration templates we used for setting the domain vocabulary (Section B-B) common to both HCTs, relational tables, SQL queries, and natural language questions. Then we detail the templates used for generating the synthetic HCTs (Section B-C), and those used for generating the SQL queries and the Natural Language questions (Section B-D).
The codes of the synthetic generator and the instructions for installing and running them are available at https://github.com/qcri/HCTQA-Benchmark/blob/main/synthetic_data_generator/README_SYNTHETIC_GENERATOR.md.
B-B Domain vocabulary templates
Figure 33: Domain vocabulary for generating synthetic tables and questions. Examples of $T_{REL}$ (a) and a corresponding $T_{HCT}$ (b) related to the amount of food import-export in some place. (c) JSON format to describe Food Import-Export domain vocabulary used to generate these tables (see Section B-C) and the related questions (see Section B-D). Not all vocabulary is used in $T_{REL}$ and $T_{HCT}$.
In this section, all terms refer to Figure 33.
Our synthetic data relies on domain vocabulary naming the dataset’s attributes and values. A synthetic $T_{HCT}$ (Figure 33b) is obtained by pivoting a relational table $T_{REL}$ (Figure 33a) composed of at least two nominal attributes and a single numerical attribute always named Value.
There are two types of nominal attributes:
  • Independent attributes like Import-Export, Years, or Quarters can be organized as row or column headers at any nesting level in $T_{HCT}$.
  • Hierarchical attributes like food Category (e.g. Meat, Cereals, Dairy, Beverage) and food Item (e.g., Cream, Soda, Milk) are dependent on each other. If both attributes appear in the HCT, the food Item must be displayed at a lower level than the food Category, i.e. closer to the inner part of the HCT, and all must appear together either in row headers or column headers.
Domain vocabulary is encoded in a JSON file (Figure 33c) with two main keys: data and values.
  • data refers to a list of attributes (e.g., ImportExport, Years, Quarters) or groups of attributes (e.g., Food), each described with three keys:
    • code is the primary key used in table and question templates to identify the independent attributes (e.g., ImportExport or Years) and the hierarchical groups of attributes (e.g., Food).
    • names is the list of attribute names displayed in $T_{REL}$ column headers but typically hidden (implicit) in $T_{HCT}$ and $Q_{NL}$. It is a list with a single name for independent attributes, or with as many names as levels for hierarchical groups of attributes (e.g., Food), ordered from top (outer) to bottom (inner) levels (e.g., Category, Item).
    • values is the list of values (e.g., 2015, 2016... for the attribute Year). It can also be the hierarchy of values of a group of attributes. For instance, in Food group, Meat is a value of the attribute Category, and Beef, Chicken…, are values of the attribute Item, whose parents are Meat in the Food hierarchy. Attribute values are always displayed as row or column headers in $T_{HCT}$.
  • values refers to the numerical values to be generated in both the core part of $T_{HCT}$ and the Value column of $T_{REL}$. This object contains shortcut keys (e.g., realUnit or intSigned) to generate specific ranges and types of numbers when called in table templates. The value associated with such shortcuts must be a list of two numbers $[m,M]$ representing the minimum $m$ and maximum $M$ possible values. If any of these min and max is a decimal number, the generated value will be a real number with two decimal digits drawn uniformly in this interval. If both min and max are integers, the generated value will be an integer drawn uniformly in this interval. The generator will ensure no replica in the generated values, reals or integers, possibly overflowing the max value if asked to generate more numeric values than the range allows.
Customization of domain vocabulary All terms in domain vocabulary are customizable except the JSON keys data, code, names, and values, following the rules above. The number of decimals of generated real numbers is specified by NUM_DECIMAL_DIGITS_REAL_FORMAT in config.R. The domain vocabulary is stored in the file PARAM_semantics.json.
B-C Table templates
The generic table templates described in the file PARAM_tableTemplate.json are used to generate a (large) set of $T_{REL}$ and $T_{HCT}$ tables from the domain vocabulary (See Section B-B).
This template contains the following keys:
  • replica is an integer setting how many random variants of each initial table template from the list tables will be generated.
  • shuffle is a list of column and row shuffling commands that will be applied to all the initial table templates listed in tables. It is used to vary the structure of the initial table. none is the default; row mixes only row attributes; cols mixes only column attributes; rowscols mixes rows and columns, keeping them as rows or columns, respectively; all mixes across rows and columns (some columns can become rows and vice-versa, except for hierarchical attributes which stay grouped in rows or columns). The shuffling process is a random permutation of the rows or columns to be mixed.
  • col_row_levels is a list of authorized depths for nested rows and columns. It must be a list of strings ”dcdc_drdr” indicating the depth of the column header ($dc$), and of the row header ($dr$), with $dc$ and $dr$ in $\{1,2,3\}$.
  • col_row_name_pos is a list of authorized displays for row and column attribute names. It must be a list of strings ”C_R” indicating where the column (C) or row (R) attribute shall be displayed. R and C can be none to hide the attribute name, left to show it on the left side of the header, or top to show it at the top of the header. (Place-holder for a future version, ignored by the generator).
  • col_row_agg_pos is a list of authorized positions for aggregate rows or columns. It must be a list of strings ”C_R” indicating on which side of the header the aggregate column (C) or row (R) shall be displayed: R can be set to top or bottom, C can be set to left or right, and both can be set to none (Default) to hide the corresponding aggregate.
  • row_format can be set to new to use one column for each level of row header, or indent to create an empty row for all level $L$ and indent all rows at level $L+1$. The header of the empty row is the attribute value for that row. An aggregate can fill this row if col_row_agg_pos is set to *_top.
  • tables is a list of table content specifications (values, headers, and aggregates) with coherent domain vocabulary. Each initial table template in the list has the following keys:
    • valueName is the title displayed at the top of the $T_{HCT}$ table (e.g., ”Food import-export in tons”)
    • values is either a shortcut name referring to one of the keys (e.g., realUnit) stored in values in the domain vocabulary (see Figure 33), or directly an interval of the form $[m,M]$ following the description of the main key values in Section B-B
    • valueUnit (Place-holder for a future version, not yet implemented) is a character string to be added to give units to values in the core part of the table (e.g., $\%$ sign after each value in the core part of $T_{REL}$).
    • rowCodes is a list of one or more of the data code keys (e.g., Food or Quarters) listed in the domain vocabulary. The generator will pick among these codes to generate the rows of $T_{HCT}$, and one part of the nominal attributes of $T_{REL}$.
    • rowSamples is a list containing as many intervals $[m,M]$ as there are codes listed in rowCodes, in the same order. Each interval specifies the minimum $m$ and maximum $M$ number $n$ of values to select among the list $L$ of values of the specified attribute. The generator first picks $n$ at random in $[m,M]$, then selects $n$ consecutive values in $L$ starting at a random position in $L$ but ensuring there are always $n$ values in the result if the list $L$ is long enough, or return $L$ itself otherwise. For instance, let’s consider an independent attribute Years with a list of values: $L=[2001,2002,2003,2004,2005,2006]$, if we set rowSamples to $[1,5]$, and the generator randomly picks $n=3$, it then picks a random index $i$ between $1$ and $4$, e.g., $i=2$, to generate the final subset $[2002,2003,2004]$ of years to use in the table. For hierarchical attributes, the same process applies recursively to all levels of the hierarchy, starting from the top level. If the interval is set to $[0]$ instead, all the values of the corresponding independent attribute are generated, or all the values at all hierarchy levels are generated for hierarchical attributes. Note that the col_row_levels parameter limits the depth of row and column headers, so the attributes finally displayed in $T_{HCT}$.
    • colCodes is the same as rowCodes but for columns. The generator will pick among these codes to generate the columns of $T_{HCT}$, and the other part of the nominal attributes of $T_{REL}$.
    • colSamples is the same as rowSamples but for columns.
    • agg_name1 is the name of the aggregate to display in the table.
    • agg_fun1 is the aggregate function to use among sum (default), avg (average), min, max.
These generic table templates PARAM_tableTemplate.json are used with the domain vocabulary PARAM_semantics.json to generate individual templates using the script 1_geneAllTablesJSONfromPatterns.R. The individual templates generated and described below are stored in the file PARAM_tablesToGenerate.json.
Customization of generic table templates These generic table templates are customizable except for all the JSON keys and following the rules above.
Figure 34: Two examples of individual table templates (a,b), each with two random samples (1, 2) of $T_{REL}$ (a’, b’) and $T_{HCT}$ (a”, b”). Templates a and b differ by the depth of column headers (col_row_levels), row indentation (row_format), presence and position of aggregates (agg_pos1). The sampled $T_{HCT}$ (1, 2) of a given template (a or b), further differs by the randomized order of the attributes (shuffle), number and names of the attribute values (sample), core numerical values (values), and presence of table cell borders (random).
An individual table template is a JSON object instantiating one combination of the parameters shuffle, col_row_levels, col_row_name_pos, col_row_agg_pos, and row_format common to all initial table templates, and set in the generic table template. It contains the following keys:
  • name is the identifier of the family of $T_{HCT}$ and $T_{REL}$ tables specified by this individual table template.
  • replica is identical to replica set in the generic table template.
  • agg_fun1 is identical to agg_fun1 set in the generic table template.
  • agg_name1 is identical to agg_name1 set in the generic table template.
  • values is identical to values set in the generic table template.
  • valueName is identical to valueName set in the generic table template.
  • shuffle is one of the shuffling commands listed in shuffle of the generic table template.
  • row_format is one of the row format commands listed in row_format of the generic table template.
  • columns and rows are lists of attribute codes obtained after filtering the first $dr$ row attributes and $dc$ column attributes and setting the aggregate position using the col_row_levels and agg_pos1 parameters from the parameter combination of this instance. sample are identical to colSamples and rowSamples of the corresponding attribute set in the generic table template.
Finally, replica instances of $T_{HCT}$ are generated using the script 2_geneHCTfromJSONtables.R which takes as input the files PARAM_semantics.json, PARAM_tableTemplate.json, and PARAM_tablesToGenerate.json. It returns HTML files named [name]_n_DB.html ($T_{REL}$) and [name]_n_HCT.html ($T_{HCT}$) with $n$ the replica number.
This process draws samples from each individual template, randomizing all the remaining parameters: the shuffle (stochastic) command is applied to rows and columns attributes, then the values of each attribute are sampled using the sample parameter, and the values in the core of the table are sampled from the values parameter. The absence or presence of cell borders in the $T_{HCT}$ layout is randomized for each final sample.
Two examples of individual templates with two of the five sampled $T_{REL}$ and corresponding $T_{HCT}$ are shown in Figure 34.
Customization of individual table templates There is no possible customization of these individual table templates automatically generated from the generic table templates.
B-D Natural Language question templates
We defined $15$ SQL templates in Table III and transcribes them into the HCT domain using $15$ domain-specific Natural Language templates (Figures 36 and 37 show all the natural language question templates in JSON format related to the seven generated HCT domains).
Figure 35: Example of a food import-export domain $T_{REL}$ (left) and $T_{HCT}$(right) to explain $Q_{NL}$ templates. All the values (dashed color frames) of columns A, B, C, and D of the relational table are factorized as row and column headers in the HCTs (HCT1, HCT2). In contrast, the whole value column E (blue frame) is reproduced in the core area of the HCT (HCT3). In this instance, the ground-truth answer AA to the question "What is the amount of Export of Milk in 2017?" ($Q_{NL}$) related to $T_{HCT}$ is "57" (Dashed blue frame). This answer AA is obtained by running the SQL query $Q_{SQL}$ paired with $Q_{NL}$ onto $T_{REL}$.
For instance, the $Q_{NL}1$ template below transcribes the $SQL_{1}$ template (Table III) for the Food Import-Export domain HCT displayed in Figure 35c:
What is the amount of_ \$Import_Export of__ \$Item__ \$Category in__ \$Quarter==of==\$Year?
This template is instanciated with detailed row and column attribute values of the HCT, to generate the following natural language question:
What is the amount of Export of Milk in 2017?
Notice that not all the table attributes may be displayed in the HCT, and the values vary due to random selection of the HCT instance.
This question will be used as a prompt for the evaluated LLM along with the $T_{HCT}$ instance. The corresponding SQL query instance is:
SELECT Value FROM DBdata
WHERE ((Import_Export = ’Export’ AND Year = ’2017’))
AND ((Category = ’Dairy’ AND Item = ’Milk’));
and it will be run on the $T_{REL}$ instance paired with $T_{HCT}$ to get the ground-truth answer A=57A="57".
Here are the $15$ types of natural language questions $Q_{NL}$ generated for the table $T_{HCT}$ in Figure 35 instantiated from the template of Figure 36a (”,” and ”;” in answers, are column and row separators respectively):
  • $\mathbf{Q_{NL}1}$: one column, one row selection: ”What is the amount of Export of Milk in 2017?” $\rightarrow$ ”57”
  • $\mathbf{Q_{NL}2}$: one column, many rows selection: ”What is the amount of Export of Butter, Coffee, Milk, or Tea in 2017?” $\rightarrow$ ”57; 146 ; 980; 749”
  • $\mathbf{Q_{NL}3}$: many columns, one row selection: ”What is the amount of Export of Tea in 2017 or 2018?” $\rightarrow$ ”749; 622”
  • $\mathbf{Q_{NL}4}$: $Q_{NL}3$ + expression not in the table: ”What are the total and minimum amounts of Export or Import of Milk in 2017 or of Export of Milk in 2018?” $\rightarrow$ ”1442,57”
  • $\mathbf{Q_{NL}5}$ (opt): $Q_{NL}4$ + expression from the table: (only if an aggregate exists in the HCT) ”What is the average amount of Export or Import of Coffee?” $\rightarrow$ ”533.75”
  • $\mathbf{Q_{NL}6}$: many rows, many columns selection: ”What is the amount of Export or Import of Butter, Cream, or Tea in 2017 or of Export of Butter, Cream, or Tea in 2018?” $\rightarrow$ ”444; 146; 443; 736; 44; 821; 751; 749; 622”
  • $\mathbf{Q_{NL}7}$: $Q_{NL}2$ + aggregation not in the table: ”What are the total and minimum amounts of Import of Butter, Coffee, Cream, Milk, or Tea in 2017?” $\rightarrow$ ”2996,266”
  • $\mathbf{Q_{NL}8}$: $Q_{NL}6$ + grouping per column + reporting: ”What are the minimum and maximum amounts for each Import-Export and Year, of Export or Import of Butter, Coffee, Cream, Milk, or Tea in 2017 or 2018? Please, report the corresponding Import-Export and Year.” $\rightarrow$ ”Export,2017,44,980; Export,2018,121,821; Import,2017,266,799; Import,2018,179,879”
  • $\mathbf{Q_{NL}9}$: $Q_{NL}2$ + grouping per row: ”What is the minimum amount for each Category, of Import of Beverage or Dairy in 2018?” $\rightarrow$ ”673;179”
  • $\mathbf{Q_{NL}10}$: $Q_{NL}6$ + grouping per rows + reporting: ”What is the minimum amount for each Category, of Import of Beverage or Dairy in 2018? Please, report the corresponding Category.” $\rightarrow$ ”Beverage,673; Dairy,179”
  • $\mathbf{Q_{NL}11}$: $Q_{NL}6$ + grouping per rows and columns + reporting: ”What is the minimum amount for each Category, Import-Export, and Year, of Import of Beverage or Dairy in 2017 or of Export or Import of Beverage or Dairy in 2018? Please, report the corresponding Category, Import-Export, and Year.” $\rightarrow$ ”Beverage,Export,2018,121; Beverage,Import,2017,266; Beverage,Import,2018,673; Dairy,Export,2018,443; Dairy,Import,2017,444; Dairy,Import,2018,179”
  • $\mathbf{Q_{NL}12}$: $Q_{NL}2$ + top-K + row filter: ”What are the bottom 5 amounts of Export of Beverage or Dairy in 2018?” $\rightarrow$ ”121; 443; 586; 622; 821”
  • $\mathbf{Q_{NL}13}$: $Q_{NL}2$ + row filter + ordering: ”What are the amounts ordered by increasing values of Import of Beverage or Dairy in 2017?” $\rightarrow$ ”266; 444; 736; 751; 799”
  • $\mathbf{Q_{NL}14}$: $Q_{NL}2$ + operation on column: ”What are the Category and Item for which the amount of Import in 2018 is greater than 513.3?” $\rightarrow$ ”Dairy,Milk; Beverage,Coffee; Beverage,Tea”
  • $\mathbf{Q_{NL}15}$: $Q_{NL}14$ with condition + reporting: ”What is the amount of Export in 2018 of Category and of Item for which the amount of Import in 2018 is greater than 513.3? Please, report the corresponding Category, Item, and amount of import-export.” $\rightarrow$ ”Dairy,Milk,586; Beverage,Coffee,121; Beverage,Tea,622”
Figure 36: Synthetic natural question templates for the Accidents (a), Constructions (b), Food (c), and Graduations (d) domains.
Figure 37: Synthetic natural question templates for the Pollution (a), Students (b), and Weather (c) domains.
Figures 36 and 37 show the natural language question templates related to the seven generated HCT domains, as JSON objects with the following keys:
  • tableName is identical to the valueName key in the generic table template (see Section B-C).
  • rowColCodes gather the rowCodes and colCodes lists of attribute codes used in rows and columns of the generic table template.
  • valueMeaning is a text characterizing a value in the table continuing the sentence: ”A value in this table represents the…” (e.g., ”amount of import-export”).
  • simplifyNested lists all the names group of hierarchical attributes used in the generic table template.
  • template_report is a sentence to report a specific list of attributes \$REPORTATTR selected for the query result in templates $8,10,11,15$.
  • template_N a list of templates for questions of type N (See the list of $15$ $Q_{NL}$ templates above).
    • A template is a natural language sentence with some place-holder variables to be filled by the generator depending on the table instance content and the type of question. All the templates must show all the possible attributes set in the generic table template, but the question generator will randomly pick the ones it will use, depending on the template type; the others will be absent.
    • We use \$attribute_name to specify variables that will be automatically replaced by the generator with values from that attribute present in the table instance (e.g., \$Category could be replaced by Meat, Dairy, Cereals, or Beverage). These attribute_name must appear in the rowColCodes list for independent attributes, and in the simplifyNested list for groups of hierarchical attributes.
    • All words in capital letters \$RESERVED and their connected words should not be modified, and must appear as-is in their respective template.
    • In templates $4,5,7,8,9,10,11$, an expression \$EXPR will be replaced by one of the aggregation names agg_name1 (min, max...) in the generic table template. Some question can ask several expressions at once, like the min and the max, calling for a plural version of the sentence. We use // to separate the singular and plural versions of a word that the generator will pick automatically depending on the number of expression terms generated. For instance, the template $7$: What is/are_the_\$EXPR_amount/amounts of_\$Import_Export of... can generate ”What is the minimum amount of Export of…” or ”What are the minimum and maximum amounts of Export of…”
    • In templates $8,9,10,11$, \$GROUPBY will be replaced by for each followed by a list of attribute names. For instance, the template $9$: What is/are_the_\$EXPR_amount/amounts \$GROUPBY of_\$Import_Export of__\$Item__\$Category in__\$Quarter==of==\$Year ? can generate the question ”What is the minimum amount for each Category, of Import of Beverage or Dairy in 2018?”
    • In template $12$, top/bottom_\$TOPK will be replaced by an integer and either top or bottom automatically selected, e.g., ”What are the top 5 amounts of Import of…
    • In template $13$, the group \$ORDERBYDESC_decreasing/increasing_values must be used as-is and can generate the question ”What are the amounts ordered by increasing values of Import of Beverage or Dairy in 2017?” or ”What are the amounts ordered by decreasing values of…”
    • In template $14$, the word \$OPPATTR will be replaced by a list of attribute names, and the word \$OPANDVAL by either greater than or lower than and a certain value, leading to a question like: ”What are the Category and Item for which the amount of Import in 2018 is greater than 513.3?”
    • We use _ to link words like of, in to their attribute. If an attribute is not present in an instance of a question, all the words connected to it with _ will be absent too (e.g., of_\$Import_Export).
    • We use __ to link several attributes together. For instance, we use __\_\_ in of__\$Item__\$Category so if \$Item is absent, the of word stays connected to \$Category like in of Meat. This pattern can generate of Beef and of Meat, but will not generate of Beef Meat because these attributes are part of the same hierarchical group.
    • We use == to link several attributes separated by a word that only exist if both surrounding attributes are present. For instance: we can say in Q2 of 2022 or in 2022 or in Q2. Thus, we tell that of is only present if both \$Quarter and \$Year are present together, by coding in__\$Quarter==of==\$Year.
    • We use ((not_an_attribute)) to ensure the presence of a word which is not an attribute. For instance, we want to enable the generation of the number of English male students but also the number of students if the \$Nationality and \$Gender attributes are absent from the generated question. As students is not an attribute value, we use the template of_((students)) of__\$Nationality__\$Gender__students to enable this option. The generator will only produce the first part of_((students)) if the word students does not appear in the rest of the generated sentence (See Figure 36b).
    • The order of the attributes used in the template is always preserved.
    • Additional rules, a template shall never start with a \$ and a space must be set before the final question mark. Attribute names composed of several words separated by - should replace it by _.
All $14$ templates are repeated with variations for the $7$ domains in the PARAM_NLquestionTemplates.json file, guiding the creation of the same templates for new domains. Note that the question $Q_{NL}15$ is automatically generated from templates $13$ and $14$.
The script 3_geneSQLandNLQfromJSONtemplates.R reads the question template file PARAM_NLquestionTemplates.json and the other parameter files, to generate the natural language questions $Q_{NL}$ and corresponding SQL queries $Q_{SQL}$ for each question template as well as their ground truth answers $A$ from the tables generated from PARAM_tablesToGenerate.json.
Customization of question templates These question templates are customizable except for all the JSON keys and following the rules above.
Appendix C Additional Experiments and Details
C-A Additional Trend Analysis & Observations
(RQ8) How does model verbosity affect the output correctness? Our initial hypothesis was that if models simply take the kitchen-sink approach and output more values in their answers then they would have a higher probability of matching values in the ground truth and thereby inflate their recall scores. As shown in Figure 38, there is no clear correlation between answer length and recall. Most models, including the best performing ones, seem to be clustered around the middle on the verbosity scale. The average number of values in the ground-truth answers for all queries is  4, yet most models tend to provide between 2-3 values in their answers. This suggests that models may in fact be conservative in their answer verbosity as which is antithetical to the overly verbose nature of these generative models.
Figure 38: Recall Score vs. Verbosity for Text-Only Models on Real World HCTs.
(RQ9) How do model perform on different synthetic QA templates? The synthetic queries we generated have 15 types including various types of selection, projection, expression, and aggregation. The model performance per template type is shown in Figure 39.
Figure 39: Recall Score for Text-Only Models on all Synthetic HCTs per Question Template.
As expected, models tend to do best on selection queries be it single-cell or multi-cell selection. This is clear as templates 1, 2, and 3 have the highest average recall across all models. Models tend to do better on queries involving expressions on columns (templates 5, 14) than queries involving aggregations (template 7). In fact, models struggle the most with template 7 which requires doing an actual aggregation (rather than retrieving an existing one from the HCT). Models like LLaMA-3.3-70B significantly outperform even models like GPT-4o and GPT-35 on this template. It is also evident that models clearly have their strengths and weaknesses in terms of templates. GPT-4o gets 58% on template 11, whereas other models like LLaMA and Mistral do not even cross 10%. However, this dominance completely disappears in Templates 4, 7, and 9 where the GPT-4o fails to score more than 8% (compared to other models that get up to 42%). There are also unexpected trends across model generations. Gemma-2-27B scores a 52% recall on template 14 which is only topped by GPT-4o and Qwen2.5 72B, both of which are relatively larger models. However, on the exact same template (14), Gemma-3-27B scores exactly half of its predecessor with a recall score of 26%. This shows that newer generations of models are not necessarily improving their capabilities on this task across the board.
(RQ10) In what cases do VLMs perform better than LLMs? Recent advances in the image understanding and reasoning capabilities of VLMs make them a suitable candidate for the task at hand. VLMs also allow to skip the extraction and conversion to CSV process as they can directly consume the images of the tables from their source documents. On average, VLMs score 25.5% recall whereas text-only LLMs score 25% as shown in (Table IV). GPT4o when prompted using the text-modality achieves an average recall of 61.6%, compared to 62.3% when using the image-modality. This illustrates the progress vision models have made in terms of document understanding and reasoning tasks such as this one. We have established that HCTs inherently have visual characteristics that are crucial to understand the content in the table. A lot of those visual clues are lost when the HCT is converted to text format (regardless of which text format its represented in). VLMs are able to overcome this information loss by taking advantage of those visual cues. This is further shown by the top performing small sized VLM model, Pixtral-12B with a recall score of 45.9% which is nearly 8% better than the best performing text-only LLM of that size range (Qwen2.5-7B with recall 38%). Furthermore, Microsoft’s Phi-3-Vision-128k (image-modality), which is only a 4.2B parameter model outperforms every text-only LLM in the small size category (many of which are 2x in size). These results clearly highlight the potential of VLMs in visual-reasoning tasks such as HCT-QA. Another interesting observation is that VLMs struggle considerably more than LLMs with larger HCTs. GPT4o performs almost identically in both the text and vision modalities except in the US CENSUS dataset which on average has 437 cells per table (nearly 3x more compared to the other datasets. This trend can also be seen in the smaller VLMs as most top performing models drop in performance when dealing with US CENSUS HCTs.
Figure 40: LLMs vs. VLMs on all Real World HCTs — Recall Score.
C-B Prompting Details & Experiments
For certain models (like Gemma-2) that do not accept a system prompt we add the system prompt as a prefix to the user prompt and send it as one prompt with the ”user” role.
C-B1 Prompts
Prompts are used with LLMs to give them context and role, and to format the results to allow automatic evaluation.
System Prompt for Inference You are a table question answering assistant. You are given a table and a question. Your task is to answer the question based on the information in the table. The table structure may be complex and not a standard relational table, so try to understand the structure of the table when answering the question. If the question cannot be answered using information from the table, return ‘No Answer‘. Do not provide any explanations, equations, code, or text explaining intermediate steps in figuring out the answer. Return only the final answer itself. If there are multiple values, separate each value with the ‘ || ‘ token.
User Prompt for Inference Answer the question given the input table. Do not provide any explanations, equations, code, or text explaining intermediate steps in figuring out the answer. Return only the final answer itself. # INPUTS: Table: ‘{table}‘ Question: ‘{question}‘ # OUTPUT: Answer:
C-B2 0-shot vs. 1-shot prompting
The majority of the models we used in our experiments can handle context lengths of 8,000+ tokens. This means that it is possible to do 1-shot prompting with a full table-question-answer triplet as the shot. This approach could help leverage in-context learning to improve the model performance. We experimented on a small set of 1,000 queries using Real World HCTs from all four data sources. Our 1-shot prompt consisted of a fixed example for all the questions in the test set. We avoided testing the models given an example question on the same HCT the test question was on as that reflects a very unrealistic scenario. The results are shown in Table VIII. We can see that adding the 1-shot example does not significantly affect the recall score as for gemma-2-9b-it it only increases by 1.3% and for Llama-3.1-8B-Instruct it actually decreases by less than 1%. Upon manual inspection, we found that the models were able to follow the output format stated in the prompt (values split by ‘ || ‘ and no explanations) in the 0-shot scenario almost perfectly. As the 1-shot example was on a different table, it provided no ”learning” for the model to answer the test question it was being asked. Due to these results, our remaining experiments use 0-shot prompting.
TABLE VIII: 0-shot vs 1-shot Prompting - Recall Score
Model 0-Shot Recall 1-Shot Recall
gemma-2-9b-it 0.367 0.373
Llama-3.1-8B-Instruct 0.247 0.239
C-C Evaluation Metrics
We experimented using Meta-Llama-3-70B as judge model to gauge the correctness of a model’s answer. The prompt given to the judge model outlined a clear set of instructions: (1) minor spelling mistakes or typos in textual values can be ignored, (2) rounding numerical values is fine if done reasonably, (3) question must be answered fully in order to be ”correct”. We compared this model’s ”correct” labels to our CC Score (Complete Containment, which we defined as 1 when Recall Score == 1, else 0) on a subset of 500 real-world HCT QA pairs. The comparison is shown in Figure 41. It is evident that the model’s judgment and the CC score closely align with each other. Since running an LLM-as-a-Judge is expensive in terms of computational resources, we opt to use Recall, Precision, and CC Score as our evaluation metrics.
Figure 41: CC Score vs. LLM-as-a-Judge Score on Real World HCTs.
C-D VLMs vs LLMs
We can see from the results in Table IV that all VLMs are sensitive to table size. HCTs arXiv and Pak have relatively lower number of cells and number of words on average and vision models tend to do the best on these tables. It is evident that all vision models struggle with US Census tables (which are the largest tables on average with good margin). Gpt-4o scores 39.6% recall on US Census which is less than majority of medium sized text-only models which are considerably smaller in size.
This is further clear by the fact the gpt-4o-vision which outperforms its text-only version (gpt-4o) in all 3 datasets (QNPC, Pak, arXiv) but only scores a recall of 39.6% for US Census which is 20% less than that of gpt-4o (text-only) at 59%. Some examples of cases where gpt-4o gets answers a question correctly in one modality but gets it wrong in the other are provided in Figure 42 and Figure 43.
Figure 42: gpt-4o-vision is correct but gpt-4o text-only is wrong.
Figure 43: gpt-4o-vision is wrong but gpt-4o text-only is correct.
C-E Finetuning Details
We fine-tuned 2 models (Meta-Llama-3-8B-Instruct and FanarC-9B) using LLaMA Factory [^2] https://github.com/hiyouga/LLaMA-Factory using the following hyperparameters:
Finetuning Configuration Summary Finetuning Type: Full-parameter finetuning Max Sequence Length: 4096 tokens Precision: bfloat16 (bf16) Training Epochs: 1 Batch Size: 2 per device Gradient Accumulation: 2 steps Learning Rate: 5e-7 Learning Rate Schedule: Cosine with min LR = 5e-8
The dataset itself, with the exact train + validation + test splits used in our finetuning experiments, can be found on HuggingFace [^3] https://huggingface.co/datasets/qcri-ai/HCTQA.